A Complete Bibliography of Publications in Numerische Mathematik (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 November 2023 Version 3.32

Title word cross-reference

(m, s) [AdST07]. 1 [CM06, CFJ03a, CFJ03b, EZ08, FMS07, Mic02]. 2 [KS07]. 3 [Deu09, KNWN09, LSW05, RS04]. 8 [FG02, KP08, ZK00]. ^{0} [BW05]. BV [EK00]. BV_{t} [BKRT04]. C^{0} [AD01]. C^{1} [QQ03, Nov01]. cos(z) [VC01]. D [AD01, Hac08]. D^{2} [AdST05]. F E [BH03b]. G^{2} [FG02]. H [AFW00, BH03b, Lac00, CMS01, GHK02]. H^{2} [LSW02]. H_{0}(curl; ω) [HWZ06]. HH [GKB09]. ILU [Yun03]. ∫_{a}^{b} f(x)e^{iωg(x)} dx [Xia07]. J_{ν}(νa) [WW02]. k [ZOS00, dFP05]. L^{1} [Ber06, CDRG^{+}07, GP08, LT01]. L^{2} [MSS02, McS02]. L^{8} [BH03b]. L_{2} [Ste01a, Ste02a]. L_{p} [Gei07]. L_{p}(0, T) [SW01]. M [Bou07]. M(a, b, x) [Mul01]. R^{3} [Nov01, GS02]. V_{b,g}^{⋆}(ker(C)) [Chu03]. p [AG00b, ABH05, Bab00, BP08, BH05a, BH07, BS06b, CS00a, EL05, GH04, Heu98, Heu01a, Heu01b, Kim05a, KS07, LY01, MBS00, TS00, XG00]. P_{1} [Osw08]. Ψ [ZCW06]. Q_{1} [AM06]. QR [VVM09, CLM03]. R^{2} [BG01]. R^{3} [Sid05]. sin(z) [VC01]. θ [AG01]. V [AD07]. W^{1,p} [AM06]. W^{2,∞} [AdST05]. ZZ [ZL03].

-adaptive [CMS01]. -Approximations [MSS02]. -coefficients [BH03b]. -continuous [FG02]. -cycle [AD07]. -cyclic [XG00]. -d [EZ08, CM06, CFJ03a, CFJ03b, Mic02]. -dimensional [KNWN09]. -FEM [BS06b, MBS00]. -harmonic [BP08]. -iteration [VVM09]. -Laplacian [ABH05, EL05, Kim05a, LY01]. -LU [GKB09]. -matrices [Bou07]. -Matrix [GHK02, BH03b]. -method [AG01]. -minimization [GP08]. -node [ZK00]. -semi-norms [AdST05]. -sided [FG02]. -spline [AdST07]. -splines [AdST05]. -Stability [LT01]. -stable [Ber06]. -tetrahedra [KP08]. -Toeplitz [dFP05]. -type [CLM03]. -version [AG00b, Bab00, BH05a, BH07, GH04, Heu98, Heu01a, Heu01b, KS07, TS00].

2 [AD01, CCM04].

389 [TW01]. 396 [Heu01b]. 3D [GHS00a, KK00].

79 [Heu01b].

a-posteriori [BGS02]. ABS [EMAS01]. Absorbing [Sze06, LMV02]. abstract [BM00]. acceleration [Mat01]. Accuracy [HJL09, AGS09, CX08, DKM09, LYZ01, McS09a, NTW01]. Accurate [ARPR04, DK04, Pas00, Mar01, MM07, zSlZ04]. acoustic [CWRR02, DGS07, HR04a, LS08, NW05]. action [RS04]. actions [CHL08]. active [ZLW04]. adapted [CX08]. Adaption [PW05b]. Adaptive [BV00, BDD04, CCF08, DRSZ08, HKS07, ZCW06, BDS07, BC08, BM08, BD04, BC00, BCH07, BQR05, CMS01, CCS05, CH06, CFA06, CM00, DSX00, DXZ08, Fae02, Grü04, HLW03, MPRS01, MSTZ03a, MSTZ03b, Moo03, PM08, Vee02, Wil03]. Adaptivity [Moo03]. added [BRS00]. Additive [BDR03, Heu01a, RXH05, TS00, AG00b, Bre03, BW05, Nab03]. adiabatic [JL03]. adjoint [Hag00]. admissibility [Erv09]. advection [Ber01, BMS00, San01, San04]. advection-diffusion [BMS00, San01, San04]. affine [Zha04]. affinely [GK01]. after [Can06]. aggregation [VBM01]. algebra [AB02, Boc01, DDH07, Boc01]. Algebraic [BFNS01, BCD05, CFM06, GLX06, KS02c, KMS04, Stö01, VBM01]. algebras [AD07, DFLZ03]. algorithm [ACC^{+}02, AD01, AKK^{+}08, Ari04, Bey00, BHV03, CMS01, CFA06, CG04, Dry05, FP09, GL05, GLX06, Hor04, Kim05b, Kre06, MNS08, MS00b, Nie02, PCM02, Per05, RS07, Tim00, ZLW04]. Algorithms [BB06b, BGP05, BH03c, CC04, CK02, CDLP07, EMAS01, HXZL08, HZ02, JJL02, TS00]. Allen [FP03a]. Allometry [Fou09]. allow [LP08]. allows [PW05b]. alloy [BB01]. Almost [LPSV03, Chu03]. alternating [BCGJ07]. alternative [Peñ03]. American [AW07, IT08, LHJB07, dFL04]. analogues [Gro09]. analyses [CP01b]. Analysis [Bak02, Ber01, Ber03, CR05, CB00, DK01, DI01, DV08b, ER09, HHW00, HSZ06, LN04a, OR06, PS09a, San04, Sch09, Sid05, WL02, WLJ04, ABC03, ACC^{+}02, AD01, AMtTB03, BNT05, BLS05a, BC07a, BCD04, Bar05, BGHNR03, BRS05, Ber02, BCP07, BN01, BCH07, BMS00, BJ00, Cap02, CP01a, CS04, CH06, DF07, DHNL^{+}00, ESW02, FP03a, FP04, FNP07, FGSn02, Gal02, GLRvdE05, HCC09, HSS08b, HP01, HL04b, HMMC04, Joh04, KL00a, KR09, KK00, Küg05, KW08b, KW08a, LLP07, LM00, MN06, McS09a, MR03, Neg08, NN06, NCV06, PC09, Reu02, RAVS07, SPSF00, SBL06, ST05, Tsu03, WO00, dVLM09, CFJ03b]. Analytic [HSZ06, NR03, BGM01, DDP07, SS00, de 05]. analyticity [TW00, TW01]. angle [Str08]. Anisotropic [FP03b, Ang00, ANS01, BGN08, BCD05, DE06, FP01, KV00, Kun00]. annuities [CF08]. any [DE06, MX06, WcSX07]. Appending [DK01]. application [BBD06, BLV06, CWRR02, Fae02, FG02, LS08, OPR06, SL02, TJ07, PMC01]. Applications [Gei07, Moo03, WR05, AXY02, AdST07, Ber09, BGCP05, BSS04, Erv09, GHK02, GLRS09, LS09, MSS02, Med01, NCV06, SZ09, Yan02, GPC05, BSS05]. applied [Gri05c, HWW02, Koc05, Mas02, San01]. Approach [CFM06, AC01, BFS01, BGS02, BR04, BMVX07, BJ03, CK07, CLRM01, CIS06, CS07, Che01, GZ01, HR04b, PY02]. approximability [BBO04]. approximants [BH03b, BR02, Mat01]. Approximate [CJMR04, HKT08, OWW04, WR05, BKW07, JKC09, Jer08, Küt03, LT01, MP02, NTW01, RS04]. approximately [ZLW04]. Approximating [AM08, BH04, KM02]. Approximation [Beb00, BLM05, CZ03, FHV03, Hac07, Mos01, AG00b, AP01, ADM04, Bab00, BC07a, BH03a, BB01, BLN04a, BB04, BB08, BRM^{+}04, BD04, BG05, BC00, BMV09, BC03b, CR00, CDRG^{+}07, CB00, DMC00, DGH09, DR04a, DR03, EL05, FP03a, FH07, Gal02, GW05, GT07, GHK02, GM06, GHL^{+}05, GT04, HPPS02, HR04a, JKN09, Joh00, KW06, KP01a, LSW02, Liu00, LY01, LFLPS05, MS00a, MSTZ03a, MSTZ03b, Neg03, Pam01, Put02, Sti05, SHS08, Wal00, hW02, Yse07, Zha03, Zha04]. Approximations [MSS02, ABH05, Bak02, CM08, DW00, DH06, EK00, FvOP05, FNP07, GGM02, GGM09, GPW03, GJ02, HR00, JU05, LRW06, LM08, LPSV03, NCV06, Ohl01, TS03, Usc08, ZR05]. arbitrary [Chr07, Hes06, KMS04, MT05, dlCYGV07]. argument [VC06]. arising [BLN04b, FNP07, GJ02, Liu00, dSB08]. Arithmetic [BB06b]. Armijo [ZZL06]. Armijo-type [ZZL06]. Aronszajn [Fae02]. artificial [HHW00, TZ03]. aspects [LS08]. assemblies [AC01]. associated [JKC09, KL00b, Noë04]. assumption [DN02]. Asymptotic [BD01, CC04, GT04, HL04b, LM05, Max07, WW03, BC07b, Cao06a, DGS07, GMZ02, Pam01, WW02]. Asymptotic-preserving [GT04]. Asymptotically [Koc05, Ova06]. Atkinson [Sid05]. atmosphere [Med01]. atoms [DY09]. attraction [BD02]. augmented [BPW05, Che01]. automata [CC00]. automatic [PMC01]. Auxiliary [HWZ06]. average [KW08b]. Averaging [BC04, Huh07, CCC^{+}05, FV06]. axiomatic [CIS06]. axis [Zhe06]. axisymmetric [AP05, BBD06, Lac00, CJK^{+}06].

Babuska [Tsu03, XZ03]. backprojection [RS07]. Backward [MR03, Mos01]. balance [BAL07, Hac07, KW08b, KW08a, STWW07]. balanced [GT04]. Balancing [GPW03, LLP07, dVLP06]. Banach [CM08]. band [NCV06]. banded [SC07]. barrier [FO09]. bars [MT03]. Barycentric [FH07, Gau01]. Barzilai [DF05]. based [AM01, AS02, BFLW09, BKW07, CMPS04, CS06, DV00, EW03, GHM09, GKB09, Kan01, LRW06, MM09, Max07, Moo01, MS00b, Ova06, OM01, Pro00, RT06, RS02, VBM01, Voh08, Yun03, dV08a]. Bases [HSZ06, LSW05, LS01a, Peñ02, Peñ06, SS05]. basis [Joh04, TGSS08]. BDF [DV08b]. BDMS [LV04a]. beam [Cop08]. behavior [HL04b]. Bellman [Grü04, JKC09]. BEM [EW03, EH06, GM05, HPPS02]. BEM-discretizations [EW03]. Bénard [KNWN09]. bending [Lyl00, PS00, dVNS07]. benefit [CF08]. Bergman [BR02]. Bermúdez [PCM02]. Bermúdez-Moreno [PCM02]. Bernstein [AKR09, Not06]. Berrut [Gau01]. Besov [ABH05]. Bessel [Pas00]. best [CNY02, DQQ01, QQ03]. between [Lyl00, Nab03]. BGK [Bou03]. bidiagonal [BOL02]. bidimensional [Gal02]. bifurcating [KNWN09]. biharmonic [AD01]. bimaterial [Liu00]. Bingham [Zha03]. binomial [LHJB07]. Biorthogonal [HPPS02, SS00]. Bivariate [BMVX07, vGLS02]. Black [AW07]. Block [NCV06, BN03, Cao01a, CG04, Kre09, XG00]. block-diagonal [CG04]. Blockwise [XG00]. blow [BQR05]. blow-up [BQR05]. Blowing [FG02]. Boltzmann [JY09, MP05, PTV03]. Borwein [DF05]. Boson [MP05]. bound [AdST07, BB01, Hes06]. boundaries [Zha07]. Boundary [BCS02, BHvPS03, CM06, Sze06, AG00b, AM03, AM02, AKK^{+}08, ADM04, Beb00, BH05a, BH07, CR05, Cao06a, CM00, CS00b, CS03, DK01, DGS07, DHN07, ES08, Fae02, GGLV01, GHS00a, GHS00b, GS02, GNP08, GH04, HKS07, HHW00, Heu98, Heu01a, Heu01b, HS09a, HS05, KK05, Koc05, KS02c, Kwe00, LW00, Lin09, Mar01, Mic02, Moo01, MS00b, Nov01, SA07, Ste00, Ste03a, SU09, Stö01, ST05, TW00, TW01, TS03, TS00, BC03a]. bounded [JY09, Vov02]. Bounds [PS02a, Bab00, BH03a, BH03c, Che07, GH08, Gri05c, HWW02, McS02, PS00, XY09, de 05, vGLS02]. Boussinesq [WLJ04]. box [DF05]. box-constrained [DF05]. bratwurst [KL00b]. breakage [KW08b, KW08a]. Brezzi [XZ03]. Bubble [GGRG05]. bubbles [BMS00, San01]. Bulirsch [Fis02]. bundle [Sim00]. BVM [SC07]. BVMs [AT03].

C [BW05]. Cahn [FP03a, FP04, Fur01, GW05]. calculation [Sch09, Weg02]. can [CDLP07]. canonical [OM01]. capacity [JV01, zSlZ04]. capillary [Bän01, Weg02]. capillary-gravity [Weg02]. cardinal [BD01, Gau01]. Carlo [CDLP07]. Cartesian [LLW03, VK09]. cascade [JJL02]. cascadic [Ste02b, Tim00]. Case [CFM06, AJMN02, ACP02, BNT05, CNY02, ER09, EPS01, Fae02, GR04, GP08, Hes06, LN04a]. Cassini [Peñ03]. Cauchy [BM00, KL03]. cell [HCC09, KW08b]. cement [AJMN02]. centered [CK02, HCC09]. central [CK08, KP01b]. central-upwind [CK08]. certain [FL06, Küg05]. CG [AK04]. chains [KNS08]. characteristic [BMV09, DMC00, RT02]. characteristic-Galerkin [DMC00]. characteristics [Mah09]. Characterizing [GS08]. Chebyshev [AM01, AKR09, DGV02, Maz08a, dFGAN00]. chemically [Pro00]. chemotaxis [CK08, Fil06]. choice [KS02a, PMC01, VVVM06b]. Christoffel [OST07]. circle [DGV02, SL02, dlCYGV07]. circles [Nie02, Peñ03]. Circulant [CC00, CNY02]. clarifier [BKRT04]. clarifier-thickener [BKRT04]. class [AG00a, AMN09, BS06a, Ber01, But04, CP06, Cao06b, CWRR02, CLM05, EMAS01, MM09, WcSX07, dV04]. classes [AT03, Bey00, FL06]. classical [GLRvdE05, HJL09, SBL06]. Clément [BBD06]. Clenshaw [WT07]. close [JL03]. closed [GS02]. closure [BBC^{+}05]. Clough [AS02, LS01a]. cluster [Sch09]. clustering [HKS07]. coalescence [Hac07]. coefficient [BM00]. coefficients [BH03b, BV00, Dry05, EW02, JKN09, Lin09, RXH05, TT05, WP00]. coercive [AG00a]. collision [PTV03]. collisions [Mau06]. Collocation [Stö01, CS00b, CS03, ED02, Koc05, KS02c, KMS04, Mar01, MFR03, SA07]. Combined [BH05b, EHV06]. Comments [Pyt02]. commutative [AT01]. Commuting [DST05]. compact [Lin08]. compactness [Bof00]. companion [BGP05]. Comparison [Son02]. Comparisons [Nab03]. Complement [CJK^{+}05, CJK^{+}06]. complementarity [ACP02, Kan01, MNS08, ZLW04]. complementary [HMS03]. Complex [DH09, HS06, Put02]. Complexity [MSW06, DRSZ08, DXZ08]. compliance [FGSn02]. complicated [CS04, RSS06]. component [BB01]. Componentwise [CP01b]. Composite [mWS05, RSS06]. Composition [MS00b]. compositions [CJMR04]. compound [AB02]. compressibility [KL00a]. compressible [FMS07, GMZ02, Kwe00, Yan02, Yse01]. Computation [BLV06, GHNP03, HSZ06, PR09, SL02, BDS07, BP04, BD04, DRSZ08, Mat01, MC04, Pas00, PGL07]. Computational [Dzi08, BB00, FGSn02]. computations [BGM01, DXZ08, Fou09]. Computing [DHN07, Mul01, BL09, CLM03, CLM07, DV05b, ED02, GST04, Li01, Pro00, RS02, VVM04]. concentrated [JV01, zSlZ04]. concentration [BB01]. concepts [DDP07]. Condition [CW03, BC07a, Bec00, BGL07, DK01, GS08, KNS08, Kwe00, Leb02]. conditioned [BRZRS03]. conditioning [BLO07]. Conditions [Sze06, AJMN02, BCD04, BBCM00, BLV06, CM00, DK01, DHN07, Erv09, EH08, HHW00, HR00, KK05, LN04b, MS00b, SS08, TMO04, TW00, ZOS00, TW01]. conducting [KM02]. confluent [Mul01]. conformal [KL00b, MSS02]. conforming [AJMN02, BCH07, Lac00, MT05, WcSX07]. congruence [Bey00]. conjugate [Ari04, Dai01, EN00, LP08, NE00, ZZL06]. connecting [CKP07, Pam01]. connection [Lyl00]. conquer [CG04]. Conservation [CHL08, ADM04, Ber05, Bou01, CHC01, CL01, DGMM01, Des01, GR04, HKR01, Hac07, HJL09, HWW02, KMR09, KP01b, Küt03, LS01b, LWW01, OWW04, PW05a, Vov02]. Conservative [CFJ03a, CFJ03b, Ber01, Fur01, Mah09, MP05, PMC01, Voh08]. Conserved [Can06, CLL08]. consistency [LSW02]. Consistent [MKV04, Leb02]. constant [BM00, PCM02, Pot04]. constants [Fou09, Wal00]. Constrained [Hor04, DF05, Hai03, HR04b, HTY08, Kor02, MS00b, SZ09]. constraint [BP08, Cao06b, CLRM01, Tai03]. constraints [DGH09, RT06, ZLW04]. Constructing [Xu00]. Construction [Chu06, CS00a, CFJ03a]. Contact [FHV03, Che01, Cop08, EW03, FGSn02, RAVS07, SŻ05]. continuation [NR03]. Continuous [BF07, BKRT04, EZ08, FG02]. control [BC04, BV07, BLV06, BY05, CCS05, Car05, CH07, CGJ09, CF08, DH04, DGH09, GLQ01, Hag00, HR04b, HTY08, LY03, MST06a, NSV03, NAE07, SW00]. Controllability [CZ03, CM06, Mic02]. controlled [GK01, JKC09]. convection [Ang00, BV07, BGCP05, DV08b, Hol08, LS01c, Mah09, Ohl01, Reu02, RT02, TT05, ZR05]. convection-diffusion [BV07, BGCP05, Hol08, LS01c, Mah09, Ohl01, Reu02, RT02, ZR05]. convection-dominated [Ang00]. Convergence [ACC^{+}02, AW07, AMtTB03, BC07a, BP08, Ber05, BDD04, BC09, Bou01, BCH07, BLN04b, Cao01a, Cap02, CH06, DXZ08, DR03, DQQ01, EGHM02, FMDH03, FP09, GJ02, HKR01, HR00, JJL02, JY09, KMR09, Küg05, KW08b, KW08a, LRW06, LS01b, LX09, LWW01, MSTZ03a, PM08, PGL07, Reu02, ST05, VBM01, VV03, Vov02, Zha07, dVLM09, BNZ05, BCR04, BCD05, BKRT04, BF07, CDLP07, hDxY99, EN00, FvOP05, GGJ03, GH04, JKN09, Jer08, JV01, Kim05b, KW06, Kwe00, LR09, LZ09, LHJB07, Lin08, LS09, MT01, Mat00, NE00, PCM02, PT06, Pyt02, Rie01, SW03, SS05, Tai03, VVVM06a, ZOS00, ZZL06]. Convergent [PS09b, Vee02, ACMR06, AK04, BC08, BM08, BCGJ07, GR00, MT03, Obe04, PK05, ZLW04, Boc01]. convex [AM08, AP01, CR00, DW00, DQQ01, GLRS09, Küt03, QQ03, Usc08]. convexity [CLRM01]. convolution [CCP07, Hac08, LS08, PSN04, SW00]. convolutional [CS00b]. Coordinates [Gal03, KMKZ09, OM01]. correct [Koc05]. correction [AMtTB03, ZZ03]. corrector [CC05]. corrosion [AKK^{+}08, CHB08]. cosine [FO09]. Cotes [WS08]. Couplage [BBP01]. coupled [BRS00, BRM^{+}04, DI01, DHNL^{+}00, Sch09]. Coupling [EH06, GM05, GLQ01, GR04, HPPS02]. covolume [CK02]. crime [Sim00]. crimes [KK00]. criteria [ALW05]. criterion [Ari04]. Critique [Mül03]. cross [BB04, BG05, GGJ03, Yse07]. cross-diffusion [BB04, GGJ03]. Crouzeix [ANS01, HS09a, RXH05]. crystalline [BP04]. crystals [BCG06]. cubature [DST05, Hes06, Pet03, Xu00]. curl [AFW00, BB08, BCLS08, BL05]. curl-curl [BCLS08]. Current [ZCW06, BLS05a, BRS05, RVH09]. Curtis [WT07]. curvature [DD02, DH06, FP03a, FNP07, Obe04]. curve [HMS03]. curved [HR04a]. curves [CS03]. curvilinear [CS00a]. cusps [BNZ05]. cut [PTV03]. cut-off [PTV03]. cycle [AD07]. cyclic [XG00]. cylinder [GST04]. cylindrical [Zha03, dV04].

d [EZ08, CM06, CFJ03a, CFJ03b, Deu09, FMS07, KS07, LSW05, Mic02, RS04]. D-Doppler [RS04]. DAEs [CCM04]. Dai [Pyt02]. damped [TZ03]. Darcy [ABC03, GW08]. data [Bak02, CIS06, DN02]. DDM [BBS07]. decay [EZ08, TZ03]. Decomposition [BB06a, Ber03, BH03c, Buf02, BW04, CM00, DT07, FML00, GGM02, GK02, GLS07, GKB09, KV01, Lac00, LW00, Nab03, Str08, Tai03, dSB08]. decompositions [Chr07, HN00]. Deconvolution [CS07]. decoupling [Chu03]. defect [AMtTB03]. defined [FL06]. defining [AT03]. definite [Bec00, BGCP05, Cao01a, dVLP06]. definitions [BC03b]. Degenerate [CE01, AM06, BLN04a, BLN04b, EK00, EGHM02, EHV06, PY02, RPK08]. degree [BFLW09, MCS09b, Pet03]. Delay [Hua09, BM00, ED02, Mas00, Mas02]. Delay-dependent [Hua09]. delayed [WP06]. delta [ABR06]. dependent [AGS09, BB01, Bar05, BCP07, BCGJ07, BF07, HP01, Hua09, MW04, MW06, OPR06, TT05, Zha03, ZCW06]. derivative [Küg05]. derivatives [SŻ05, Yse04]. derived [CM06, EFG06, NCV06]. design [BC08, Maz08b]. detecting [AKK^{+}08]. DGFE [DV08b]. DGFEM [SW03]. diagonal [AM03, BN03, CG04, KPS03, Put02, VVVM06a, VVVM06b]. diagonal-plus-semiseparable [VVVM06b]. diagonalizable [Che07]. diagonally [AXY02, Bou07, DK04]. Difference [CZ03, AT01, BLS^{+}05b, BD02, BKRT04, Da08, Fur01, JV01, Lin08, LSY09, Obe04, zSlZ04, WP00, WL02, WW02, WW03, WLJ04, WP06, dVLM09, dV08a]. differentiability [EH08]. Differential [GP03, BM00, BMV09, BLN04b, BY05, CYY07, Chr07, DR04a, ED02, GM06, GPHA04, Gri05c, Gri05b, GS00, Hei09, HK05, HMY07, JKN09, KS02c, KMS04, Mas00, Mas02, MSW06, MST06b, MSTZ03a, MSTZ03b, SW08, Stö01]. differential-algebraic [KS02c, KMS04, Stö01]. Diffraction [GHNP03, BBP01]. diffusion [Ang00, AMtTB03, BB04, BV07, BGCP05, BMS00, BCGJ07, DR03, DM04, DV08b, DE06, GGJ03, HS05, Hol08, Jer08, Lin09, LS01c, Lin08, Mah09, MR01, Ohl01, Reu02, RT02, San01, San04, Ste05, WP00, WP06, ZR05, dFL04]. diffusions [JKC09]. diffusive [CL01]. diffusive-dispersive [CL01]. digital [CDLP07]. Dimension [HO08, Chr07, Hes06, Maz08a, Moo01, Pet03]. Dimensional [LM05, Bou01, BH03c, BCLS08, CR05, ER09, EFG06, Fae02, Fou09, GHM09, GP08, HKR01, KNWN09, LN04a, ST05, Sze06]. dimensions [AG00b, Bab00, BGN08, GH04, Hac08, HHW00, Heu98, Heu01b, MX06, Tos00, TS03, WcSX07]. diminishing [PW05a]. Diophantine [EMAS01]. Dirac [ABR06]. direct [BHV03]. direction [BCGJ07]. Dirichlet [TW01, BCD04, CC04, EFG06, GPC05, NR03, RSS06, SU09, TW00, BC03a]. disc [AP05, Joh00]. Discontinuous [BL05, Lin09, RSWW03, CGJ09, Dry05, GK03a, GK03b, Kim05b, LN04a, LPSV03, RXH05, ZR05, GNP08]. discrepancy [JT09]. Discrete [DSS02, EK00, GHS00a, Hip01, JU05, KK05, AdST05, Bak02, BGCP05, Bof00, CM06, CL01, DD04, DD02, ESW02, EZ08, FO09, Gei07, GS02, Grü04, KP01b, LWW01, Mic02, RS07, Vec01, Zha07, ZZ09]. discretization [AG00a, ABC06, Bän01, BP08, BV04, BBD06, BBCM00, BF07, Can06, Cas01, Dry05, DT07, GGJ03, GW08, Grü04, HL04b, HLS05, Hol08, MPRS01, MST06b, RPK08, RG04, SKR04, TZ03]. discretizations [ABC03, CD04, CHL08, EW03, HJL09, Mar02, PK05, PS09b]. discretized [AGS09, MST06a]. disk [Boc01, Boc01]. Disk-Algebra [Boc01]. dislocation [ACMR06]. dispersive [CL01]. dissipative [MN06]. distance [PS02a]. distributed [DH04]. distribution [AC01, Bak02]. distribution-valued [Bak02]. disturbance [Chu03]. div [AFW00, BB08, BCLS08, BL05]. div-curl [BB08]. divergence [CDRG^{+}07, DR04b]. divergence-free [DR04b]. divide [CG04]. divide-and-conquer [CG04]. Domain [BB06a, CZ03, GGM02, GLS07, GKB09, BH03a, Ber03, BH03c, Buf02, CK07, CM00, DT07, FML00, GK02, GGLV01, LW00, Nab03, BC03a]. Domains [BHvPS03, Olv06, BNZ05, BCS02, BCD05, BCJ09, CCSS09, CK07, CC04, Cao06a, CS04, CJK^{+}05, CJK^{+}06, CD02, GLRS09, HSS08b, HO09, HR04a, JKN09, JY09, RSS06, Vov02]. dominant [AXY02, Bou07, DK04]. dominate [KV00]. dominated [Ang00]. Doppler [RS04]. double [TMO04]. doubling [GLX06]. doubly [EK00]. Drazin [SK03]. Drazin-inverse [SK03]. DtN [NN06]. Dual [LSW05, CCS05, LP09a, MT01, QQ03, Li05]. dual-mixed [CCS05]. dual-primal [MT01]. Duality [PMC01, GPC05]. d’un [BMV03]. DW [HS08]. dynamical [dSB08]. Dynamics [Gal03, ACMR06, CCC^{+}05, JL03, KQW04, SHS08].

early [FO09]. early-exercise [FO09]. Eddy [ZCW06, BLS05a, BRS05, RVH09]. Edge [KV00, Bof00, BCG06, BCD05, KP08, TS03]. effect [SS05]. Effective [KO00]. Effects [BC03a]. efficiency [GS07]. Efficient [Hac08, HSZ06, NTW01, Xia07, PW05b]. eigendecomposition [CG04]. eigenfunctions [CC04]. eigensystems [Ste01b]. eigenvalue [BGL07, BGP05, BCJ09, DRSZ08, DXZ08, Kre09, LN04b, Noë04, SU09]. eigenvalues [BS06a, CC04, Che07]. eigenvectors [dSB08]. Elastic [FHV03, BB08, Che01, FGSn02, Liu00]. elastic-plastic [Liu00]. Elastic-Viscoplastic [FHV03, FGSn02]. elasticity [AW02, BGS02, CD04, CS00a, GPW03, LWC09, LV04b, RSWW03, Ste03a, Yan02, dVLP06]. elastoacoustic [BGHNR03]. elastoacoustics [ARPR04]. electric [BRS05, BC03b, RVH09]. electromagnetic [BCJ09, KM02]. électromagnétique [BBP01]. electronic [Sch09, Yse04, Yse05, Yse07]. electrostatic [AKK^{+}08]. Element [BB06a, BDD04, BDR03, BHvPS03, LSW05, ZCW06, AG00b, AM03, ARPR04, AD01, Ang00, Ari04, ALW05, Bab00, BBO07, BNZ05, Bän01, BH03a, BV02, BB01, BLN04a, BB04, BGS02, BC04, BCD04, BC08, BP08, Beb00, BV04, BV07, BM08, BBD06, BBCM00, BL05, BRS00, BGHNR03, BV00, BH05a, BH07, BG07, BG09, BC00, BCR04, BCH07, BCLS08, BF01, BY05, BCS02, BF07, BGG04, Cao06a, CYY07, CBJ02, CD04, CCS05, Car05, CH06, CS06, CH07, Cas01, CM06, CX08, Chr07, CB00, DXZ08, DGH09, DI01, Deu09, DHNL^{+}00, DH06, EL05, EW02, ES08, Erv09, EHV06, Fae02, FP04, FvOP05, FNP07, GGM02, GT07, Gei07, Ger00, GHL^{+}05, GHS00a, GH04, GLRS09, HKS07, HLPS05, HP01, HMMC04, HXZL08, HO09, HR04a, Heu98, Heu01a, Heu01b]. element [HS05, Hol08, HS09b, Jer08, JU05, KK05, KL00a, KM02, KK00, KS07, KV00, Kun00, Lac00, LRW06, LM08, Ley04, LLW03, LYZ01, Lin09, Liu00, LV04b, Mar02, McS02, MX06, Moo03, Neg03, Neg08, RPK08, RXH05, RSS06, RSWW03, RT02, SA07, SKR04, Sch04, Ste00, Ste03a, SU09, TT05, TW00, TW01, TS03, TS00, Tsu03, hW02, WcSX07, Wil03, XC01, Yan02, YW06, ZK00, ZZ03, dVNS07]. elements [AM06, AS02, ANS01, AW02, BRM^{+}04, Bof00, BCG06, BCD05, BCJ09, CP01a, CDRG^{+}07, CS00a, CD02, FMDH03, GHL^{+}05, GH03, GHS00b, GGRG05, HS09a, KS01, LS01a, LV04a, Lyl00, McS09a, MT05, Osw08, PS00, ST03, Ste02b, Vee02, Zha04, BBP01]. elevation [Maz08a]. ellipsoids [MKV04]. Elliptic [BDR03, CZ03, ABR06, AK04, BNT05, BH03b, BV00, BG07, BG09, CC04, Cao06a, CYY07, Cap02, CS04, CDRG^{+}07, CN00, CK02, CM00, DGH09, Dry05, FP03b, GHL^{+}05, GNP08, HLPS05, HSS08b, HO09, Hor04, JU05, KK05, KK00, KS07, KL03, Küg05, LW00, LN04a, LM08, LPSV03, Liu00, LS09, Lui02, MX06, NSV03, RXH05, ST03, Tim00]. endpoint [Sid04]. Energy [CLL08, CHL08, EZ08, Sti05, vGLS02]. Energy-conserved [CLL08]. Enhanced [CS05, BCR04, NN06]. enriched [FMDH03]. entropic [DM04, MP05]. Entropy [Bou03, Gal03, Ber05, DR04a]. Entrywise [AXY02, XY09]. envelopes [DW00, LP01]. Epi [PK05]. Epi-convergent [PK05]. equal [But04]. equal-weight [But04]. Equation [CFJ03a, CFJ03b, ACMR06, AW07, BLN04a, BV07, BGCP05, BN01, BC03b, Cao06a, CMS01, CMPS04, CM06, CZ03, DD04, DF07, FP03a, FP04, Fur01, Gal02, GW05, GR00, Grü04, GLX06, JV01, Kan01, Lac00, Liu00, MP05, MST06b, MM07, Mic02, MS00a, Neg08, PTV03, Per05, SKR04, Ste05, zSlZ04, TZ03, WW02, Yse04, Yse05, Zhe06]. equation-based [Kan01]. Equations [BH05b, Gal03, MSW06, Sze06, AGS09, ABC03, AM03, AKK03, ACC^{+}02, Ang00, AMtTB03, AG01, ABC06, Bän01, BGN08, Bar05, BM00, BV00, BGP05, BMV09, BY05, BCS02, BD02, BF07, CCP07, Can06, CC04, CYY07, CCS05, CDRG^{+}07, Cas01, CC05, CWRR02, CLL08, Chu06, CB00, CHL08, CS00b, CD02, CS03, Da08, DMC00, DH04, DR04a, DY09, ESW02, ED02, EK00, EGHM02, EFG06, FMS07, GGM02, Ger00, GM06, GGLV01, Gri05c, Gri05b, GP08, GGRG05, GS00, HMS03, HO08, HPS03, Hei09, HK05, HMY07, HS08, HPSS05, JKC09, JKN09, JU05, KL00a, KW08b, KW08a, KMS04, Ley04, LYC03, Li05, LY03, LFLPS05, MT03, Mas00, Mas02, MFR03, Med01, MR01, MX06, MSTZ03a, MSTZ03b, Moo01, Mos01, NSSV06, Ohl01, OR02, OWW04, OPR06, PMC01, PY02]. equations [RPK08, RG04, Reu02, STWW07, SW08, SW00, TW00, TW01, TJ07, Tos00, TGSS08, Vec01, WP00, WL02, WW03, WLJ04, Wil03, ZZ09]. equilibrated [Max07]. equilibria [GP03, GPHA04]. equivalence [Gro09]. equivalences [BSS04, BSS05]. equivalent [AK04]. Erratum [BSS05, GK03a, Heu01b, MW06, TW01]. Error [AdST05, BH03a, BH04, Buf02, FP04, FNP07, Grü04, HWW02, JKC09, Kim05a, MT03, MV07, NN06, PC09, RPK08, TT05, WO00, Zha03, dSB08, de 05, vGLS02, AM06, AMN09, ACC^{+}02, ABH05, ABR06, AG01, BFS01, BB01, BGS02, BC04, BCD04, Bar05, BV04, Ber02, Ber09, BD01, BG09, BCR04, Bra05, BMS00, BJ03, CMS01, CBJ02, CS04, CD04, CMPS04, CCS05, Car05, CS06, CH07, CGJ09, CHC01, CN00, CE01, CM08, DH04, Deu09, DV05b, EL05, EW03, FV06, FP01, FP03b, GHM09, GLRvdE05, Gri05c, Hes06, HN05, Hol08, HS09b, Joh04, Koc05, KV00, Kun00, Küt03, LM08, Ley04, LT01, LY01, LY03, LV04b, MN06, Max07, McS02, Moo01, Moo03, MR03, NSV03, NSSV06, Noë04, Not06, Nov01, Ohl01, Ova06, PS00, SKR04]. error [SW08, SW00, SW01, SBL06, Tsu03, Voh08, Wal00, ZK00, dV04, dVNS07, dV08a]. error-based [Moo01]. errors [Bab00, SS00]. Essentially [TJ07, XY09]. estimate [BGS02, BG09, Buf02, CMS01, CD04, CMPS04, CS06, CHC01, Hol08, HS09b, MV07, Ohl01]. Estimates [DW00, AM06, AMN09, ACC^{+}02, ABH05, ABR06, AdST05, AG01, Bab00, BFS01, BH04, CBJ02, CN00, CE01, CM08, DRSZ08, DH04, Deu09, EL05, FP01, FP03b, GLRS09, JKC09, Kim05a, KV00, Küt03, LM08, Ley04, LT01, LY01, LY03, Moo03, NSSV06, PW03, RPK08, SKR04, SW01, Ste02a, TT05, Voh08, ZK00, Zha03, dV04, dVNS07, dSB08]. Estimating [SS00]. Estimation [BD02, BV04, BJ03, Grü04, Koc05, Moo01, Nov01, SW08]. Estimations [BMV03, HN05]. estimator [EW03, GHM09, Kun00, Max07, Noë04, San01, dV08a]. estimators [BCR04, Bra05, FV06, LV04b, Ova06]. Euler [FMS07, Sid04]. Eulerian [Gal03]. European [AW07]. evaluation [DSX00]. Evans [HSZ06]. evolution [AKK03, AG01, BFS01, BGN08, HMS03, HO08, MN06]. exact [Ova06]. exactness [Max07, MCS09b]. exercise [FO09]. Existence [BH03b, SHS08, BFLW09]. expansion [Cao06a, GMZ02, WW02]. Expansions [LM05, BDS07, BJ00, CC04, DSX00, EvP04, FO09, Sid04, WW03]. Expected [CW03]. experiments [FGSn02]. Explicit [dFP05, CP06, CC05, TJ07, Tsu03, VV03]. Exponential [SW03, CP06, GHK02, Gri05c, HMY07, Huh07, TZ03]. exponentials [XY09]. exponents [BL09, DV05b]. Extended [Maz08a, GT07, AKR09]. extension [AdST07, Tsu03]. extensions [DST05]. exterior [AB02, BG01, BGG04, DI01, Deu09, ES08, FML00, GM05, HHW00, NN06, Nov01]. extra [AB03]. Extrapolation [Wen01]. extremal [Sti05].

Faber [KL00b]. Factorization [BGM01, CP01b, PR01, Yun03]. family [LN04a]. Fast [BDS07, BD04, BGP05, DDH07, DDHK07, LFLPS05, MP05, OR02, PSN04, CWRR02, DSX00, MNS08]. FDTD [CLL08]. FE [NN06]. feasible [Kan01]. Fejér [WT07]. FEM [BC07a, Ber02, BS06b, CGJ09, EH06, GM05, HPPS02, MBS00]. FEM-BEM [HPPS02]. FETI [Bre03, Li05, PS09a]. few [Pet03]. fictitious [CK07, GGLV01]. Field [BH05b, BRS05, BC03b, Noë04]. fields [Ber03]. filling [FG02]. film [GR00]. Filon [Xia07]. Filon-type [Xia07]. filter [Mos01]. filtered [RS07]. filters [BCM06]. filtration [ST05]. financial [BLN04b]. finis [BBP01, BMV03]. finis-potentiels [BBP01]. Finite [Bän01, BLN04a, BB04, BB06a, BGHNR03, BDD04, BDR03, BY05, CYY07, CDRG^{+}07, CHC01, CZ03, DGH09, Des08, Deu09, DHNL^{+}00, DH06, GT07, GHL^{+}05, HR04a, KK00, LSW05, Liu00, QQ03, WP00, mWS05, ZCW06, dV08a, AT01, AJMN02, ABC03, ARPR04, AD01, ABH05, Ang00, AW07, ANS01, Ari04, ALW05, AW02, Bab00, BBO07, BNZ05, BH03a, BV02, BB01, BGS02, BC04, BCD04, BC08, BP08, BV04, BV07, BM08, BBD06, BL05, BMV03, BRS00, BV00, BRM^{+}04, BLS^{+}05b, Ber09, BG07, BG09, BCG06, BC00, BCR04, BCH07, BCLS08, BF01, BCJ09, BF07, BGG04, Cao06a, CP01a, CBJ02, CD04, CCS05, Car05, CH06, CS06, CH07, CM06, CWRR02, CX08, Chr07, CB00, CD02, DXZ08, DI01, DE06, EL05, EW02, Erv09, EGHM02, EHV06]. finite [EFG06, FP04, FvOP05, FNP07, FMDH03, Fil06, Fou09, Fur01, GGM02, Gei07, Ger00, GH03, GGRG05, GLRS09, HCC09, HLPS05, HP01, HMMC04, HXZL08, HO09, Hol08, Jer08, JV01, JU05, KK05, KMR09, KL00a, Kim05a, KM02, KS01, KS07, KV00, Kun00, Lac00, LRW06, LM08, Ley04, LLW03, LYZ01, Lin09, Lin08, LSY09, LV04b, MP02, MT05, MV07, MR01, Moo03, Neg03, Neg08, Nie02, Ohl01, Osw08, PS00, RPK08, RXH05, RSS06, RSWW03, RT02, SKR04, Sch04, ST03, Ste02b, TT05, TW00, TW01, TS03, Tsu03, Vee02, VV03, Voh08, Vov02, WL02, hW02, WLJ04, WP06, Wil03, WS08, Yan02, YW06, ZK00, ZX02, Zha04, ZZ03, dVLM09]. finite-dimensional [Fou09]. finite-element [BC00, Hol08]. finite-infinite [DI01]. finite-part [WS08]. finite-volume [BMV03, HCC09]. First [Ber09, BC07a, Pro00, VV03]. fit [Nie02]. fitted [AW07]. fitting [MKV04]. fixed [Chu03, KW08a, MFR03]. Fletcher [ZZL06]. flow [BP08, BBC^{+}05, Ber02, BCP07, CR05, CE01, CS00a, DD02, Deu09, Dzi08, FvOP05, FNP07, SW03, Yan02, YW06, Zha03]. flows [BC07a, FP03a, JY09, Pro00, VK09]. fluctuation [Bou01]. fluctuation-splitting [Bou01]. fluid [BB00, BRS00, DHNL^{+}00, GHM09, LM00, Zha03, dSB08]. fluid-solid [BRS00]. fluid-structure [LM00]. fluids [BRM^{+}04, Yse01]. flux [Ber05, Bou03, KW06, LSY09, Voh08]. flux-based [Voh08]. Fock [MT03]. fold [BGS02]. Forchheimer [GW08]. form [CDRG^{+}07, CLM07, DF07, PMC01, VVVM06a]. forms [Chr07]. formula [Ber09, BD01, BP01, BP03, MC04, Tsu03]. formulae [DST05, Gau01, MCS09b, Not06, Xu00, dlCYGV07, de 05]. formulas [DGV02, PGL07, SL02]. Formulation [ZCW06, AB03, ABC06, BG01, BGN08, BRS00, BGHNR03, CF08, KS02b, Kwe00, LRW06, LLW03, RVH09, SW01, WL02]. formulations [BRS05, CS00a, Med01]. Fortin [Bof00]. Fourier [CJK^{+}05, CJK^{+}06, FO09, Lac00]. Fourier-cosine [FO09]. fourth [Lin08, Mar01, MX06, STWW07, WLJ04]. fourth-order [Lin08, STWW07]. fourth/second [Mar01]. fractional [MM07, San04, Ste01a]. fracture [Neg03]. frame [CS07, CCF08]. frames [HSS08a]. framework [Ari04, BFLW09, BKW07]. free [Bän01, BMS00, CR05, DR04b, Küg05, ST05, VVVM06b, Zha07, San01]. French [BMV03]. frequency [BW04, DGS07, ER09, FML00, GJ02, Mos01]. Freudenthal [Bey00]. frictional [RAVS07]. frictionless [FGSn02]. Friedrichs [HKR01]. Froude [VK09]. full [AGS09, Can06]. Fully [GS02, Bak02, CL01, CM08, DD02, RS07]. Function [HSZ06, DV05a, DSS02, DHN07, Joh04, MC04, MC06, Mul01, NCV06, VC06]. function-vorticity [DSS02]. functional [BC00, BMV09, GT07, Ova06]. functionals [DSX00]. Functions [PP03, AM08, AdST07, BBO04, BS06b, BGM01, DGV02, GGL00, GST04, GLRS09, HS08, MM09, MSS02, Not06, Peñ02, Peñ06, PW05b, PR01, SS00, TGSS08, WP06, de 05]. fundamental [Vec01]. fundamentals [PS02b].

Galerkin [AMN09, CGJ09, DMC00, GK03a, GK03b, GHS00b, GNP08, GS00, HKS07, HMMC04, HLS05, KV01, LN04a, Lin09, Mah09, RSWW03, TS00, dFGAN00, dSB08]. Galerkin-characteristics [Mah09]. gamma [VC06]. Gas [Gal03, KQW04]. Gauss [Gau05, MC06, Not06, OST07, PP03]. Gaussian [BP01, BP03, DH09, MC04]. Gautschi [Gri05c, Gri05b]. Gautschi-type [Gri05c, Gri05b]. General [Hil06, BMS00, BJ03, CLM03, DSWW06, FNP07, GHS00a, LR09, dSB08]. Generalized [KMKZ09, MBS00, SPSF00, AKR09, BBO07, BPW05, BV02, BGL07, BAL07, BGP05, Che07, GPC05, KO00, OPR06, Pas00, Ste02a]. Generating [HS08, MM09, NCV06]. genuinely [KP01b]. geometric [BGN08, LS09, TS03]. geometrically [Nie02]. geometry [Lac00]. geostrophic [STWW07]. German [Boc01]. Gerschgorin [Peñ03]. Ginzburg [Bar05]. given [Pet03]. gives [DKM09]. Glimm [HWW02]. Global [hDxY99, Hai03, MS00a, Pyt02, ZZL06, JKN09]. globally [ZLW04]. glue [AJMN02]. GMRES [CLR02, LZ09]. GMWB [CF08]. GO [KPS03]. Godunov [Gal03]. Godunov-type [Gal03]. Good [DSWW06]. governed [BY05, LY03]. governing [AW07]. grad [BCLS08]. grad-div [BCLS08]. graded [Ste01b]. gradient [Ari04, Dai01, DGH09, EN00, FV06, LP08, NE00, Ova06, ZZL06]. Graeffe [MZ01]. Gram [GLRvdE05, SBL06]. Gratings [GHNP03]. gravity [Weg02]. grazing [PTV03]. Green [DHN07, GLRS09]. Gregory [DV00]. grid [AGS09, Bey00, BG07, Cap02, CWRR02, DE06, KS02a, LLW03, Ste02b, VK09, Yse05, Moo03]. Grid-to-Grid [Moo03]. grids [AJMN02, BLS^{+}05b, Buf02, GH03, HKR01, HCC09, KW06, LPSV03, SPSF00, WO00]. Griffith [Neg03]. group [AT01]. groups [KMKZ09, Wen01]. growth [BLO07]. guaranteed [CF08].

H [BNA07, ZCW06]. H- [ZCW06]. Hadamard [mWS05, WS08]. half [CWRR02, DHN07]. half-plane [CWRR02, DHN07]. Hamilton [ACMR06, Grü04, GP08, HS08, LYC03, LT01]. Hamiltonian [CCC^{+}05, Can06, CFM06, CLM07, Hai03, HJL09, Sun07]. Hancock [Ber06]. hand [CDRG^{+}07]. handle [BQR05]. Hankel [Bec00, BGL07]. haptotaxis [CK08]. harmonic [AP05, BP08, HPSS05]. harmonics [DSS02]. Hartree [MT03]. heat [BP08, BN01, HS05, JV01, zSlZ04, YW06]. Helmholtz [DI01, DHN07, ES08, FML00, GM05, GS02]. Herglotz [SL02]. Hermitian [BGCP05, BGP04, CNY02]. heteroclinic [BH04]. hétérogène [BBP01]. Heterogeneous [GLQ01, GPW03, HO09]. hexahedral [Zha04]. High [McS09a, BNZ05, BCGJ07, DGS07, DKM09, ER09, FML00, FH07, GJ02, Hua09, LYZ01, NTW01, dVLM09]. high-frequency [DGS07, ER09]. high-order [dVLM09]. Higher [GK01, LSW05, BGN08, JKN09, MN06]. Highly [Olv06, EW02]. Hilbert [Da08, EN05, Neu00, Yse04]. Hilliard [FP04, Fur01]. historical [Gau05]. HJB [ZZ09]. Hodge [Chr07, Hip01]. Hölder [GLRS09]. Holling [GT07]. holomorphic [MP02]. homogeneous [AP01]. homogenization [BJ00, HO09, MBS00]. homotopy [BFLW09]. Householder [Str09]. hp [GNP08, SW03, Moo03, TS03, SW08]. HP-Adaptivity [Moo03]. hp-DGFEM [SW03]. hp-Discontinuous [GNP08]. HP-finite [TS03]. hp-version [SW08]. HSS [BGCP05]. Hu [LRW06]. Hybrid [BG05, GHS00b, BL09, DGS07, EHV06, HMY07, Ste00]. hydrodynamic [BBC^{+}05]. hydrodynamics [BC07b]. Hyperbolic [BAL07, Ber05, BH04, DR04a, GR04, KP01b, LS01b, LMV02, Sti05, TJ07, Yse07]. hyperbolic-parabolic [TJ07]. hypercube [But04]. hypergeometric [Mul01]. hyperplanes [Nie02]. Hypersingular [LM05, AG00b, BH05a, CMPS04, Heu98, Heu01b]. hyperspheres [Nie02].

ideal [BMVX07, BKW07, BKRT04]. identification [BV04, KS02b, Küg05]. identity [CJMR04]. II [CNY02, CJK^{+}06, CFJ03b, Fae02, GT07, KW08b, NE00, VC01]. III [BRM^{+}04, CE01, EN00, RS04]. ill [BRZRS03, KS02a, Neu00, RT06]. ill-conditioned [BRZRS03]. ill-posed [KS02a, Neu00, RT06]. image [CCSS09, HMS03, MR01]. imaginary [VC06]. impedance [BHV03, DHN07]. Implementation [BC00, GHS00b, WY06]. Implicit [DKM09, AKK03, BP08, HMS03, MR01]. implies [DN02]. improve [CDLP07]. Improved [EK02, BB01, Joh00]. Improvements [GS07]. Improving [BNZ05]. impulse [CF08]. incomplete [VC06]. incompressible [BC07a, BH03a, CE01, Deu09, Ger00, GGLV01, GMZ02, LRW06, Li05, Sch04, Ste03a, WLJ04, dVLP06]. inconsistent [ZOS00]. increasing [Pet03]. indefinite [HPSS05]. independence [HTY08]. independent [CC05]. index [KMS04, CCM04]. Index-2 [CCM04]. inelastic [Mau06]. inequalities [GPC05, HXY08, HLW03, Kor02, KS07, Tai03]. inequality [Ber05]. inertial [NTW01, PR09]. inexact [HZ02, LR09, Rie01]. Inf [BBCM00, MT05, BC07a, GS08, Leb02]. Inf-sup [BBCM00, MT05, BC07a, GS08, Leb02]. infinite [DI01, GS00]. infinitely [AG01]. infinity [Deu09]. inflow [Kwe00]. inhomogeneities [BHV03]. Inhomogeneous [BCD04, GJ02, LFLPS05]. initial [AM02, Bak02]. inpainting [CCSS09]. Instability [Mas02, AD00]. instationary [DH04]. integer [EPS01]. integrable [CCC^{+}05]. Integral [BLM05, BH05b, GST04, BG05, BY05, BC03b, CMS01, CWRR02, CS00b, CS03, DGS07, GS02, Heu98, Heu01b, MFR03, Mül03, WS08]. Integrals [LM05, Olv06, mWS05, DH09, Hac07, JJL02, Sid04]. integrated [BS06b]. Integration [OM01, But04, MM09, MR03, PK05, PW05b, Sid05]. integrations [GPHA04]. integrator [Gri05c]. Integrators [CFM06, CP06, Hai03, JL03, KMKZ09, PC09]. Integro [MSW06, BLN04b, BY05, DR04a, MST06b]. Integro-differential [MSW06, BLN04b, BY05, DR04a, MST06b]. interaction [BC09, GHM09, LM00]. interactions [GT07]. interface [AJMN02, GR04, HLPS05, LLW03, OR06, OPR06, PW03]. Interior [HPSS05, ZR05, BW05, BF07, KP08, LPSV03]. Interlacing [DJ07]. intermediate [dV04]. internal [AKK^{+}08]. interpolants [Gau01]. interpolating [AdST05]. Interpolation [AM06, BLM05, CGS06, DR07, Moo01, WR05, ZK00, AdST07, Ber09, BD01, Boc01, BMVX07, CIS06, DV00, DR04b, DQQ01, EL05, FH07, GL05, Joh00, Joh04, Ker03, MS00b, QQ03, Moo03]. Interpolationspolynome [Boc01]. Interpolatory [DGV02]. interval [BP01, BP03, GS00, MC04, MC06, WS08]. Invariant [CFM06, EvP04, PR09, RS02]. Invariants [CFM06, BAL07, Sun07]. inverse [BH03b, FNP07, JT09, LN04b, NW05, RS04, SK03, dFP05]. inversion [RS07, Tsu03, VVM09]. invertibility [KS01]. inviscid [STWW07]. irreducibly [Bou07]. Irrotational [DR04b]. isentropic [FMS07]. isoparametric [AM06, KS01]. iterates [Kim05b]. iteration [Chu06, EN05, Küg05, Lui02, MZ01, Neu00, RT06, VVM09]. iterations [ALW05, ER09, HKT08, KL03, LP08, Nab03]. Iterative [HS06, Yan02, Cao01a, FML00, GS07, Heu98, Heu01b, HN00, LP09a, Pot04, SK03, ZZ09].

Jacobi [Grü04, ACMR06, BS06b, DKM09, GP08, HS08, LYC03, LT01, LN04b, Mat00, MC04, MC06, PP03]. Jean [Gau01]. Jean-Paul [Gau01]. John [Bau07]. jump [JKC09, dFL04]. jump-diffusions [JKC09]. jumps [HK05].

Kantorovich [BB00]. KdV [Zhe06]. Keller [Fil06]. kernel [SS00, Xu00]. kernels [FL06, PSN04]. kind [MFR03, OM01, SW00]. Kinetic [ADM04, KQW04, Bou03, Bou01]. kink [WT07]. Kirchhoff [Per05]. knots [PSN04]. Koenig [BGM01]. Konvergente [Boc01]. Korobov [DSWW06]. Kronrod [Gau05, Not06, PP03]. Krylov [Bec00, Kre06, SS05]. Kummer [ST08]. Kutta [AMN09, AM02, BLV06, CJMR04, CCP07, CCM04, GPHA04, Hag00, HJL09, Hua09, LFLPS05, Mas00, Mas02, Sun07, TJ07].

Lagrange [Boc01, TW01, Ber00, Boc01, BMVX07, BGG04, DK01, HLPS05, LSW05, TW00]. Lagrange-Interpolationspolynome [Boc01]. Lagrangian [Ber01, BC09, Che01, Des01, Gal03, PC09, PM08]. Laguerre [GS00]. lambda [Bor04]. lambda-omega [Bor04]. Lamé [CCS05]. Landau [Bar05]. Landweber [EN05, Küg05, Neu00]. Laplace [HS05, MST06b, SU09]. Laplacian [ABH05, EL05, Kim05a, LY01, Vee02]. Larché [GW05]. Large [HSZ06, DF05, HWW02, Kre06, RS02]. large-scale [DF05]. laser [BC09]. lattice [DSWW06, JY09]. lattices [CGS06]. Lavrentiev [HTY08]. Lavrentiev-regularized [HTY08]. law [BBC^{+}05]. laws [ADM04, BAL07, Ber05, Bou01, CHC01, CL01, CC05, DGMM01, Des01, GR04, HKR01, HJL09, HWW02, KMR09, KP01b, Küt03, LS01b, LWW01, PW05a, Vov02]. Lax [HKR01]. layer [Heu01a, LMV02, TS03]. layers [EZ08, ZR05]. LBB [DK01]. LCP [ZX02]. Least [BL05, Hei09, Lin09, MKV04, PS02a, PS02b, YW06]. Least-Squares [BL05, Hei09, Lin09]. Legendre [Mat01, dFGAN00]. less [FMDH03]. level [BW05, FML00, HMS03, HLS05, HR04b, KS02a, Obe04]. Lie [KMKZ09, Wen01]. like [BN03, Chu06, GS07, HMS03]. like-Newton [GS07]. limit [JL03, LRW06, PTV03]. line [CWRR02, ZZL06]. linéaire [BMV03]. Linear [BR04, CW03, Hil06, MSW06, AM02, BN03, BGP04, BPW05, BM00, BMV03, BGCP05, BG07, BG09, BRZRS03, BJ03, CDRG^{+}07, CM06, CHL08, Da08, DDH07, DF07, EL05, EK02, EMAS01, GGM09, GPW03, KV00, KS02c, LWC09, LZ09, LS01c, LV04b, Lui02, LP01, Lyl00, Mas00, Mas02, MNS08, NSSV06, RSWW03, SW01, SK03, Ste03a, Stö01, WW03, ZOS00, ZLW04]. Linear-programming [BR04]. linearised [LM00]. linearization [Kor02]. Linearly [AKK03]. lines [BFS01, Nie02]. Lipschitz [BHvPS03, BC03b, JKN09]. loading [ST03]. loads [CD04]. Lobatto [MC06, Moo03]. Local [HXZL08, KPS03, LSY09, ACMR06, AMtTB03, EFG06, GM06, Hac07, Hei09, LS01a]. Localization [Fae02]. localized [Ley04]. locally [AM03, TZ03, Voh08, WO00]. locating [BHV03]. locking [HP01]. logarithmic [MSS02]. Long [CCC^{+}05, DF07, HL04a, RAVS07, SW01, TZ03]. long-term [RAVS07]. Long-time [CCC^{+}05, SW01]. look [Bra05]. Lower [BH03c, Bab00, Hes06, PS00]. LR [FP09]. LU [GKB09]. Lyapunov [BL09, DV05b].

M [AXY02, DK04]. M-matrices [AXY02, DK04]. Maclaurin [Sid04]. Macro [LS01a, AS02]. Macro-elements [LS01a, AS02]. magnetic [Noë04]. magnetohydrodynamic [Ger00]. magnetohydrodynamics [Sch04]. manifolds [EvP04, NTW01, PR09, Usc08]. map [CJMR04]. mapping [KS01, MP02, MSS02]. mappings [HXY08]. maps [AP05, BH04, KL00b]. Marching [NW05]. Markov [KNS08]. mask [PR01]. mass [BB00, BRS00, Hac07]. matched [EZ08]. matching [BLS^{+}05b, BBS07, FHV03, HLPS05, LPSV03]. material [BH03a]. materials [CD04, FGSn02, GMZ02, RAVS07]. Math [BSS05, Heu01b, MW06, TW01].

Mathematical [BCP07, CHB08, Gal02, ZLW04]. Matrices [GHNP03, AT03, AXY02, Beb00, Bec00, BS06a, BGP05, Bou07, Cao01b, Cao06b, CG04, DK04, EPS01, HKT08, LP08, Mat00, RS02, SC07, Son00, Son02, Ste01b, Tsu03, VVM04, XG00, XY09]. Matrix [BB06b, DFLZ03, GHK02, AB02, AD07, BB01, BH03b, Cap02, CLM03, DDHK07, FP09, Kre06, LN04b, OR02, PR01, Ste03b, VVVM06b, Vec01, dFP05]. matrix-sequences [Cap02]. maximal [CM08, Gei07, MCS09b]. maximization [Usc08]. maximum [JU05, KK05, TW00, TW01]. maximum-norm [TW00, TW01]. Maxwell [BC09, BCS02, BHvPS03, CLL08, CD02, HPSS05, Lac00, Tos00]. Mean [KL03, DD02, DH06, FP03a, FNP07, HMY07, Obe04]. mean-square [HMY07]. means [BGM01]. measure [KP01a]. measurements [AKK^{+}08]. measures [CR00]. mechanics [BB00, Liu00, Neg03]. media [Pot04]. Mellin [MFR03]. memoriam [Bau07]. memory [CCP07, RAVS07]. meromorphic [GL05]. Mesh [AP05, CFJ03a, CFJ03b, HTY08, FMDH03, Li01, ZL03]. Mesh-independence [HTY08]. mesh-less [FMDH03]. Meshes [FHV03, AM03, ANS01, GHS00a, HLPS05, KV00, Kun00, LYC03, LS01c, SA07, TS03]. Method [BB06a, CFJ03a, CFJ03b, ZCW06, AB02, ARPR04, Ang00, AW07, AMtTB03, Ari04, AG01, Bab00, BFS01, BBO07, BC07a, BV02, BGS02, BC08, BBC^{+}05, BM08, BL05, BLS^{+}05b, Ber00, Ber03, BBS07, BH05a, BH07, BL09, BG07, BG09, BCR04, BCH07, BQR05, Bre03, BCLS08, Buf02, BF07, BGG04, CLR02, CS05, CH06, CS06, CM06, CWRR02, CK02, CLM07, CM00, CS03, hDxY99, Dai01, DI01, DGS07, DQQ01, EH08, FML00, FP04, GPC05, Ger00, GGLV01, GK03a, GK03b, Gri05b, GS00, GH04, HCC09, HXY08, HLPS05, HMMC04, HLS05, Hei09, HO09, Heu98, Heu01a, Heu01b, HPSS05, KS02a, KL00a, Kim05a, KNWN09, KM02, KS07, Kre09, KP08, Kun00, LP09a, LWC09, Li01, Li05, LX09, LYZ01, LP09b, Mah09, MT01, Mar02, Mar01, Mat00, Max07, MM07, Moo03]. method [NN06, Pot04, Pyt02, QQ03, RSS06, Reu02, STWW07, San01, Sch09, Ste00, Ste03a, SU09, Ste02b, TS00, VVM04, VK09, WLJ04, Wil03, Xu00, XC01, YW06, Zha03, ZZL06, ZZ09, dFL04, dV08a, dVLM09, dFGAN00, CJK^{+}05, CJK^{+}06, NP04]. Methods [BDD04, BDR03, BHvPS03, HS06, Hil06, LSW05, AM01, AKK03, AMN09, AM02, ALW05, AK04, BGP04, BPW05, BV07, BFNS01, BV00, BGCP05, Bor04, BW05, BF01, BY05, BCS02, BJ03, CJMR04, CCP07, Cao01a, Cao06a, CYY07, Cap02, CBJ02, CH07, CCM04, CCF08, CLL08, CX08, Chu06, Da08, DDP07, DF05, DFLZ03, EK02, ES08, Fae02, FMDH03, GGM02, GLQ01, GP03, GHS00a, GS02, GS07, GNP08, GP08, GZ01, GLRS09, HKS07, Hag00, HL04a, HLW03, HK05, HTY08, HS05, HRW02, HS08, Hua09, HN00, IT08, JT09, JY09, Kan01, Koc05, KW08b, KW08a, KV00, KV01, KS02c, KPS03, LW00, LN04a, LLW03, Lin09, LSY09, LS09, LV04b, Lui02, MT03, MN06, MP05, Mas00, Mas02, MR03, Nab03, NAE07, NTW01, OM01, PMC01, PTV03, PS09a, RXH05]. methods [RSWW03, Sch04, SW08, SC07, SK03, SS05, SS08, Sti05, Stö01, Sun07, ST05, TT05, Tai03, TJ07, Tos00, Voh08, WY06, Wen01, Xia07, Yan02, ZX02, DV05b]. MHD [BKW07, Gal03]. micromagnetics [CP01a, KP01a, LX09]. microstructure [BP04]. microstructures [Li01]. midrange [Nie02]. mild [EH08]. mimetic [BLS^{+}05b, LSY09, dVLM09, dV08a]. Mindlin [CS06, Lyl00]. Minimal [HN00, BF01, HR00, SS08, Sti05, vGLS02]. Minimally [Wal00]. minimization [DFLZ03, GP08]. minimum [CF08, JU05]. MITC3 [McS02]. Mixed [AW02, GH03, Sch04, TS03, BG01, BGS02, BM08, BF01, BGG04, Cao06a, CCS05, CS06, CFA06, CM06, CS00a, DE06, FP04, GGM02, GPW03, Hol08, Kan01, KK05, Kim05a, LM08, LV04b, RPK08, SKR04, SW03, Yan02, Yse04, GM05]. mixed-hybrid [EHV06]. mobility [BB01]. model [AB03, BB01, BB04, BRS05, BRM^{+}04, CHB08, Fil06, GGJ03, GW08, HP01, Neg03]. Modelling [SŻ05, ACMR06]. models [BBC^{+}05, Ber02, Bou03, BC07b, CK08, Hac07, Yse01, dSB08]. Modified [BCG06, Hai03, HMMC04, ZK00, ZZL06, Zhe06]. moment [HSS08b]. momentum [CHL08, OWW04]. Monge [BB00]. monotone [AG01, Ber01, HLW03, Hol08, JU05, Lui02, NSSV06, Obe04, Vov02, ZX02]. Monotonic [MST06a]. Monotonicity [Bou07, NAE07]. Monte [CDLP07]. Moreno [PCM02]. Morley [HS09b, MX06, dVNS07]. Mortar [BDR03, LSW05, BBCM00, BLS^{+}05b, Dry05, Mar02, RXH05, XC01, GGM02]. mothers [BBS07]. motion [Obe04]. mountain [Hor04]. Multi [BRZRS03, HLS05, BB01, CJMR04, GHL^{+}05, HS08, KS02a, KW06, MR03, Ste02b, Sun07]. multi-component [BB01]. multi-grid [KS02a, Ste02b]. Multi-level [HLS05]. Multi-parameter [BRZRS03]. multi-revolution [CJMR04]. multi-scale [GHL^{+}05]. multi-symplectic [HS08, MR03]. multi-symplecticity [Sun07]. multidimensional [KP01b]. Multigrid [AFW00, CK02, HRW02, XC01, AD07, Bor04, Kor02, LWC09, Reu02, Tim00, VBM01]. multiharmonic [BLS05a]. Multilevel [AM03, DDP07, HSS08a, AD07, GK03a, GK03b, Osw08]. multiple [CNY02, ER09, MCS09b]. multiple-zero [CNY02]. multiplication [DDHK07]. multiplicative [BFNS01, Nab03]. Multiplier [LSW05, Ber00, GGLV01, HLPS05]. multiplier/fictitious [GGLV01]. multipliers [BGG04, DK01]. Multiresolution [BSS04, BSS05, DGMM01, SPSF00]. Multiscale [BP04, Cao06a, EW02, GLS07, HO09, Neg08, PS09a]. multisplittings [Yun03]. multistep [CP06, GZ01, HL04a]. Multivariate [BR02, Olv06]. multiwave [BKW07]. Mumford [BC00]. MUSCL [Ber06, HCC09]. MUSIC [AKK^{+}08]. MUSIC-type [AKK^{+}08].

Navier [AGS09, ACC^{+}02, ABC06, Bän01, BF07, Cas01, CC05, CB00, DH04, ESW02, GGLV01, HL04b, HLS05, JY09, KL00a, WL02]. Near [Ker03, CJMR04]. Near-interpolation [Ker03]. nearly [Ste03a, dVLP06]. Necessary [ZOS00, LN04b, TMO04]. Negative [SW00, HPS03, XY09]. Neighborhood [BCM06]. Nernst [PS09b]. Nessyahu [PT06]. nested [Zha04]. nets [CDLP07]. networks [CC00]. Neumann [Dry05, GPW03, NR03, Nov01, dVLP06]. Nevanlinna [GL05]. Newton [Chu06, DFLZ03, DQQ01, GS07, Hac08, HTY08, JT09, Kor02, Kre09, LR09, QQ03, Rie01, WY06, WS08]. Newton-like [Chu06]. no [FH07]. Nodal [BB06a, BCJ09, CD02, DV00]. node [ZK00]. nodes [EPS01, MCS09b, Pet03]. noisy [HPS03]. Non [BB06a, FHV03, Olv06, AT01, AJMN02, ACMR06, AP01, BGP04, BMV03, BV00, BLS^{+}05b, BGCP05, BBS07, BH04, BCS02, CR00, CHL08, HLPS05, HL04b, JU05, LPSV03, MT05, Mül03, PTV03, Pot04, SS05, SS08, WcSX07, XY09]. non-conforming [AJMN02, MT05, WcSX07]. non-convex [AP01, CR00]. non-Hermitian [BGCP05, BGP04]. non-homogeneous [AP01]. non-hyperbolic [BH04]. non-iterative [Pot04]. non-linear [CHL08]. Non-matching [FHV03, BLS^{+}05b, BBS07, HLPS05, LPSV03]. non-monotone [JU05]. non-negative [XY09]. non-optimal [SS05]. Non-overlapping [BB06a, BBS07]. Non-polytope [Olv06]. non-quasiconvex [Mül03]. non-smooth [BV00, BCS02]. non-stagnation [SS08]. non-stationary [HL04b]. nonautonomous [PR09]. Nonconforming [Ste02b, BCLS08, CBJ02, CH06, CH07, LY01, McS09a, Osw08, XC01]. nonconforming/mixed [EHV06]. nonconvex [BR04, GGM09, LWW01]. nonequispaced [PSN04]. nonhomogeneous [CHC01]. Nonlinear [DSX00, Hil06, Sze06, AKK03, ACP02, BLS05a, BLN04a, BB04, BGS02, BG07, BD02, BGG04, CD04, CCF08, Chu06, CM08, Dai01, DD04, DM04, DV08b, EK00, EGHM02, EFG06, GGJ03, GPC05, GR04, GK01, GNP08, GS00, HKR01, HMMC04, Hei09, HK05, HTY08, HZ02, JT09, KS02a, KK05, KS07, Kre09, KL03, KMS04, Liu00, Lui02, Mah09, MR01, Neu00, Ohl01, Per05, RT06, Tai03, Vee02, WP00]. nonmatching [Buf02]. nonnegative [Son00]. Nonnegativity [GR00]. nonoverlapping [BH03c]. nonquasimonotone [WP06]. nonsmooth [BNA07, GGM09]. nonstationary [TT05]. nonsymmetric [AK04, BN03, Cao06b, GLX06, LP08]. nonuniform [SA07]. nonzero [Deu09]. Norm [EL05, BSS04, BSS05, Fae02, LY01, Not06, SW00, TW00, TW01]. Normal [DF07, FGSn02]. norms [AdST05, BB08, San04]. Note [Pet02, Gau01, Cao01b, EN00, Gau05, Gri05b, NE00, Peñ06, SBL06]. nuclear [AC01]. Num [BSS05, MW06]. Number [CW03, Bec00, Bey00, VK09]. numbers [KNS08]. Numer [Heu01b, TW01]. Numerical [ACP02, AB02, AP01, BLS05a, BM00, BRS05, BMV09, CR00, CP01a, DR04a, EvP04, FP03a, FGSn02, FHV03, GW05, GW08, HMMC04, HK05, HS05, JL03, LM00, LYC03, Neg08, Pam01, RAVS07, Weg02, Zhe06, AD01, Ang00, BGL07, BCP07, BQR05, BLN04b, BC03b, BJ00, CK07, CC04, CLRM01, CHB08, CF08, CLM07, CHL08, DD02, DW00, DGS07, FMS07, GGJ03, GR04, GP03, GK01, Hor04, KNWN09, KMKZ09, LLP07, LX09, MT03, MP02, MP05, MSS02, MM07, PTV03, Per05, PY02, STWW07, Sid05, TMO04, TZ03, Yse05, Zha03, BRM^{+}04]. numerical-asymptotic [DGS07]. Numerically [ST08, DHN07].

object [KM02]. observability [Erv09]. observations [XZ03]. obstacle [BBP01, BC04, Bra05, BCH07, CN00, NSV03, hW02, Zha07]. ocean [GGRG05]. oceanography [RG04]. off [PTV03]. Oldroyd [AB03]. omega [Bor04]. One [PT06, Sze06, BFS01, CR05, DW00, DJ07, EFG06, FP09, GP08, LN04a, Moo01, VVVM06b]. One-dimensional [Sze06, CR05, EFG06, GP08]. one-parameter [DJ07]. one-point [FP09]. One-sided [PT06]. open [BH05a, BH07]. Operator [DM04, IT08, AG00b, BBD06, Bof00, DHN07, GHK02, HXY08, HLW03, Heu01a, Noë04, San04, SU09]. Operators [BLM05, AKR09, AD00, AK04, BH03b, BH05a, BH07, BG05, DR07, Gei07, GM06, Heu98, Heu01b, Hip01, Huh07, NR03]. Optimal [AMN09, Bab00, BV07, FL06, HS06, McS02, NP04, PW03, SW01, AC01, AD01, BC08, Bey00, BLV06, BY05, DXZ08, GH04, Hag00, HR04b, HTY08, KNS08, Kwe00, LPSV03, LY03, Peñ02, Peñ06, SS05, TJ07]. Optimality [Osw08, RS07, AD07]. optimally [BM08, Leb02]. Optimization [EH06, AM08, AP05, BNA07, GGM09, SZ09, SHS08, WY06]. optimized [LS09]. option [AW07]. options [FO09, IT08, LHJB07, dFL04]. orbits [BH04, CKP07, Pam01]. Order [BLM05, LSW05, AGS09, AM01, AMN09, AM02, BNZ05, BC07a, BLN04a, BNA07, BLV06, BCGJ07, CC04, Cao06a, CDRG^{+}07, CK08, CM00, DM04, Gri05c, Gri05b, GK01, HPS03, Hei09, Hua09, JKN09, Joh00, KP01b, Kwe00, Lin08, MN06, Mar01, MT05, MM07, MX06, Pro00, RT02, STWW07, San04, Sti05, zSlZ04, VV03, WW03, WLJ04, dVLM09]. ordinary [MSTZ03a, MSTZ03b]. Orlicz [DR07]. Orthogonal [SK03, VVVM06b, AM01, DJ07, KV01, Usc08, VVVM06a, dSB08]. orthogonalization [GLRvdE05]. Orthomin [ZOS00]. Orthonormal [DV05a]. Osborn [Tsu03]. oscillation [DN02, FL06, PW05a]. Oscillatory [Olv06, DH09, EW02, Gri05c, Gri05b]. ovals [Peñ03]. Overlapping [Tos00, BB06a, BBS07, Kim05b]. overrelaxation [BPW05].

P1 [XC01]. Padé [BR02, Mat01, Put02]. Padé-Bergman [BR02]. Padé-Legendre [Mat01]. Padua [BMVX07]. pairs [Che07]. panel [HKS07]. paper [Mül03]. Parabolic [MSW06, Bak02, BLN04a, BQR05, BLN04b, CFA06, CS00b, CS03, CM08, DT07, EK00, EGHM02, EHV06, EFG06, GGM02, GST04, GK02, KV01, Ley04, LY03, LFLPS05, MST06b, Moo01, Mos01, PY02, RPK08, SW08, TW00, TW01, TJ07, Wil03, ZR05]. parallel [ACC^{+}02, BOL02, HXZL08]. parameter [BV04, BRZRS03, DJ07, HP01, KS02b, Küg05, MPRS01, NP04, OPR06]. Parameters [WR05, PMC01, PCM02]. parametric [Dzi08]. Part [mWS05, WS08, BRM^{+}04, CJK^{+}05, CJK^{+}06, CFJ03a, CFJ03b, EN00, Fae02, LYZ01, NE00]. partial [BMV09, CYY07, GS00, Hei09, SW08, VC01]. partially [HXY08]. particle [BBO04, Yse01]. partition [DDP07, KPS03]. partitioned [BLV06, Sun07]. partitioning [KP08]. pass [Hor04]. patch [FG02, ZL03]. patches [GHL^{+}05]. Pathwise [JKN09]. Patlak [Fil06]. Paul [Gau01]. PDE [SZ09]. PDE-constrained [SZ09]. PDEs [BCM06, CS04, HJL09, Küg05, Lui02, PS09a, Sun07, CK07, Cap02, GLS07].

Peaceman [DR03]. PEERS [LV04a]. penalised [CP01a]. penalized [AW07]. penalty [BW05, BF07, HPSS05, LPSV03, LP09a, ZR05, dFL04]. Pencils [BB06b]. Perfectly [EZ08]. perforated [CC04, Cao06a, HO09]. Performance [ESW02, Yun03]. periodic [ED02, JY09, Kre06]. perspective [Sid04]. Perturbation [GH08, AXY02, CP01b, Che07, DM04, XG00, XY09]. perturbations [Bou07, Ste03b]. perturbed [BCGJ07, CX08, Lin09, Mar01, Ohl01, Ste05]. phase [BB01, BBC^{+}05, CE01]. phenomenon [WT07]. Phillips [MPRS01]. photonic [BCG06]. Pick [GL05]. Piecewise [GGM09, BH05a, BH07, EL05, ED02, Pot04]. pipes [Zha03]. pivot [KW08a]. planar [LWC09]. Planck [PS09b]. plane [BH05a, BH07, CWRR02, DHN07, GM05, Heu01a]. planes [Nie02]. planetary [STWW07]. plasma [BC09]. plastic [Liu00]. plasticity [HR00]. plate [CS06, DHNL^{+}00, Joh00, Lyl00, McS09a, Mar02, PS00, dVNS07]. plate-bending [PS00]. plus [BN03, CG04, VVVM06a, VVVM06b]. Point [SZ03, BGS02, BS06a, Cao06b, ESW02, FP09, HZ02, KS02b, KW06, LP08, Mar01, Moo01, OPR06, SZ09, Sti05, WW03]. points [BMVX07, TMO04, ZL03]. Pointwise [Ley04, NSV03, NSSV06]. Poisson [CJK^{+}05, CJK^{+}06, PM08, PS09b]. polar [KMKZ09]. poles [Chu03, FH07, dlCYGV07]. polyconvex [CD04]. polygons [BNZ05, MS00a, SW03]. polyharmonic [BB08]. polyhedral [BCD05, BCJ09, CD02, GLRS09]. Polynomial [GM06, ED02, MFR03, MS00a, Pet03, vGLS02]. Polynomials [PP03, AM01, BS06b, Boc01, DJ07, GH08, KR09, KL00b, Moo03, Pas00, SK03, dFGAN00]. polytope [Olv06]. population [BB04, GGJ03, Hac07, KW08b, KW08a]. ports [RVH09]. posed [CCP07, KS02a, LMV02, Neu00, RT06]. posedness [BKRT04, Usc08]. Positive [Gal03, dVLP06, BGP04, Bec00, BGCP05, Cao01a, Cao01b]. positivity [CK08]. post [ARPR04]. post-process [ARPR04]. posteriori [ABC03, AMN09, ABR06, BFS01, BGS02, BC04, BCD04, Bar05, BV04, BMV03, Ber02, BG09, BCR04, Bra05, CMS01, CBJ02, CS04, CMPS04, CCS05, Car05, CS06, CH07, CGJ09, CN00, CM08, FV06, GHM09, HN05, HS09b, KS02a, KV00, Kun00, LM08, LY01, LY03, LV04b, MPRS01, MN06, Max07, Moo03, NSV03, NSSV06, Noë04, Ohl01, San01, SW08, Voh08, dVNS07, Moo01]. postprocessed [dFGAN00]. Potential [HS06, DF07, Hac08, Heu01a, WO00]. potentiels [BBP01]. pour [BBP01, BMV03]. power [MSS02]. powers [Ste03b]. pp [TW01]. Precise [Tsu03]. Preconditioned [BGP04, BGCP05]. preconditioner [AG00b, Bre03, RS02]. Preconditioners [BN03, TGSS08, AM03, BW05, Cao06b, Cas01, CC00, CNY02, ESW02, GPW03, MW04, MW06, NCV06, OPR06, Yun03, dVLP06]. Preconditioning [EN05, AK04, DDP07, GKB09, HWZ06, KO00, Osw08]. predator [GT07]. predator-prey [GT07]. predictor [CC05]. predictor-corrector [CC05]. prescribed [DH06]. presence [RVH09]. Preserving [CFM06, HMY07, AKR09, BP08, BC07b, CK08, Gal02, GT04, GR00, GLX06]. pressure [ARPR04, ABC06, BGHNR03, CC05, CB00]. pressure-stabilized [CB00]. Prewavelet [BN01]. prey [GT07]. Pricing [FO09, AW07, CF08, IT08]. Primal [Li05, BG01, HR00, MT01]. primitive [GGRG05, Med01, RG04]. principal [Str08]. principle [Jer08, JT09, LLP07, MSTZ03b]. principles [JU05, KK05]. priori [ABC03, BCD04, CD04, FP01, Hol08, KS02a, Küt03, LY01, PW03, SKR04]. prismatic [KS01, CJK^{+}05]. Problem [FHV03, AD01, AB03, ADM04, BV02, BC08, BGL07, BBD06, BAL07, BBCM00, BB00, BL05, BMV03, BGHNR03, BG09, BCH07, BCLS08, CR05, CC04, CFA06, CWRR02, Che01, Chu03, CJK^{+}05, CJK^{+}06, Cop08, DSS02, DHNL^{+}00, FL06, FGSn02, GM05, GHM09, HR00, HP01, HL04b, HLS05, HXZL08, HN05, HR04b, HS05, Hol08, Kim05a, KO00, KK00, LM00, LN04b, Mar02, MW04, Med01, Noë04, Nov01, OR06, OPR06, RAVS07, RVH09, SW01, SU09, ST05, hW02, WO00, MW06]. problème [BMV03]. Problems [BDR03, BHvPS03, CW03, SZ03, ZCW06, AG00a, AM08, AM02, AP01, ABR06, BFS01, BNT05, BLS05a, Bak02, BG01, BC04, BR04, BV04, BM00, BG07, Bra05, BMS00, BLN04b, BY05, BCJ09, BCGJ07, BJ00, BGG04, CP06, Cao06a, CLRM01, CR00, CCF08, CN00, CX08, CK02, CM00, CM08, DGH09, DI01, DV08b, DE06, Dry05, DT07, EL05, ER09, EW03, EW02, ES08, EHV06, EFG06, FML00, FP03b, GPC05, GK02, GHL^{+}05, GPW03, GS02, GNP08, HHW00, HLPS05, HSS08b, HO09, HR04a, HTY08, Hor04, HZ02, JT09, KS02a, Kan01, KK05, KNWN09, Koc05, Kre09, KL03, KV01, KS02c, KP01b, LW00, LN04a, LM08, LPSV03, LLP07, LWC09, LLW03, Lin09, LS01c, Lin08, LMV02, Liu00, LY03, LS09, Mah09, MN06, Mar01, Moo01, MNS08, NW05, Neu00]. problems [NN06, NSV03, PW03, RXH05, RT06, RSS06, RT02, San01, ST03, SZ09, SŻ05, Sti05, Stö01, SHS08, TT05, Tim00, Usc08, Yan02, ZR05, Zha07, dV04, BC03a]. procedure [BF01]. procedures [FvOP05]. process [ARPR04, GLRvdE05]. processes [dFL04]. processing [HMS03, MR01]. product [CLM03, GS08, HSS08a, Usc08]. product/quotient [CLM03]. products [Kre06]. Programming [CW03, BR04, DF05]. programs [EK02, PK05, ZLW04]. Projected [DF05, Sch09]. projection [BC07a, Gro09, HPS03, KS02b, Pro00, RT06, Ste01a, Ste02a, VK09, ZX02, dSB08]. projection-based [Pro00]. projection-regularized [KS02b]. projection-type [ZX02]. projections [Gri05a]. proof [AD07]. propagation [Yse01]. proper [KV01, dSB08]. properties [AT03, AKR09, CLR02, Cao01b, Dai01, FL06, Gro09, Mat01, VVVM06a]. property [PW05a, Ste05]. Pseudo [BB08, GM06]. pseudo-differential [GM06]. Pseudo-polyharmonic [BB08]. pseudodifferential [HPS03, TGSS08]. pseudospectral [BLO07]. Publisher [Pet02]. pyramidal [KS01].

QR [BGP05, CP01b, DV05b, VVM04]. QR-method [VVM04]. Quadratic [CFM06, Mat00, Sun07, DF05, DV00, hW02]. quadratically [MT03].

Quadrature [GGL00, MCS09b, Olv06, PP03, BP01, BP03, CCP07, DGV02, DV00, DH09, FL06, Fis02, Gau05, LS08, MC04, MC06, Not06, OST07, PGL07, SL02, SS00, WT07, dlCYGV07, de 05]. quadratures [PK05]. quadrilateral [Zha04]. quantities [Can06]. quantum [JL03, MST06a]. Quasi [EL05, LY01, YW06, Ber01, BG09, Cop08, CDLP07, WY06, DFLZ03, Maz08a]. Quasi-least-squares [YW06]. quasi-linear [BG09]. quasi-monotone [Ber01]. quasi-Monte [CDLP07]. quasi-Newton [WY06, DFLZ03]. Quasi-Norm [EL05, LY01]. quasi-static [Cop08]. quasiconvex [Mül03]. quasilinear [KK00]. quasistatic [RSWW03, SW01]. quotient [CLM03].

Rachford [DR03]. Radau [MC06]. radial [Joh04, PSN04, TGSS08]. radiative [BC07b, GT04]. radius [GZ01]. rank [DW00]. Rate [FvOP05, Tai03, BNZ05, CDLP07, GH04, Küg05, LZ09, LHJB07, Pam01]. Rates [BDD04, FH07, JKN09, LWW01, MSTZ03a, Rie01]. Rational [VVM09, dlCYGV07, DV05a, FH07, GGL00, LSW02, PGL07, de 05]. Raviart [ANS01, HS09a, RXH05]. Rayleigh [KNWN09]. reacting [Pro00]. reaction [BCGJ07, DR03, DM04, Jer08, Lin09, Lin08, Ste05, WP00, WP06]. reaction-diffusion [BCGJ07, DR03, DM04, Lin09, Ste05, WP00, WP06]. reaction/diffusion [Jer08]. reactor [AC01]. real [Bec00, CWRR02, GST04, Zhe06]. reconstruction [DSX00, Pot04]. recovery [FV06, LP09b, MP02, Ova06, ZL03]. rectangular [EPS01]. recurrence [ST08, VC06]. Recursive [Mat01]. reduced [BC09]. reduction [AM02, CLM03, DM04]. Reeves [ZZL06]. refinable [PR01]. refined [AM03, WO00]. Refinement [CFJ03a, CFJ03b, Bey00, Zha04]. region [BD02, WY06]. regularity [ABH05, CM08, Gei07, HR00, Yse04]. regularization [BRZRS03, CD02, FvOP05, HPS03, KS02a, MPRS01, Zha03]. regularizations [LR09, Rie01]. Regularized [BH05b, HTY08, KS02b]. regularizing [CLR02]. rehabilitation [CD02]. Reissner [Lyl00, CS06]. related [KP01b, Med01, Ste02a, Voh08]. relations [ST08]. relative [DKM09, XY09]. relativistic [KQW04]. relativity [FNP07]. relaxation [BBC^{+}05, BKW07, LS01b, Mül03, NP04]. relaxed [CP01a]. Reliable [BOL02, BC04]. representations [GST04, SS00, Wal00]. reproducing [Xu00]. Residual [BMS00, CMPS04, CS06, CN00, Voh08, BG09, EW03, GHM09, HN05, Kun00, Max07, SS08, dV08a, San01]. Residual-based [CMPS04, CS06, EW03, GHM09]. Residual-free [BMS00, San01]. residual-type [BG09]. residuals [hDxY99, KV00, Pyt02]. resolution [BP04, Neg08]. response [GT07]. result [Jer08, SC07]. Results [HS06, AD01, GJ02, SHS08]. retarded [BMV09]. retardés [BBP01]. Reversible [CCM04]. revolution [CJMR04]. RFB [CS05]. Riccati [RS02, GLX06]. Riccati-based [RS02]. Richards’ [SKR04]. Riemann [BAL07, BKW07]. Riesz [SL02]. right [CDRG^{+}07]. right-hand [CDRG^{+}07]. RK [HS08]. Robin [AJMN02]. Robust [KW06, LWC09, Lin08, San01, Ste03a]. Robustness [Ber02]. Romberg [Fis02]. root [Str09]. Rosseland [GT04]. rot [Lac00]. rotation [YW06]. rotor [SHS08]. rough [Bak02, KW06]. Rounding [GLRvdE05]. Rule [mWS05]. rules [But04, DSWW06, GGL00, MC06, WS08]. Runge [AMN09, AM02, BLV06, CJMR04, CCP07, CCM04, GPHA04, Hag00, HJL09, Hua09, LFLPS05, Mas00, Mas02, Sun07, TJ07].

Saddle [SZ03, BGS02, BS06a, Cao06b, ESW02, HZ02, KS02b, LP08, OPR06, SZ09]. saddle-point [HZ02]. satisfactory [ST08]. satisfying [Bou03, FL06]. saturation [DN02, Ste05]. scalar [Bou01, CIS06, CHC01, GR04, HWW02, PW05a, Vov02]. scale [DF05, FMS07, GHL^{+}05, RSS06]. Scaled [PS02b, Mat00, PS02a]. scales [EN05, Neu00]. scaling [AM03]. Scattering [GHNP03, CWRR02, DGS07, ER09, KM02, LS08, NW05, NN06]. scenario [BNT05]. schema [BMV03, BMV03]. Scheme [Ber06, AGS09, AT01, ACMR06, BC09, BC07b, BCGJ07, BKRT04, CL01, CF08, CK08, DD04, DD02, DV08b, DE06, EGHM02, EHV06, EFG06, Fil06, Fur01, HKR01, HMS03, HWW02, KP01b, LS01b, LHJB07, Lin08, Mau06, MV07, MR01, Neg08, Obe04, PM08, PT06, Pro00, RT02, zSlZ04, VV03]. Schemes [Gal03, Ber01, Ber05, BLV06, Bor04, Bou01, BLN04b, CHC01, CC05, DGMM01, Des08, DF07, Erv09, GT04, Gro09, GR00, GK01, JV01, KMR09, KO00, KQW04, LYC03, MST06a, MM09, NW05, Vov02, WL02, Yan02]. Schmidt [GLRvdE05, SBL06]. Scholes [AW07]. Schrödinger [AD00, DD04, DF07, Neg08, Noë04, Sze06, Yse04, Yse05]. Schur [CLM07, Kre06]. Schwarz [AG00b, BFNS01, BDR03, Bre03, BW05, Cas01, Heu01a, Kim05b, LW00, LS09, Lui02, Nab03, RXH05, SZ03, SZ09, Tos00, TS00]. Schwarz-type [SZ03, SZ09]. screen [Heu01a]. screens [BC03b]. SDFEM [LS01c]. search [ZLW04, ZZL06]. Second [AM01, BNA07, AGS09, CC04, Cao06a, CDRG^{+}07, CK08, CM00, Gri05c, Gri05b, HSS08b, Hei09, Mar01, MFR03, MM07, OM01, RT02, SW00, zSlZ04, VC06, WW03]. second-kind [SW00]. Second-order [BNA07, CK08, Gri05c, Gri05b, MM07, WW03]. section [CWRR02]. sectional [KW08b, KW08a]. secular [BGP05]. sedimentation [BKRT04]. Segel [Fil06]. Selecting [WR05]. selection [BBO04, MPRS01]. Self [HPS03, HLW03]. Self-adaptive [HLW03]. Self-regularization [HPS03]. Semi [GGJ03, HS06, HMS03, MR01, AG00a, AdST05, BB08, Ber01, BC09, CM06, DD04, EZ08, GP03, GPHA04, GS00, KP01b, Mic02, NSSV06, PM08, SK03, TZ03]. semi-coercive [AG00a]. semi-discrete [CM06, DD04, EZ08, KP01b, Mic02]. Semi-discretization [GGJ03, TZ03]. Semi-implicit [HMS03, MR01]. semi-infinite [GS00]. Semi-Iterative [HS06, SK03]. semi-Lagrangian [Ber01, BC09, PM08]. semi-linear [NSSV06]. semi-norms [BB08]. semi-stable [GP03, GPHA04]. Semiconvergence [Son00]. semidefinite [BGP04, Cao01b]. semidiscrete [SW01]. Semidiscretization [DH04]. semidiscretizations [OWW04, TW00, TW01]. semilinear [CP06, CYY07, Hor04, OWW04, Tim00]. semiseparable [CG04, VVM04, VVVM06a, VVVM06b]. semismooth [HTY08]. Sensitivity [BJ00, KR09, OST07]. separation [BB01]. sequence [Fis02]. sequences [Cap02, DJ07]. serendipity [ZK00]. series [FO09, Lac00]. set [HMS03, HR04b, MP02]. sets [Obe04]. several [CGS06, CC05]. shadowing [CKP07]. Shah [BC00]. shallow [DMC00, PMC01, VK09]. Shape [AKR09, CZ03, EH06, BBO04, BS06b, SHS08]. Sharp [CE01]. shell [dV04]. shifted [DJ07]. Shishkin [LS01c]. shocks [LWW01]. shortest [hDxY99, KP08, Pyt02]. shortest-interior-edge [KP08]. side [CDRG^{+}07]. sided [FG02, PT06]. Signorini [Che01, Cop08, HN05]. similarity [VVVM06a, VVVM06b]. Simplicial [Bey00, TW00, TW01]. Simplification [SC07]. simplified [BCP07]. simulation [FMS07, HMY07, MT03, Zhe06]. simultaneous [PGL07]. Simultaneously [CCSS09]. sinc [Ber09, Gau01]. SINC- [Gau01]. single [Ber05, Heu01a]. Singular [CJK^{+}05, CJK^{+}06, AP05, BH07, Cao01b, DM04, Koc05, Pot04, SK03, Son00, VVM04, Wil03, ZOS00]. singularities [Lin09, MFR03, PW05b, Sid04]. singularly [BCGJ07, CX08, Lin09, Mar01, Ohl01, Ste05]. sixth [BLN04a]. skew [BGP04]. skew-Hermitian [BGP04]. Slobodeckij [Fae02]. Small [DN02, BHV03, CD04, DF07]. Smolyak [Pet03]. Smooth [AS02, BV00, BCS02, CS03, GS02, Sid05]. smoothed [VBM01]. Smoothers [SZ03, SZ09]. Smoothing [WR05, AdST07, EK02, QQ03]. smoothing-type [EK02]. Smoothness [Gro09]. Sobolev [AdST07, DR07, JJL02, Ste01a, Ste02a]. solid [BRS00, GHM09]. Solution [Bor04, Lac00, Mah09, MSW06, Ang00, BOL02, BM00, BB00, BRS00, BGCP05, CK07, EK02, FML00, HLPS05, HR04b, HS05, Hor04, MNS08, SK03, Yse05, BC03a]. solutions [ACP02, BLN04b, DR04a, ED02, EK00, GGJ03, GM06, JKC09, Jer08, KK05, KNWN09, Küt03, LT01, OR02, ST08]. solved [BFS01]. solver [BKW07, Bou01]. Solving [BCJ09, ACC^{+}02, Chu06, GPC05, HMS03, Hei09, JT09, KW08b, KW08a, MR01, Stö01, ZLW04]. Some [XZ03, AT03, AD00, Bou07, Can06, DDP07, GPC05, GS07, JT09, LLP07, Lyl00, Mar02, Mat01, RPK08, Ste02a, Tai03]. SOR [HS06]. sound [Yse01]. source [ABR06]. sources [Pot04]. Space [CFA06, CFJ03a, CFJ03b, GK02, BF07, HWZ06, Moo01, RG04, TZ03, Yse07]. spaces [AKR09, AdST07, Chr07, CM08, Da08, DSWW06, DR07, Fou09, GS08, Gri05a, HSS08a, JJL02, Maz08a, Maz08b, Ste01a, Ste02a, Yse04, Yse05]. Sparse [HSS08b, LP09b, ST03, Yse05, GH03, HSS08a, LYZ01]. sparsity [RT06]. spatial [OWW04, HLS05]. spatially [GT07]. Spectral [AD00, ABC06, BLO07, Erv09, PTV03, BBCM00, Cas01, GS02, GZ01, HLS05, HR04a]. spectrum [FP09]. sphere [AP05, GM06, Sim00, TGSS08]. spheres [BP08, Hes06, Nie02]. Spline [CS00b, AdST07, CS03, DV00, Gri05a, Joh00, SA07]. splines [AdST05, HRW02, LS01a, LP01, Mar01, vGLS02]. splits [AS02]. splitting [BGP04, Ber05, Bou01, CLL08, DM04, DF07, GGM02, HXY08, HO08, HLW03, IT08, Pro00]. splittings [Bou03, Cao01b, Huh07, Son00, Son02]. SPR [ZZ03]. Square [Str09, HMY07]. Square-root [Str09]. Squares [BL05, Hei09, Lin09, MKV04, PS02a, PS02b, YW06]. stabilisation [BF01]. stabilised [Buf02]. Stability [AG01, CX08, Chr07, Da08, ED02, HCC09, Hil06, KP08, LT01, Mas00, TW00, TW01, VK09, Chu03, GZ01, HMY07, Hua09, LX09, LV04a, Peñ02, Peñ06, PT06, SW01, Ste01a, Ste02a]. stabilization [Ber00, Leb02]. Stabilized [ES08, BV02, BV07, Ber03, CB00, Ger00, RG04]. Stable [Gal03, GPHA04, Ber06, BGP05, CS00a, DDH07, DDHK07, Fur01, GP03, LS01a, MT05, VC06]. stage [Cao01a]. staggered [HKR01]. stagnation [SS08]. staircase [MS00b]. standard [DKM09, LP09b]. state [ESW02, HR04b, HTY08]. state-constrained [HR04b]. static [Cop08]. stationary [CB00, HL04b, Neg08, Sch04]. steady [Cas01, ESW02, GGLV01, Jer08, YW06]. steady-state [ESW02]. step [GS07, HWW02]. stepping [Mau06, SW08]. stepsize [GZ01]. Stiefel [Usc08]. Stieltjes [PP03]. stiff [AG01]. stochastic [CK07, CYY07, CC00, DD04, Grü04, HSS08b, HK05, HMY07, IT08, JKN09, PK05, ST03, XG00]. Stokes [MW06, HXZL08, AGS09, ACC^{+}02, AB03, ABC06, Bän01, BV02, BBD06, BBCM00, BF07, CCS05, Cas01, CC05, CS00a, CB00, DH04, DSS02, ESW02, GGLV01, HL04b, HLS05, JY09, KL00a, KO00, Kwe00, Li05, MW04, Med01, OR06, OPR06, SW03, WL02]. Stopping [ALW05, Ari04]. strain [BCR04]. strategies [KPS03]. stream [DSS02]. stress [AB03]. stress-vorticity [AB03]. Strictly [Kan01]. string [Per05]. strongly [AG01, GNP08]. structure [DM04, Gal02, GLX06, LM00, Sch09]. structure-preserving [Gal02, GLX06]. structured [HKT08]. structures [CS04]. study [CHB08]. subdivision [Gro09]. subject [CLRM01]. subspace [Chu03, SS05, Str09, WY06]. subspaces [RS02]. substructuring [Heu98, Heu01b, LP09a, MT01]. successive [BPW05]. sufficient [LN04b, TMO04, ZOS00]. summability [Vec01]. sums [VC01]. sup [BC07a, BBCM00, GS08, Leb02, MT05]. Superapproximation [Gri05a]. Superconvergence [BBO07, mWS05, FV06, WS08]. superconvergent [Ova06]. Superlinearly [AK04]. supported [Wal00]. Surface [GHNP03, Bän01]. surfaces [BH05a, BH07, CMS01, CMPS04, CIS06, DH06, GS02, Sid05]. SVD [CLM03]. SVDs [DK04]. Symm [MS00a]. Symmetric [HL04a, KS02c, KMS04, Cao01a, Cao01b, CG04, LPSV03, MNS08, VVVM06b, VV03]. symmetry [ZL03]. Symm’s [CMS01]. symplectic [BLV06, Hai03, HS08, MM09, MR03]. symplecticity [Sun07]. synthesis [BNA07]. system [BC09, BQR05, DMC00, GJ02, Hag00, Kwe00, LZ09, PM08, PS09b, YW06, dSB08]. Systems [HSZ06, BN03, BGP04, BPW05, BOL02, BGCP05, Ber05, Bor04, BRZRS03, Cao01a, CNY02, CLM05, DR03, DM04, Des01, Erv09, EMAS01, GR04, GP03, GPHA04, GK01, Jer08, KMR09, LS01b, Mas02, PC09, SS00, SK03, VV03, WP06, ZOS00]. Szeg [KR09]. Szego [Not06].

Tadmor [PT06]. Tangent [MZ01, Sim00]. Taylor [BJ00, EvP04]. Tchebycheff [CLM05]. technique [KW08b, KW08a, PR01, ZZ03]. techniques [BC04, BRZRS03]. temperature [TT05]. temperature-dependent [TT05]. tension [Mar01]. tensor [HSS08a, Usc08]. tent [CDLP07]. term [LP09a, RAVS07]. termination [QQ03, ZX02]. terms [ABR06]. tests [BFLW09]. tetrahedra [KP08]. tetrahedral [KV00, Kun00]. their [Gro09, GLRS09]. theorem [BGM01]. theorems [Son02, Tsu03]. Theoretical [LS08, AD01, BKW07]. theories [XZ03]. Theory [BB06b, HS06, AG00a, AXY02, BFNS01, BMVX07, BLN04b, Car05, CH07, CGJ09, Fou09, GHS00b, Kim05b, LR09, XG00]. thermal [TT05]. thermoviscoelastic [Cop08]. Theta [Bor04]. Theta-schemes [Bor04]. thickener [BKRT04]. thin [GR00, Joh00]. thin-plate [Joh00]. third [KP01b]. third-order [KP01b]. three [AG00b, Ber03, BH03c, Fae02, GS07, HHW00, Heu98, Heu01b, Tos00, TS03]. three-dimensional [BH03c, Fae02]. three-fields [Ber03]. three-step [GS07]. Tight [LP01]. Tikhonov [LP09b, MPRS01, RT06]. Tikhonov-based [RT06]. Time [MST06b, WP06, ZCW06, AGS09, Bar05, BCP07, BCGJ07, BF07, CCC^{+}05, CFA06, CFJ03a, CFJ03b, DY09, DF07, GGJ03, GK02, HL04b, HPSS05, HWW02, MST06a, MW04, MW06, Mau06, Pro00, RG04, SW08, SW01, TZ03, Zha03]. Time-delayed [WP06]. Time-dependent [ZCW06, AGS09, Bar05, BCP07, BF07, Zha03]. time-discretized [MST06a]. time-harmonic [HPSS05]. time-splitting [Pro00]. time-stepping [Mau06, SW08]. times [HL04a]. Tocher [AS02, LS01a]. Todd [Bau07]. Toeplitz [BN03, BGM01, Cap02, CNY02, LZ09, NCV06, SC07, dFP05]. Toeplitz-like-plus-diagonal [BN03]. tomographic [RS07]. tomography [BHV03, RS04]. topological [BFLW09, SŻ05]. torus [SA07]. Total [PW05a, FvOP05, PS02a, PS02b]. tracking [Str09]. transfer [BB00, BC07b, GT04, YW06]. transform [HS05, SL02]. transformation [CDLP07, Li01, MST06b, Sid05, VVVM06b]. transformations [VVVM06a]. transformed [CCSS09, Hag00]. Transmission [BHvPS03, BG01, BGG04]. transport [Des08]. Transversal [CKP07]. Trapezoidal [mWS05]. trapping [Jer08]. tree [BD04, LHJB07]. trial [ST05]. triangle [AS02]. triangles [MT05]. triangular [BS06b, KMR09, KV00, Lyl00]. triangulations [LS01a]. tridiagonal [FP09, LZ09, Tsu03, VVVM06a, dFP05]. trigonometric [MCS09b]. trust [WY06]. tuning [KPS03]. turbulent [BRM^{+}04]. turning [TMO04, WW03]. Two [BG07, BW05, BW04, FMS07, HZ02, LM05, RSS06, AGS09, Bab00, BGS02, BBC^{+}05, BRM^{+}04, Bou01, BCLS08, Cao01a, Cap02, CWRR02, CE01, ER09, FML00, GHM09, GH04, HKR01, McS09a, Mar01, Moo01, Put02, ST05]. Two-Dimensional [LM05, BCLS08, ER09, GHM09, HKR01, ST05]. two-fold [BGS02]. Two-frequency [BW04]. Two-grid [BG07, AGS09, Cap02, CWRR02]. Two-level [BW05, FML00]. two-phase [BBC^{+}05, CE01]. two-point [Mar01, Moo01]. Two-scale [FMS07, RSS06]. two-stage [Cao01a]. type [AKK^{+}08, ANS01, Bar05, BG09, CCM04, CN00, CLM03, Da08, DV00, EK02, Gal03, GT07, Gri05b, HN05, JT09, KO00, MFR03, MST06b, Med01, McS02, Not06, SZ03, SZ09, WcSX07, Xia07, ZX02, ZZL06, Gri05c].

Ulm [EH08]. Ultraconvergence [ZL03]. unbounded [BH03a]. uncertainties [BC03a]. uncertainty [BNT05]. unconstrained [DFLZ03, PC09, WY06]. Uniform [BCR04, MW04, MW06, Mic02, OPR06, TZ03, WW02, dV04, Ste05]. uniformly [BCGJ07]. unifying [Car05, CH07, CGJ09]. unilateral [AG00a]. Uniqueness [BP03, MC04]. unit [DGV02, Joh00, SL02, dlCYGV07]. units [BKRT04]. unity [DDP07]. univariate [Gro09, Peñ02, Peñ06, PR01]. unknown [HXY08, PW05b]. unsteady [WL02]. unstructured [HKR01, HCC09, KMS04, LYC03, SPSF00]. unusual [BV02]. Updating [Str08]. Upper [PS00, Bab00]. upscaling [DY09]. upwind [CK08]. use [Lac00]. using [BV07, BS06b, CDLP07, Fis02, GHL^{+}05, HLPS05, HS05, Kim05b, LLW03, Mar01, NTW01, PS02a, San04, SS00]. Uzawa [HZ02, KO00, NP04].

validation [ACP02]. value [AM02, ADM04, Cao06a, GNP08, HHW00, Koc05, KL03, KS02c, LW00, Mar01, Moo01, Stö01, BC03a]. valued [Bak02]. values [BMV09, MP02, VVM04]. Vandermonde [Bec00, EPS01]. Variable [BLM05, CF08, DH06, GZ01, Sid05]. variables [CGS06, Put02]. variants [HZ02]. variation [FvOP05]. variational [AP01, BGN08, BR04, CLRM01, CR00, CCF08, GPC05, HXY08, HLW03, KS02b, Kor02, KK00, KS07, MSTZ03b, Mül03, PC09, Sim00, Tai03]. various [GJ02]. vector [AP01, Ber05, Bou03]. vectorial [BB08]. vectors [DV05a]. velocity [ABC06, Deu09]. verification [KNWN09, OR02, TMO04]. version [AG00b, Bab00, BH05a, BH07, CS00a, GH04, Heu98, Heu01a, Heu01b, KS07, SW08, TS00]. vertex [CK02]. vertex-centered [CK02]. via [AK04, GMZ02, HS08, MST06b, PK05, TZ03, VVM04, WW02]. vibration [DHNL^{+}00]. vibrations [BRS00]. virtual [GLQ01]. viscoelastic [AB03, BCP07, RAVS07]. viscoelasticity [RSWW03, SW01]. Viscoplastic [FHV03, FGSn02]. viscosity [BLN04b, TZ03]. Vlasov [BC09, PM08]. volatility [IT08]. Volterra [Da08, SW00, SW01, Vec01]. volume [AJMN02, ABC03, ABH05, AW07, BMV03, BG07, BG09, CHC01, Des08, DE06, EGHM02, EHV06, EFG06, Fil06, HCC09, HMS03, KMR09, Kim05a, MV07, MR01, NAE07, Ohl01, VV03, Voh08, Vov02]. volume-nonconforming [EHV06]. volume-nonconforming/mixed-hybrid [EHV06]. volumes [BMV03]. vorticity [AB03, ABC06, DSS02, Gal02, WL02].

Washizu [LRW06]. water [DMC00, PMC01, VK09]. Wave [CFJ03a, CFJ03b, DY09, Can06, CM06, CHL08, Erv09, MM07, Mic02, OWW04, TZ03]. wavefunctions [Yse07]. Wavelet [Ber00, CM00, BDS07, CS07, DSX00, HPPS02]. wavelets [JJL02]. Waves [GHNP03, EZ08, LS08, Weg02]. weak [CLM05, Kwe00]. weakly [BH07, DK04, FMS07]. web [HRW02]. web-splines [HRW02]. Weight [PP03, But04, DGV02, MC04, MC06, Not06]. Weighted [BBD06, CD02, LSW02, BSS04, BSS05, DD02, DSWW06]. weights [DSWW06]. Well [BKRT04, LMV02, Usc08, CCP07, GT04]. well-balanced [GT04]. Well-posed [LMV02, CCP07]. Well-posedness [BKRT04, Usc08]. Which [Maz08b]. whole [Zhe06]. Willmore [Dzi08]. withdrawal [CF08]. without [AM02, BC07a, JKN09, VVM09]. Worst [BNT05, Hes06]. worst-case [Hes06].

Yamamoto [Tsu03]. yield [BC04]. Young [CR00, KP01a]. Yuan [Pyt02].

zero [CNY02, GZ01, VK09]. zero-stability [GZ01]. Zeros [VC01, DJ07, Pas00]. Zienkiewicz [WcSX07]. Zienkiewicz-type [WcSX07].

References

Allen:2002:NEA

[AB02] Leanne Allen and Thomas J. Bridges. Numerical exterior algebra and the compound matrix method. Numerische Mathematik, 92 (2):197–232, August 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amara:2003:ESV

[AB03] Mohamed Amara and Jacques Baranger. An extra stress-vorticity formulation of Stokes problem for the Oldroyd viscoelastic model. Numerische Mathematik, 94(4):603–622, June 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

           http://www.springerlink.com/openurl.asp?genre=article&
           issn=0029-599X&volume=94&issue=4&spage=603.

Achdou:2003:PPA

[ABC03] Y. Achdou, C. Bernardi, and F. Coquel. A priori and a posteriori analysis of finite volume discretizations of Darcy’s equations. Numerische Mathematik, 96(1):17–42, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Azaiez:2006:SDV

[ABC06] Mejdi Azaïez, Christine Bernardi, and Nejmeddine Chorfi. Spectral discretization of the vorticity, velocity and pressure formulation of the Navier–Stokes equations. Numerische Mathematik, 104 (1):1–2, July 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Andreianov:2005:BRN

[ABH05] Boris Andreianov, Franck Boyer, and Florence Hubert. Besov regularity and new error estimates for finite volume approximations of the p-Laplacian. Numerische Mathematik, 100 (4):565–592, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=100&issue=4&spage=565.

Araya:2006:PEE

[ABR06] Rodolfo Araya, Edwin Behrens, and Rodolfo Rodríguez. A posteriori error estimates for elliptic problems with Dirac delta source terms. Numerische Mathematik, 105(2):193–216, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Allaire:2001:NAO

[AC01] Grégoire Allaire and Carlos Castro. A new approach for the optimal distribution of assemblies in a nuclear reactor. Numerische Mathematik, 89(1):1–29, July 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1089001/10890001.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089001/10890001.pdf.

Albarreal:2002:CAE

[ACC^{+}02] I. Albarreal, M. C. Calzada, J. L. Cruz, E. Fernández-Cara, J. Galo, and M. Marín. Convergence analysis and error estimates for a parallel algorithm for solving the Navier–Stokes equations. Numerische Mathematik, 93(2):201–221, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alvarez:2006:CSN

[ACMR06] O. Alvarez, E. Carlini, R. Monneau, and E. Rouy. A convergent scheme for a non local Hamilton–Jacobi equation modelling dislocation dynamics. Numerische Mathematik, 104(4):413–444, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:2002:NVS

[ACP02] G. E. Alefeld, X. Chen, and F. A. Potra. Numerical validation of solutions of complementarity: The nonlinear case. Numerische Mathematik, 92(1):1–16, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aslanyan:2000:SIS

[AD00] A. Aslanyan and E. B. Davies. Spectral instability for some Schrödinger operators. Numerische Mathematik, 85(4):525–552, June 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085004/00850525.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085004/00850525.pdf.

Amara:2001:OCF

[AD01] M. Amara and F. Dabaghi. An optimal C^{0} finite element algorithm for the 2D biharmonic problem: theoretical analysis and numerical results. Numerische Mathematik, 90(1):19–46, November 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090001/10900019.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090001/10900019.pdf.

Arico:2007:CMM

[AD07] Antonio Aricò and Marco Donatelli. A V -cycle multigrid for multilevel matrix algebras: proof of optimality. Numerische Mathematik, 105(4):511–547, February 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aregba-Driollet:2004:KAB

[ADM04] Denise Aregba-Driollet and Vuk Milišić. Kinetic approximation of a boundary value problem for conservation laws. Numerische Mathematik, 97(4):595–633, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arcangeli:2005:EEW

[AdST05] Rémi Arcangéli, María Cruz López de Silanes, and Juan José Torrens. Error estimates in W^{2,∞}-semi-norms for discrete interpolating D^{2}-splines. Numerische Mathematik, 101(4): 573–599, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=4&spage=573.

Arcangeli:2007:EBF

[AdST07] Rémi Arcangéli, Mara Cruz López de Silanes, and Juan José Torrens. An extension of a bound for functions in Sobolev spaces, with applications to (m, s)-spline interpolation and smoothing. Numerische Mathematik, 107(2):181–211, August 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:2000:MHD

[AFW00] Douglas N. Arnold, Richard S. Falk, and Ragnar Winther. Multigrid in H (div) and H (curl). Numerische Mathematik, 85(2):197– 217, April 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085002/00850197.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085002/00850197.pdf.

Adly:2000:DTC

[AG00a] Samir Adly and Daniel Goeleven. A discretization theory for a class of semi-coercive unilateral problems. Numerische Mathematik, 87(1):1–34, November 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0087001/00870001.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0087001/00870001.pdf.

Ainsworth:2000:ASP

[AG00b] Mark Ainsworth and Benqi Guo. An additive Schwarz preconditioner for p-version boundary element approximation of the hypersingular operator in three dimensions. Numerische Mathematik, 85(3):343–366, May 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0085003/00850343.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0085003/00850343.pdf.

Axelsson:2001:SEE

[AG01] O. Axelsson and S. V. Gololobov. Stability and error estimates for the θ-method for strongly monotone and infinitely stiff evolution equations. Numerische Mathematik, 89(1):31– 48, July 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089001/10890031.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089001/10890031.pdf.

Abboud:2009:SOA

[AGS09] Hyam Abboud, Vivette Girault, and Toni Sayah. A second order accuracy for a full discretized time-dependent Navier–Stokes equations by a two-grid scheme. Numerische Mathematik, 114(2): 189–231, December 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=114&issue=2&spage=189.

Achdou:2002:NCG

[AJMN02] Y. Achdou, C. Japhet, Y. Maday, and F. Nataf. A new cement to glue non-conforming grids with Robin interface conditions: The finite volume case. Numerische Mathematik, 92(4):593–620, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:2004:SCC

[AK04] O. Axelsson and J. Karátson. Superlinearly convergent CG methods via equivalent preconditioning for nonsymmetric elliptic operators. Numerische Mathematik, 99(2):197–223, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Akrivis:2003:LIM

[AKK03] Georgios Akrivis, Ohannes Karakashian, and Fotini Karakatsani. Linearly implicit methods for nonlinear evolution equations. Numerische Mathematik, 94(3):403–418, May 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ammari:2008:MTA

[AKK^{+}08] Habib Ammari, Hyeonbae Kang, Eunjoo Kim, Kaouthar Louati, and Michael S. Vogelius. A MUSIC-type algorithm for detecting internal corrosion from electrostatic boundary measurements. Numerische Mathematik, 108(4):501–528, February 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aldaz:2009:SPP

[AKR09] J. M. Aldaz, O. Kounchev, and H. Render. Shape preserving properties of generalized Bernstein operators on Extended Chebyshev spaces. Numerische Mathematik, 114(1):1–25, November 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=114&issue=1&spage=1.

Arioli:2005:SCI

[ALW05] M. Arioli, D. Loghin, and A. J. Wathen. Stopping criteria for iterations in finite element methods. Numerische Mathematik, 99 (3):381–410, January 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Abdulle:2001:SOC

[AM01] Assyr Abdulle and Alexei A. Medovikov. Second order Chebyshev methods based on orthogonal polynomials. Numerische Mathematik, 90(1):1–18, November 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1090001/10900001.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1090001/10900001.pdf.

Alonso-Mallo:2002:RKM

[AM02] Isaías Alonso-Mallo. Runge–Kutta methods without order reduction for linear initial boundary value problems. Numerische Mathematik, 91(4):577–603, June 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ainsworth:2003:MDS

[AM03] Mark Ainsworth and William McLean. Multilevel diagonal scaling preconditioners for boundary element equations on locally refined meshes. Numerische Mathematik, 93(3):387–413, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Acosta:2006:IEE

[AM06] Gabriel Acosta and Gabriel Monzón. Interpolation error estimates in W^{1,p} for degenerate Q_{1} isoparametric elements. Numerische Mathematik, 104(2):129–150, August 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aguilera:2008:AOP

[AM08] Néstor E. Aguilera and Pedro Morin. Approximating optimization problems over convex functions. Numerische Mathematik, 111(1): 1–34, November 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=1&spage=1.

Akrivis:2009:OOP

[AMN09] Georgios Akrivis, Charalambos Makridakis, and Ricardo H. Nochetto. Optimal order a posteriori error estimates for a class of Runge–Kutta and Galerkin methods. Numerische Mathematik, 114(1):133–160, November 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=114&issue=1&spage=
133.

Anthonissen:2003:CAL

[AMtTB03] M. J. H. Anthonissen, R. M. M. Mattheij, and J. H. M. ten Thije Boonkkamp. Convergence analysis of the local defect correction method for diffusion equations. Numerische Mathematik, 95(3):401–425, September 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Angermann:2000:FEM

[Ang00] Lutz Angermann. A finite element method for the numerical solution of convection-dominated anisotropic diffusion equations. Numerische Mathematik, 85(2):175–195, April 2000. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0085002/00850175.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0085002/
00850175.pdf.

Apel:2001:CRT

[ANS01] Thomas Apel, Serge Nicaise, and Joachim Schöberl. Crouzeix– Raviart type finite elements on anisotropic meshes. Numerische Mathematik, 89(2):193–223, August 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/

1089002/10890193.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089002/10890193.pdf.

Aranda:2001:NAN

[AP01] Ernesto Aranda and Pablo Pedregal. Numerical approximation of non-homogeneous, non-convex vector variational problems. Numerische Mathematik, 89(3):425–444, September 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089003/10890425.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089003/10890425.pdf.

Alouges:2005:MOS

[AP05] François Alouges and Morgan Pierre. Mesh optimization for singular axisymmetric harmonic maps from the disc into the sphere. Numerische Mathematik, 101(3):391–414, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=101&issue=3&spage=391.

Arioli:2004:SCC

[Ari04] M. Arioli. A stopping criterion for the conjugate gradient algorithm in a finite element method framework. Numerische Mathematik, 97(1):1–24, March 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alonso:2004:APP

[ARPR04] Ana Alonso, Anahí Dello Russo, Claudio Padra, and Rodolfo Rodríguez. Accurate pressure post-process of a finite element method for elastoacoustics. Numerische Mathematik, 98(3):389– 425, September 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alfeld:2002:SME

[AS02] Peter Alfeld and Larry L. Schumaker. Smooth macro-elements based on Clough–Tocher triangle splits. Numerische Mathematik, 90(4):597–616, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.de/link/service/journals/00211/bibs/2090004/
20900597.htm;   http://link.springer.de/link/service/
journals/00211/papers/2090004/20900597.pdf.

Achdou:2001:FDS

[AT01] Yves Achdou and Nicoletta Tchou. A finite difference scheme on a non commutative group. Numerische Mathematik, 89(3):401–424, September 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089003/10890401.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089003/10890401.pdf.

Aceto:2003:PMD

[AT03] Lidia Aceto and Donato Trigiante. On the properties of matrices defining some classes of BVMs. Numerische Mathematik, 96(1):1– 16, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:2002:MFE

[AW02] Douglas N. Arnold and Ragnar Winther. Mixed finite elements for elasticity. Numerische Mathematik, 92(3):401–419, September 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Angermann:2007:CFF

[AW07] Lutz Angermann and Song Wang. Convergence of a fitted finite volume method for the penalized Black–Scholes equation governing European and American option pricing. Numerische Mathematik, 106(1):1–40, March 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alfa:2002:EPT

[AXY02] Attahiru Sule Alfa, Jungong Xue, and Qiang Ye. Entrywise perturbation theory for diagonally dominant M-matrices with applications. Numerische Mathematik, 90(3):401–414, January 2002. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/2090003/20900401.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/2090003/20900401.pdf.

Babuska:2000:OEL

[Bab00] Ivo Babuška. Optimal estimates for lower and upper bounds of approximation errors in the p-version of the finite element method in two dimensions. Numerische Mathematik, 85(2):219– 255, April 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085002/00850219.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085002/00850219.pdf.

Bakaev:2002:AFD

[Bak02] Nikolai Yu. Bakaev. Analysis of fully discrete approximations to parabolic problems with rough or distribution-valued initial data. Numerische Mathematik, 92(4):621–651, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ben-Artzi:2007:HBL

[BAL07] Matania Ben-Artzi and Jiequan Li. Hyperbolic balance laws: Riemann invariants and the generalized Riemann problem. Numerische Mathematik, 106(3):369–425, May 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bansch:2001:FED

[Bän01] Eberhard Bänsch. Finite element discretization of the Navier– Stokes equations with a free capillary surface. Numerische Mathematik, 88(2):203–235, April 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088002/10880203.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088002/10880203.pdf.

Bartels:2005:PEA

[Bar05] Sören Bartels. A posteriori error analysis for time-dependent Ginzburg–Landau type equations. Numerische Mathematik, 99 (4):557–583, February 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:2007:MJT

[Bau07] Friedrich L. Bauer. In memoriam John Todd (1911–2007). Numerische Mathematik, 108(1):1–6, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Benamou:2000:CFM

[BB00] Jean-David Benamou and Yann Brenier. A computational fluid mechanics solution to the Monge–Kantorovich mass transfer problem. Numerische Mathematik, 84(3):375–393, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0084003/00840375.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0084003/00840375.pdf.

Barrett:2001:IEB

[BB01] John W. Barrett and James F. Blowey. An improved error bound for a finite element approximation of a model for phase separation of a multi-component alloy with a concentration dependent mobility matrix. Numerische Mathematik, 88(2):255– 297, April 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088002/10880255.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088002/10880255.pdf.

Barrett:2004:FEAb

[BB04] John W. Barrett and James F. Blowey. Finite element approximation of a nonlinear cross-diffusion population model. Numerische Mathematik, 98(2):195–221, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bendali:2006:NOD

[BB06a] Abderrahmane Bendali and Yassine Boubendir. Non-overlapping domain decomposition method for a nodal finite element method. Numerische Mathematik, 103(4):515–537, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Benner:2006:AMP

[BB06b] Peter Benner and Ralph Byers. An arithmetic for matrix pencils: Theory and new algorithms. Numerische Mathematik, 103(4):539– 573, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Benbourhim:2008:PPV

[BB08] Mohammed-Najib Benbourhim and Abderrahman Bouhamidi. Pseudo-polyharmonic vectorial approximation for div-curl and elastic semi-norms. Numerische Mathematik, 109(3):333–364, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baudin:2005:RMT

[BBC^{+}05] Michaël Baudin, Christophe Berthon, Frédéric Coquel, Roland Masson, and Quang Huy Tran. A relaxation method for twophase flow models with hydrodynamic closure law. Numerische Mathematik, 99(3):411–440, January 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

BenBelgacem:2000:ISC

[BBCM00] F. Ben Belgacem, C. Bernardi, N. Chorfi, and Y. Maday. Infsup conditions for the mortar spectral element discretization of the Stokes problem. Numerische Mathematik, 85(2):257– 281, April 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085002/00850257.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085002/00850257.pdf.

Belhachmi:2006:WCO

[BBD06] Zakaria Belhachmi, Christine Bernardi, and Simone Deparis. Weighted Clément operator and application to the finite element discretization of the axisymmetric Stokes problem. Numerische Mathematik, 105(2):217–247, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:2004:ASP

[BBO04] Ivo Babuška, Uday Banerjee, and John E. Osborn. On the approximability and the selection of particle shape functions. Numerische Mathematik, 96(4):601–640, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:2007:SGF

[BBO07] Ivo Babuška, Uday Banerjee, and John E. Osborn. Superconvergence in the generalized finite element method. Numerische Mathematik, 107(3):353–395, September 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bachelot:2001:CEF

[BBP01] Alain Bachelot, Laurent Bounhoure, and Agnès Pujols. Couplage éléments finis-potentiels retardés pour la diffraction électromagnétique par un obstacle hétérogène. Numerische Mathematik, 89(2):257–306, August 2001. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1089002/10890257.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089002/10890257.pdf.

Bertoluzza:2007:MMN

[BBS07] S. Bertoluzza, F. Brezzi, and G. Sangalli. The method of mothers for non-overlapping non-matching DDM. Numerische Mathematik, 107(3):397–431, September 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bourdin:2000:IAF

[BC00] Blaise Bourdin and Antonin Chambolle. Implementation of an adaptive finite-element approximation of the Mumford– Shah functional. Numerische Mathematik, 85(4):609–646, June 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085004/00850609.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085004/00850609.pdf.

ka:2003:EUD

[BC03a] Ivo Babuška and Jan Chleboun. Effects of uncertainties in the domain on the solution of Dirichlet boundary value problems. Numerische Mathematik, 93(4):583–610, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buffa:2003:EFI

[BC03b] A. Buffa and S. H. Christiansen. The electric field integral equation on Lipschitz screens: definitions and numerical approximation. Numerische Mathematik, 94(2):229–267, April 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:2004:ATY

[BC04] S. Bartels and C. Carstensen. Averaging techniques yield reliable a posteriori finite element error control for obstacle problems. Numerische Mathematik, 99(2):225–249, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Badia:2007:CAF

[BC07a] Santiago Badia and Ramon Codina. Convergence analysis of the FEM approximation of the first order projection method for incompressible flows with and without the inf-sup condition. Numerische Mathematik, 107(4):533–557, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buet:2007:APS

[BC07b] Christophe Buet and Stéphane Cordier. An asymptotic preserving scheme for hydrodynamics radiative transfer models. Numerische Mathematik, 108(2):199–221, December 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:2008:CAF

[BC08] Sören Bartels and Carsten Carstensen. A convergent adaptive finite element method for an optimal design problem. Numerische Mathematik, 108(3):359–385, January 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bostan:2009:CSL

[BC09] Mihai Bostan and Nicolas Crouseilles. Convergence of a semi-Lagrangian scheme for the reduced Vlasov Maxwell system for laser plasma interaction. Numerische Mathematik, 112(2):169– 195, April 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=2&spage=169.

Bartels:2004:IDC

[BCD04] S. Bartels, C. Carstensen, and G. Dolzmann. Inhomogeneous Dirichlet conditions in a priori and a posteriori finite element error analysis. Numerische Mathematik, 99(1):1–24, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buffa:2005:ACA

[BCD05] A. Buffa, M. Costabel, and M. Dauge. Algebraic convergence for anisotropic edge elements in polyhedral domains. Numerische Mathematik, 101(1):29–65, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=1&spage=29.

Boffi:2006:MEF

[BCG06] Daniele Boffi, Matteo Conforti, and Lucia Gastaldi. Modified edge finite elements for photonic crystals. Numerische Mathematik, 105 (2):249–266, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bujanda:2007:HOU

[BCGJ07] B. Bujanda, C. Clavero, J. L. Gracia, and J. C. Jorge. A high order uniformly convergent alternating direction scheme for time dependent reaction-diffusion singularly perturbed problems. Numerische Mathematik, 107(1):1–25, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:2007:CAC

[BCH07] Dietrich Braess, Carsten Carstensen, and Ronald H. W. Hoppe. Convergence analysis of a conforming adaptive finite element method for an obstacle problem. Numerische Mathematik, 107(3): 455–471, September 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buffa:2009:SEE

[BCJ09] Annalisa Buffa, Patrick Ciarlet, and Erell Jamelot. Solving electromagnetic eigenvalue problems in polyhedral domains with nodal finite elements. Numerische Mathematik, 113(4):497–518, October 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=4&spage=497.

Brenner:2008:NFE

[BCLS08] S. C. Brenner, J. Cui, F. Li, and L.-Y. Sung. A nonconforming finite element method for a two-dimensional curl-curl and grad-div problem. Numerische Mathematik, 109(4):509–533, June 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buades:2006:NFP

[BCM06] Antoni Buades, Bartomeu Coll, and Jean-Michel Morel. Neighborhood filters and PDEs. Numerische Mathematik, 105(1):1–34, November 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bonito:2007:MNA

[BCP07] Andrea Bonito, Philippe Clément, and Marco Picasso. Mathematical and numerical analysis of a simplified time-dependent viscoelastic flow. Numerische Mathematik, 107(2):213–255, August 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:2004:UCP

[BCR04] D. Braess, C. Carstensen, and B. D. Reddy. Uniform convergence and a posteriori error estimators for the enhanced strain finite element method. Numerische Mathematik, 96(3):461–479, January 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buffa:2002:BEM

[BCS02] A. Buffa, M. Costabel, and C. Schwab. Boundary element methods for Maxwell’s equations on non-smooth domains. Numerische Mathematik, 92(4):679–710, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beska:2001:AFE

[BD01] M. Beśka and K. Dziedziul. Asymptotic formula for the error in cardinal interpolation. Numerische Mathematik, 89(3):445–456, September 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089003/10890445.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089003/10890445.pdf.

Bulgak:2002:ERA

[BD02] H. Bulgak and G. Demidenko. Estimation for the region of attraction of nonlinear difference equations. Numerische Mathematik, 92(3):421–431, September 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Binev:2004:FCA

[BD04] Peter Binev and Ronald DeVore. Fast computation in adaptive tree approximation. Numerische Mathematik, 97(2):193–217, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Binev:2004:AFE

[BDD04] Peter Binev, Wolfgang Dahmen, and Ron DeVore. Adaptive finite element methods with convergence rates. Numerische Mathematik, 97(2):219–268, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bjorstad:2003:ASM

[BDR03] Petter E. Bjørstad, Maksymilian Dryja, and Talal Rahman. Additive Schwarz methods for elliptic mortar finite element problems. Numerische Mathematik, 95(3):427–457, September 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barinka:2007:FCA

[BDS07] A. Barinka, W. Dahmen, and R. Schneider. Fast computation of adaptive wavelet expansions. Numerische Mathematik, 105(4): 549–589, February 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bebendorf:2000:ABE

[Beb00] Mario Bebendorf. Approximation of boundary element matrices. Numerische Mathematik, 86(4):565–589, October 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0086004/00860565.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0086004/
00860565.pdf.

Beckermann:2000:CNR

[Bec00] Bernhard Beckermann. The condition number of real Vandermonde, Krylov and positive definite Hankel matrices. Numerische Mathematik, 85(4):553–577, June 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085004/00850553.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085004/00850553.pdf.

Bertoluzza:2000:WSL

[Ber00] Silvia Bertoluzza. Wavelet stabilization of the Lagrange multiplier method. Numerische Mathematik, 86(1):1–28, July 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0086001/00860001.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0086001/
00860001.pdf.

Bermejo:2001:ACQ

[Ber01] Rodolfo Bermejo. Analysis of a class of quasi-monotone and conservative semi-Lagrangian advection schemes. Numerische Mathematik, 87(4):597–623, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087004/10870597.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087004/10870597.pdf.

Berrone:2002:RPE

[Ber02] Stefano Berrone. Robustness in a posteriori error analysis for FEM flow models. Numerische Mathematik, 91(3):389–422, May 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=91&issue=3&spage=
389.

Bertoluzza:2003:AST

[Ber03] Silvia Bertoluzza. Analysis of a stabilized three-fields domain decomposition method. Numerische Mathematik, 93(4):611–634, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berthelin:2005:CFV

[Ber05] F. Berthelin. Convergence of flux vector splitting schemes with single entropy inequality for hyperbolic systems of conservation laws. Numerische Mathematik, 99(4):585–604, February 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berthon:2006:WMS

[Ber06] Christophe Berthon. Why the MUSCL–Hancock scheme is L^{1}- stable. Numerische Mathematik, 104(1):27–46, July 2006. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berrut:2009:FAF

[Ber09] Jean-Paul Berrut. First applications of a formula for the error of finite sinc interpolation. Numerische Mathematik, 112(3):341– 361, May 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=3&spage=341.

Bey:2000:SGR

[Bey00] Jürgen Bey. Simplicial grid refinement: on Freudenthal’s algorithm and the optimal number of congruence classes. Numerische Mathematik, 85(1):1–29, March 2000. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085001/00850001.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085001/00850001.pdf.

Brezzi:2001:MSP

[BF01] F. Brezzi and M. Fortin. A minimal stabilisation procedure for mixed finite element methods. Numerische Mathematik, 89(3): 457–491, September 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1089003/10890457.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1089003/10890457.pdf.

Burman:2007:CIP

[BF07] Erik Burman and Miguel A. Fernández. Continuous interior penalty finite element method for the time-dependent Navier–Stokes equations: space discretization and convergence. Numerische Mathematik, 107(1):39–77, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beelitz:2009:FET

[BFLW09] Thomas Beelitz, Andreas Frommer, Bruno Lang, and Paul Willems. A framework for existence tests based on the topological degree and homotopy. Numerische Mathematik, 111(4): 493–507, February 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=4&spage=493.

Benzi:2001:ATM

[BFNS01] Michele Benzi, Andreas Frommer, Reinhard Nabben, and Daniel B. Szyld. Algebraic theory of multiplicative Schwarz methods. Numerische Mathematik, 89(4):605–639, October 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1089004/10890605.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1089004/
10890605.pdf.

Babuska:2001:OAP

[BFS01] Ivo Babuška, Miloslav Feistauer, and Pavel Solín. On one approach to a posteriori error estimates for evolution problems solved by the method of lines. Numerische Mathematik, 89(2):225–256, August 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089002/10890225.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089002/10890225.pdf.

Barrenechea:2001:PMF

[BG01] Gabriel R. Barrenechea and Gabriel N. Gatica. A primal mixed formulation for exterior transmission problems in R^{2}. Numerische Mathematik, 88(2):237–253, April 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088002/10880237.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088002/10880237.pdf.

Borm:2005:HCA

[BG05] Steffen Börm and Lars Grasedyck. Hybrid cross approximation of integral operators. Numerische Mathematik, 101(2): 221–249, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=2&spage=221.

Bi:2007:TGF

[BG07] Chunjia Bi and Victor Ginting. Two-grid finite volume element method for linear and nonlinear elliptic problems. Numerische Mathematik, 108(2):177–198, December 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bi:2009:RTP

[BG09] Chunjia Bi and Victor Ginting. A residual-type a posteriori error estimate of finite volume element method for a quasilinear elliptic problem. Numerische Mathematik, 114(1):107–132, November 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=114&issue=1&spage=107.

Bertaccini:2005:PHM

[BGCP05] Daniele Bertaccini, Gene H. Golub, Stefano Serra Capizzano, and Cristina Tablino Possio. Preconditioned HSS methods for the solution of non-hermitian positive definite linear systems and applications to the discrete convection-diffusion equation. Numerische Mathematik, 99(3):441–484, January 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bustinza:2004:MFE

[BGG04] Rommel Bustinza, Galina C. Garcia, and Gabriel N. Gatica. A mixed finite element method with Lagrange multipliers for nonlinear exterior transmission problems. Numerische Mathematik, 96 (3):481–523, January 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bermudez:2003:FEA

[BGHNR03] A. Bermúdez, P. Gamallo, L. Hervella-Nieto, and R. Rodríguez. Finite element analysis of pressure formulation of the elastoacoustic problem. Numerische Mathematik, 95(1):29–51, July 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beckermann:2007:NCG

[BGL07] B. Beckermann, G. H. Golub, and G. Labahn. On the numerical condition of a generalized Hankel eigenvalue problem. Numerische Mathematik, 106(1):41–68, March 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bini:2001:FAF

[BGM01] Dario Andrea Bini, Luca Gemignani, and Beatrice Meini. Factorization of analytic functions by means of Koenig’s theorem and Toeplitz computations. Numerische Mathematik, 89(1):49– 82, July 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089001/10890049.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089001/10890049.pdf.

Barrett:2008:VFA

[BGN08] John W. Barrett, Harald Garcke, and Robert Nürnberg. A variational formulation of anisotropic geometric evolution equations in higher dimensions. Numerische Mathematik, 109(1):1–44, March 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bai:2004:PHS

[BGP04] Zhong-Zhi Bai, Gene H. Golub, and Jian-Yu Pan. Preconditioned Hermitian and skew-Hermitian splitting methods for non-Hermitian positive semidefinite linear systems. Numerische Mathematik, 98(1):1–32, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bini:2005:FSQ

[BGP05] Dario A. Bini, Luca Gemignani, and Victor Y. Pan. Fast and stable QR eigenvalue algorithms for generalized companion matrices and secular equations. Numerische Mathematik, 100(3):373–408, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrientos:2002:MFE

[BGS02] Mauricio A. Barrientos, Gabriel N. Gatica, and Ernst P. Stephan. A mixed finite element method for nonlinear elasticity: two-fold saddle point approach and a-posteriori error estimate. Numerische Mathematik, 91(2):197–222, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bao:2003:EBF

[BH03a] Weizhu Bao and Houde Han. Error bounds for the finite element approximation of an incompressible material in an unbounded domain. Numerische Mathematik, 93(3):415–444, January 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bebendorf:2003:EHM

[BH03b] Mario Bebendorf and Wolfgang Hackbusch. Existence of H-matrix approximants to the inverse F E-matrix of elliptic operators with L^{8}-coefficients. Numerische Mathematik, 95(1):1–28, July 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brenner:2003:LBT

[BH03c] Susanne C. Brenner and Qingmi He. Lower bounds for threedimensional nonoverlapping domain decomposition algorithms. Numerische Mathematik, 93(3):445–470, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beyn:2004:EEA

[BH04] Wolf-Jürgen Beyn and Thorsten Hüls. Error estimates for approximating non-hyperbolic heteroclinic orbits of maps. Numerische Mathematik, 99(2):289–323, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bespalov:2005:VBE

[BH05a] Alexei Bespalov and Norbert Heuer. The p-version of the boundary element method for hypersingular operators on piecewise plane open surfaces. Numerische Mathematik, 100(2):185–209, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buffa:2005:RCF

[BH05b] A. Buffa and R. Hiptmair. Regularized combined field integral equations. Numerische Mathematik, 100(1):1–19, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bespalov:2007:VBE

[BH07] Alexei Bespalov and Norbert Heuer. The p-version of the boundary element method for weakly singular operators on piecewise plane open surfaces. Numerische Mathematik, 106(1):69–97, March 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bruhl:2003:DIT

[BHV03] Martin Brühl, Martin Hanke, and Michael S. Vogelius. A direct impedance tomography algorithm for locating small inhomogeneities. Numerische Mathematik, 93(4):635–654, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buffa:2003:BEM

[BHvPS03] A. Buffa, R. Hiptmair, T. von Petersdorff, and C. Schwab. Boundary element methods for Maxwell transmission problems in Lipschitz domains. Numerische Mathematik, 95(3):459–485, September 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buscaglia:2000:SAT

[BJ00] G. Buscaglia and M. Jai. Sensitivity analysis and Taylor expansions in numerical homogenization problems. Numerische Mathematik, 85(1):49–75, March 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085001/00850049.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085001/00850049.pdf.

Butcher:2003:NAE

[BJ03] J. C. Butcher and Z. Jackiewicz. A new approach to error estimation for general linear methods. Numerische Mathematik, 95(3): 487–502, September 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Burger:2004:WPB

[BKRT04] R. Bürger, K. H. Karlsen, N. H. Risebro, and J. D. Towers. Well-posedness in BV_{t} and convergence of a difference scheme for continuous sedimentation in ideal clarifier-thickener units. Numerische Mathematik, 97(1):25–65, March 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bouchut:2007:MAR

[BKW07] François Bouchut, Christian Klingenberg, and Knut Waagan. A multiwave approximate Riemann solver for ideal MHD based on relaxation. I: theoretical framework. Numerische Mathematik, 108 (1):7–42, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bensow:2005:DLS

[BL05] Rickard Bensow and Mats G. Larson. Discontinuous least-squares finite element method for the Div–Curl problem. Numerische Mathematik, 101(4):601–617, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=4&spage=601.

Beyn:2009:HMC

[BL09] Wolf-Jürgen Beyn and Alexander Lust. A hybrid method for computing Lyapunov exponents. Numerische Mathematik, 113(3): 357–375, September 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=113&issue=3&spage=357.

Borm:2005:AIO

[BLM05] Steffen Börm, Maike Löhndorf, and Jens M. Melenk. Approximation of integral operators by variable–order interpolation. Numerische Mathematik, 99(4):605–643, February 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:2004:FEAa

[BLN04a] John W. Barrett, Stephen Langdon, and Robert Nürnberg. Finite element approximation of a sixth order nonlinear degenerate parabolic equation. Numerische Mathematik, 96(3):401–434, January 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Briani:2004:CNS

[BLN04b] Maya Briani, Claudia La Chioma, and Roberto Natalini. Convergence of numerical schemes for viscosity solutions to integrodifferential degenerate parabolic problems arising in financial theory. Numerische Mathematik, 98(4):607–646, October 2004. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Burke:2007:SCP

[BLO07] J. V. Burke, A. S. Lewis, and M. L. Overton. Spectral conditioning and pseudospectral growth. Numerische Mathematik, 107(1):27– 37, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bachinger:2005:NAN

[BLS05a] F. Bachinger, U. Langer, and J. Schöberl. Numerical analysis of nonlinear multiharmonic eddy current problems. Numerische Mathematik, 100(4):593–616, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=100&issue=4&spage=593.

Berndt:2005:MMF

[BLS^{+}05b] Markus Berndt, Konstantin Lipnikov, Mikhail Shashkov, Mary F. Wheeler, and Ivan Yotov. A mortar mimetic finite difference method on non-matching grids. Numerische Mathematik, 102(2):203–230, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=2&spage=203.

Bonnans:2006:COC

[BLV06] J. Frédéric Bonnans and Julien Laurent-Varin. Computation of order conditions for symplectic partitioned Runge– Kutta schemes with application to optimal control. Numerische Mathematik, 103(1):1–10, March 2006. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=103&issue=1&spage=1.

Bellen:2000:NSC

[BM00] A. Bellen and S. Maset. Numerical solution of constant coefficient linear delay differential equations as abstract Cauchy problems. Numerische Mathematik, 84(3):351–374, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0084003/00840351.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0084003/00840351.pdf.

Becker:2008:OCA

[BM08] Roland Becker and Shipeng Mao. An optimally convergent adaptive mixed finite element method. Numerische Mathematik, 111 (1):35–54, November 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=1&spage=35.

Brezzi:2000:RFB

[BMS00] F. Brezzi, D. Marini, and E. Süli. Residual-free bubbles for advection-diffusion problems: the general error analysis. Numerische Mathematik, 85(1):31–47, March 2000. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0085001/00850031.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0085001/
00850031.pdf.

Bergam:2003:EPD

[BMV03] A. Bergam, Z. Mghazli, and R. Verfürth. Estimations a posteriori d’un schéma de volumes finis pour un problème non linéaire. (French) [A posteriori estimations of a finite-volume schema for a linear problem]. Numerische Mathematik, 95(4):599–624, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Breda:2009:NAC

[BMV09] D. Breda, S. Maset, and R. Vermiglio. Numerical approximation of characteristic values of partial retarded functional differential equations. Numerische Mathematik, 113(2):181–242, August 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=2&spage=181.

Bos:2007:BLI

[BMVX07] Len Bos, Stefano De Marchi, Marco Vianello, and Yuan Xu. Bivariate Lagrange interpolation at the Padua points: the ideal theory approach. Numerische Mathematik, 108(1):43–57, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bourgeois:2001:PAH

[BN01] Christian Bourgeois and Serge Nicaise. Prewavelet analysis of the heat equation. Numerische Mathematik, 87(3):407–434, January 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1087003/10870407.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1087003/10870407.pdf.

Bai:2003:PNB

[BN03] Zhong-Zhi Bai and Michael K. Ng. Preconditioners for nonsymmetric block Toeplitz-like-plus-diagonal linear systems. Numerische Mathematik, 96(2):197–220, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bompart:2007:SON

[BNA07] Vincent Bompart, Dominikus Noll, and Pierre Apkarian. Second-order nonsmooth optimization for H [] synthesis. Numerische Mathematik, 107(3):433–454, September 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:2005:WCS

[BNT05] I. Babuška, F. Nobile, and R. Tempone. Worst case scenario analysis for elliptic problems with uncertainty. Numerische Mathematik, 101(2):185–219, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=2&spage=185.

Bacuta:2005:IRC

[BNZ05] Constantin Băcu tă, Victor Nistor, and Ludmil T. Zikatanov. Improving the rate of convergence of ‘high order finite elements’ on polygons and domains with cusps. Numerische Mathematik, 100 (2):165–184, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Boche:2001:KLI

[Boc01] Holger Boche. Konvergente Lagrange-Interpolationspolynome für die Disk-Algebra. (German) [Convergent Lagrange interpolation polynomials for disk algebra]. Numerische Mathematik, 88(1):1– 7, March 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088001/10880001.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088001/10880001.pdf.

Boffi:2000:FOD

[Bof00] Daniele Boffi. Fortin operator and discrete compactness for edge elements. Numerische Mathematik, 87(2):229–246, December 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087002/00870229.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087002/00870229.pdf.

Bar-On:2002:RPS

[BOL02] Ilan Bar-On and Mauro Leoncini. Reliable parallel solution of bidiagonal systems. Numerische Mathematik, 90(3):415–440, January 2002. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/2090003/20900415.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/2090003/20900415.pdf.

Borzi:2004:SLO

[Bor04] Alfio Borzì. Solution of lambda-omega systems: Theta-schemes and multigrid methods. Numerische Mathematik, 98(4):581–606, October 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bourdarias:2001:CFS

[Bou01] C. Bourdarias. Convergence of fluctuation-splitting schemes for two dimensional scalar conservation laws with a kinetic solver. Numerische Mathematik, 87(4):645–662, February 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1087004/10870645.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1087004/
10870645.pdf.

Bouchut:2003:ESF

[Bou03] F. Bouchut. Entropy satisfying flux vector splittings and kinetic BGK models. Numerische Mathematik, 94(4):623–672, June 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=94&issue=4&spage=
623.

Bouchon:2007:MSP

[Bou07] François Bouchon. Monotonicity of some perturbations of irreducibly diagonally dominant M-matrices. Numerische Mathematik, 105(4):591–601, February 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bojanov:2001:GIQ

[BP01] Borislav D. Bojanov and Petar P. Petrov. Gaussian interval quadrature formula. Numerische Mathematik, 87(4):625–643, February 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1087004/10870625.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1087004/10870625.pdf.

Bojanov:2003:UGI

[BP03] B. Bojanov and P. Petrov. Uniqueness of the Gaussian interval quadrature formula. Numerische Mathematik, 95(1):53–62, July 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:2004:MRC

[BP04] Sören Bartels and Andreas Prohl. Multiscale resolution in the computation of crystalline microstructure. Numerische Mathematik, 96(4):641–660, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:2008:CIC

[BP08] Sören Bartels and Andreas Prohl. Convergence of an implicit, constraint preserving finite element discretization of p-harmonic heat flow into spheres. Numerische Mathematik, 109(4):489–507, June 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bai:2005:GSO

[BPW05] Zhong-Zhi Bai, Beresford N. Parlett, and Zeng-Qi Wang. On generalized successive overrelaxation methods for augmented linear systems. Numerische Mathematik, 102(1):1–38, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=102&issue=1&spage=
1.

Brandle:2005:ANM

[BQR05] Cristina Brändle, Fernando Quirós, and Julio D. Rossi. An adaptive numerical method to handle blow-up in a parabolic system. Numerische Mathematik, 102(1):39–59, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=1&spage=39.

Boichu:2002:MPB

[BR02] Daniel Boichu and Vincent Robin. Multivariate Padé-Bergman approximants. Numerische Mathematik, 93(2):223–238, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:2004:LPA

[BR04] Sören Bartels and Tomás Roubícek. Linear-programming approach to nonconvex variational problems. Numerische Mathematik, 99(2):251–287, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:2005:PEE

[Bra05] Dietrich Braess. A posteriori error estimators for obstacle problems — another look. Numerische Mathematik, 101 (3):415–421, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=3&spage=415.

Brenner:2003:ASP

[Bre03] Susanne C. Brenner. An additive Schwarz preconditioner for the FETI method. Numerische Mathematik, 94(1):1–31, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernardi:2004:MTC

[BRM^{+}04] C. Bernardi, T. Chacón Rebollo, M. Gómez Mármol, R. Lewandowski, and F. Murat. A model for two coupled turbulent fluids. Part III: Numerical approximation by finite elements. Numerische Mathematik, 98(1):33–66, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bermudez:2000:FES

[BRS00] Alfredo Bermúdez, Rodolfo Rodríguez, and Duarte Santamarina. A finite element solution of an added mass formulation for coupled fluid-solid vibrations. Numerische Mathematik, 87(2):201–227, December 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087002/00870201.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087002/00870201.pdf.

Bermudez:2005:NAE

[BRS05] Alfredo Bermúdez, Rodolfo Rodríguez, and Pilar Salgado. Numerical analysis of electric field formulations of the eddy current model. Numerische Mathematik, 102(2):181–201, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=102&issue=2&spage=
181.

Brezinski:2003:MPR

[BRZRS03] C. Brezinski, M. Redivo-Zaglia, G. Rodriguez, and S. Seatzu. Multi-parameter regularization techniques for ill-conditioned linear systems. Numerische Mathematik, 94(2):203–228, April 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Benzi:2006:ECS

[BS06a] Michele Benzi and Valeria Simoncini. On the eigenvalues of a class of saddle point matrices. Numerische Mathematik, 103(2):173–196, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=2&spage=173.

Beuchler:2006:NSF

[BS06b] S. Beuchler and J. Schöberl. New shape functions for triangular p-FEM using integrated Jacobi polynomials. Numerische Mathematik, 103(3):339–366, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=103&issue=3&spage=339.

Beuchler:2004:MWN

[BSS04] S. Beuchler, R. Schneider, and C. Schwab. Multiresolution weighted norm equivalences and applications. Numerische Mathematik, 98(1):67–97, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [BSS05].

Beuchler:2005:EMW

[BSS05] S. Beuchler, R. Schneider, and C. Schwab. Erratum: Multiresolution weighted norm equivalences and applications, Num. Math. 98(1) 67–97 (2004). Numerische Mathematik, 102 (2):365–366, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=2&spage=365. See [BSS04].

Buffa:2002:EES

[Buf02] Annalisa Buffa. Error estimate for a stabilised domain decomposition method with nonmatching grids. Numerische Mathematik, 90 (4):617–640, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/2090004/20900617.htm;
 http://link.springer.de/link/service/journals/00211/
papers/2090004/20900617.pdf.

Butler:2004:NCE

[But04] Neil A. Butler. A new class of equal-weight integration rules on the hypercube. Numerische Mathematik, 99(2):349–363, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernardi:2000:AFE

[BV00] C. Bernardi and R. Verfürth. Adaptive finite element methods for elliptic equations with non-smooth coefficients. Numerische Mathematik, 85(4):579–608, June 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085004/00850579.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085004/00850579.pdf.

Barrenechea:2002:USF

[BV02] Gabriel R. Barrenechea and Frédéric Valentin. An unusual stabilized finite element method for a generalized Stokes problem. Numerische Mathematik, 92(4):653–677, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Becker:2004:PEE

[BV04] Roland Becker and Boris Vexler. A posteriori error estimation for finite element discretization of parameter identification problems. Numerische Mathematik, 96(3):435–459, January 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Becker:2007:OCC

[BV07] Roland Becker and Boris Vexler. Optimal control of the convection-diffusion equation using stabilized finite element methods. Numerische Mathematik, 106(3):349–367, May 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buzdin:2004:TFD

[BW04] A. Buzdin and G. Wittum. Two-frequency decomposition. Numerische Mathematik, 97(2):269–295, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brenner:2005:TLA

[BW05] Susanne C. Brenner and Kening Wang. Two-level additive Schwarz preconditioners for C^{0} interior penalty methods. Numerische Mathematik, 102(2):231–255, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=2&spage=231.

Brunner:2005:FEM

[BY05] Hermann Brunner and Ningning Yan. Finite element methods for optimal control problems governed by integral equations and integro-differential equations. Numerische Mathematik, 101(1):1–27, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=1&spage=1.

Cano:2006:CQS

[Can06] B. Cano. Conserved quantities of some Hamiltonian wave equations after full discretization. Numerische Mathematik, 103(2):197–223, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=2&spage=197.

Cao:2001:CBT

[Cao01a] Zhi-Hao Cao. Convergence of block two-stage iterative methods for symmetric positive definite systems. Numerische Mathematik, 90(1):47–63, November 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1090001/10900047.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1090001/10900047.pdf.

Cao:2001:NPS

[Cao01b] Zhi-Hao Cao. A note on properties of splittings of singular symmetric positive semidefinite matrices. Numerische Mathematik, 88(4):603–606, June 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1088004/10880603.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1088004/10880603.pdf.

Cao:2006:MAE

[Cao06a] Li-Qun Cao. Multiscale asymptotic expansion and finite element methods for the mixed boundary value problems of second order elliptic equation in perforated domains. Numerische Mathematik, 103(1):11–45, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=103&issue=1&spage=11.

Cao:2006:CCP

[Cao06b] Z.-H. Cao. A class of constraint preconditioners for nonsymmetric saddle point matrices. Numerische Mathematik, 103 (1):47–61, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=1&spage=47.

Capizzano:2002:CAT

[Cap02] Stefano Serra Capizzano. Convergence analysis of two-grid methods for elliptic Toeplitz and PDEs matrix-sequences. Numerische Mathematik, 92(3):433–465, September 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2005:UTP

[Car05] C. Carstensen. A unifying theory of a posteriori finite element error control. Numerische Mathematik, 100(4):617–637, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=100&issue=4&spage=
617.

Casarin:2001:SPS

[Cas01] Mario A. Casarin. Schwarz preconditioners for the spectral element discretization of the steady Stokes and Navier– Stokes equations. Numerische Mathematik, 89(2):307–339, August 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089002/10890307.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089002/10890307.pdf.

Codina:2000:APS

[CB00] Ramon Codina and Jordi Blasco. Analysis of a pressurestabilized finite element approximation of the stationary Navier– Stokes equations. Numerische Mathematik, 87(1):59–81, November 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087001/00870059.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087001/00870059.pdf.

Carstensen:2002:PEE

[CBJ02] Carsten Carstensen, Sören Bartels, and Stefan Jansche. A posteriori error estimates for nonconforming finite element methods. Numerische Mathematik, 92(2):233–256, August 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:2000:CPS

[CC00] Raymond H. Chan and Wai Ki Ching. Circulant preconditioners for stochastic automata networks. Numerische Mathematik, 87 (1):35–57, November 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0087001/00870035.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0087001/00870035.pdf.

Cao:2004:AEN

[CC04] Li-Qun Cao and Jun-Zhi Cui. Asymptotic expansions and numerical algorithms of eigenvalues and eigenfunctions of the Dirichlet problem for second order elliptic equations in perforated domains. Numerische Mathematik, 96(3):525–581, January 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chalons:2005:NSE

[CC05] C. Chalons and F. Coquel. Navier–Stokes equations with several independent pressure laws and explicit predictor-corrector schemes. Numerische Mathematik, 101(3):451–478, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0029-599X&volume=101&issue=3&spage=
451.

Cances:2005:LTA

[CCC^{+}05] Eric Cancès, François Castella, Philippe Chartier, Erwan Faou, Claude Le Bris, Frédéric Legoll, and Gabriel Turinici. Longtime averaging for integrable Hamiltonian dynamics. Numerische Mathematik, 100(2):211–232, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=100&issue=2&spage=211.

Charina:2008:AFM

[CCF08] Maria Charina, Costanza Conti, and Massimo Fornasier. Adaptive frame methods for nonlinear variational problems. Numerische Mathematik, 109(1):45–75, March 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:2004:RMR

[CCM04] R. P. K. Chan, P. Chartier, and A. Murua. Reversible methods of Runge–Kutta type for Index-2 DAEs. Numerische Mathematik, 97(3):427–440, May 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvo:2007:RCQ

[CCP07] M. P. Calvo, E. Cuesta, and C. Palencia. Runge–Kutta convolution quadrature methods for well-posed equations with memory. Numerische Mathematik, 107(4):589–614, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2005:PDM

[CCS05] Carsten Carstensen, Paola Causin, and Riccardo Sacco. A posteriori dual-mixed adaptive finite element error control for Lamé and Stokes equations. Numerische Mathematik, 101(2):309–332, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=2&spage=309.

Cai:2009:SII

[CCSS09] Jian-Feng Cai, Raymond H. Chan, Lixin Shen, and Zuowei Shen. Simultaneously inpainting in image and transformed domains. Numerische Mathematik, 112(4):509–533, June 2009. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=112&issue=
4&spage=509.

Costabel:2002:WRM

[CD02] Martin Costabel and Monique Dauge. Weighted regularization of Maxwell equations in polyhedral domains: A rehabilitation of nodal finite elements. Numerische Mathematik, 93(2):239–277, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2004:PEE

[CD04] Carsten Carstensen and Georg Dolzmann. An a priori error estimate for finite element discretizations in nonlinear elasticity for polyconvex materials under small loads. Numerische Mathematik, 97(1):67–80, March 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cristea:2007:TTC

[CDLP07] Ligia L. Cristea, Josef Dick, Gunther Leobacher, and Friedrich Pillichshammer. The tent transformation can improve the convergence rate of quasi-Monte Carlo algorithms using digital nets. Numerische Mathematik, 105(3):413–455, January 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Casado-Daz:2007:FEA

[CDRG^{+}07] J. Casado-Daz, T. Chacón Rebollo, V. Girault, M. Gómez Mármol, and F. Murat. Finite elements approximation of second order linear elliptic equations in divergence form with righthand side in L^{1}. Numerische Mathematik, 105(3):337–374, January 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:2001:DTP

[CE01] Zhangxin Chen and Richard E. Ewing. Degenerate two-phase incompressible flow III. Sharp error estimates. Numerische Mathematik, 90(2):215–240, December 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1090002/10900215.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1090002/10900215.pdf.

Chen:2008:NSI

[CF08] Zhuliang Chen and Peter A. Forsyth. A numerical scheme for the impulse control formulation for pricing variable annuities with a guaranteed minimum withdrawal benefit (GMWB). Numerische Mathematik, 109(4):535–569, June 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cascon:2006:STA

[CFA06] J. M. Cascón, L. Ferragut, and M. I. Asensio. Space–time adaptive algorithm for the mixed parabolic problem. Numerische Mathematik, 103(3):367–392, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=103&issue=3&spage=367.

Collino:2003:CSTa

[CFJ03a] F. Collino, T. Fouquet, and P. Joly. A conservative space–time mesh refinement method for the 1-D wave equation. Part I: Construction. Numerische Mathematik, 95(2):197–221, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Collino:2003:CSTb

[CFJ03b] F. Collino, T. Fouquet, and P. Joly. A conservative space–time mesh refinement method for the 1-D wave equation. Part II: Analysis. Numerische Mathematik, 95(2):223–251, August 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chartier:2006:AAI

[CFM06] Philippe Chartier, Erwan Faou, and Ander Murua. An algebraic approach to invariant preserving integrators: The case of quadratic and Hamiltonian invariants. Numerische Mathematik, 103(4):575–590, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandrasekaran:2004:DCA

[CG04] S. Chandrasekaran and M. Gu. A divide-and-conquer algorithm for the eigendecomposition of symmetric block-diagonal plus semiseparable matrices. Numerische Mathematik, 96(4):723–731, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2009:UTP

[CGJ09] Carsten Carstensen, Thirupathi Gudi, and Max Jensen. A unifying theory of a posteriori error control for discontinuous Galerkin FEM. Numerische Mathematik, 112(3):363–379, May 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=3&spage=363.

Carnicer:2006:ILS

[CGS06] J. M. Carnicer, M. Gasca, and T. Sauer. Interpolation lattices in several variables. Numerische Mathematik, 102(4): 559–581, February 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=4&spage=559.

Carstensen:2006:CAA

[CH06] Carsten Carstensen and Ronald H. W. Hoppe. Convergence analysis of an adaptive nonconforming finite element method. Numerische Mathematik, 103(2):251–266, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=103&issue=2&spage=251.

Carstensen:2007:UTP

[CH07] C. Carstensen and Jun Hu. A unifying theory of a posteriori error control for nonconforming finite element methods. Numerische Mathematik, 107(3):473–502, September 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chainais-Hillairet:2008:MNS

[CHB08] Claire Chainais-Hillairet and Christian Bataillon. Mathematical and numerical study of a corrosion model. Numerische Mathematik, 110(1):1–25, July 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chainais-Hillairet:2001:FVS

[CHC01] Claire Chainais-Hillairet and Sylvie Champier. Finite volume schemes for nonhomogeneous scalar conservation laws: error estimate. Numerische Mathematik, 88(4):607–639, June 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088004/10880607.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088004/
10880607.pdf.

Chen:2001:ALA

[Che01] Zhiming Chen. On the augmented Lagrangian approach to Signorini elastic contact problem. Numerische Mathematik, 88(4): 641–659, June 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088004/10880641.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088004/10880641.pdf.

Chen:2007:PBG

[Che07] Xiao Shan Chen. On perturbation bounds of generalized eigenvalues for diagonalizable pairs. Numerische Mathematik, 107(1):

79–86, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cohen:2008:CEM

[CHL08] David Cohen, Ernst Hairer, and Christian Lubich. Conservation of energy, momentum and actions in numerical discretizations of non-linear wave equations. Numerische Mathematik, 110(2): 113–143, August 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=110&issue=2&spage=113.

Christiansen:2007:SHD

[Chr07] Snorre H. Christiansen. Stability of Hodge decompositions in finite element spaces of differential forms in arbitrary dimension. Numerische Mathematik, 107(1):87–106, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chu:2003:FPD

[Chu03] Delin Chu. The fixed poles of the disturbance decoupling problem and almost stability subspace V_{b,g}^{⋆}(ker(C)). Numerische Mathematik, 96(2):221–252, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chun:2006:CNL

[Chu06] Changbum Chun. Construction of Newton-like iteration methods for solving nonlinear equations. Numerische Mathematik, 104(3): 297–315, September 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Caselles:2006:AAS

[CIS06] V. Caselles, L. Igual, and O. Sander. An axiomatic approach to scalar data interpolation on surfaces. Numerische Mathematik, 102(3):383–411, January 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=3&spage=383.

Ciarlet:2005:FSC

[CJK^{+}05] P. Ciarlet, Jr., B. Jung, S. Kaddouri, S. Labrunie, and J. Zou. The Fourier Singular Complement Method for the Poisson problem. Part I: Prismatic domains. Numerische Mathematik, 101(3):423–450, September 2005. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=3&spage=423.

Ciarlet:2006:FSC

[CJK^{+}06] P. Ciarlet, B. Jung, S. Kaddouri, S. Labrunie, and J. Zou. The Fourier Singular Complement Method for the Poisson problem. Part II: Axisymmetric domains. Numerische Mathematik, 102(4):583–610, February 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=4&spage=583.

Calvo:2004:ACN

[CJMR04] Manuel Calvo, Laurent O. Jay, Juan I. Montijano, and Luis Rández. Approximate compositions of a near identity map by multi-revolution Runge–Kutta methods. Numerische Mathematik, 97(4):635–666, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chou:2002:MAV

[CK02] So-Hsiang Chou and Do Y. Kwak. Multigrid algorithms for a vertex-centered covolume method for elliptic problems. Numerische Mathematik, 90(3):441–458, January 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/2090003/20900441.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/2090003/
20900441.pdf.

Canuto:2007:FDA

[CK07] Claudio Canuto and Tomas Kozubek. A fictitious domain approach to the numerical solution of PDEs in stochastic domains. Numerische Mathematik, 107(2):257–293, August 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chertock:2008:SOP

[CK08] Alina Chertock and Alexander Kurganov. A second-order positivity preserving central-upwind scheme for chemotaxis and haptotaxis models. Numerische Mathematik, 111(2):169–205, December 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=2&spage=169.

Coomes:2007:TCO

[CKP07] Brian A. Coomes, Hüseyin Koçak, and Kenneth J. Palmer. Transversal connecting orbits from shadowing. Numerische Mathematik, 106(3):427–469, May 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chalons:2001:FDS

[CL01] C. Chalons and P. G. LeFloch. A fully discrete scheme for diffusive-dispersive conservation laws. Numerische Mathematik, 89(3):493–509, September 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1089003/10890493.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089003/10890493.pdf.

Chen:2008:ECS

[CLL08] Wenbin Chen, Xingjie Li, and Dong Liang. Energy-conserved splitting FDTD methods for Maxwell’s equations. Numerische Mathematik, 108(3):445–485, January 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chu:2003:QTR

[CLM03] Delin Chu, Lieven De Lathauwer, and Bart De Moor. A QR-type reduction for computing the SVD of a general matrix product/ quotient. Numerische Mathematik, 95(1):101–121, July 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Costantini:2005:CWT

[CLM05] Paolo Costantini, Tom Lyche, and Carla Manni. On a class of weak Tchebycheff systems. Numerische Mathematik, 101(2):333–354, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=2&spage=333.

Chu:2007:NMC

[CLM07] Delin Chu, Xinmin Liu, and Volker Mehrmann. A numerical method for computing the Hamiltonian Schur form. Numerische Mathematik, 105(3):375–412, January 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvetti:2002:RPG

[CLR02] D. Calvetti, B. Lewis, and L. Reichel. On the regularizing properties of the GMRES method. Numerische Mathematik, 91(4):605– 625, June 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carlier:2001:NAV

[CLRM01] G. Carlier, T. Lachand-Robert, and B. Maury. A numerical approach to variational problems subject to convexity constraint. Numerische Mathematik, 88(2):299–318, April 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088002/10880299.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088002/10880299.pdf.

Cohen:2000:WAM

[CM00] A. Cohen and R. Masson. Wavelet adaptive method for second order elliptic problems: boundary conditions and domain decomposition. Numerische Mathematik, 86(2):193–238, August 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086002/00860193.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086002/00860193.pdf.

Castro:2006:BCL

[CM06] Carlos Castro and Sorin Micu. Boundary controllability of a linear semi-discrete 1-D wave equation derived from a mixed finite element method. Numerische Mathematik, 102(3):413–462, January 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=3&spage=413.

Cuesta:2008:PEE

[CM08] E. Cuesta and Ch. Makridakis. A posteriori error estimates and maximal regularity for approximations of fully nonlinear parabolic problems in Banach spaces. Numerische Mathematik, 110(3):257– 275, September 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=110&issue=3&spage=257.

Carstensen:2004:RBP

[CMPS04] Carsten Carstensen, M. Maischak, D. Praetorius, and E. P. Stephan. Residual-based a posteriori error estimate for hypersingular equation on surfaces. Numerische Mathematik, 97(3):397– 425, May 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2001:PEE

[CMS01] C. Carstensen, M. Maischak, and E. P. Stephan. A posteriori error estimate and h-adaptive algorithm on surfaces for Symm’s integral equation. Numerische Mathematik, 90(2):197–213, December 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090002/10900197.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090002/10900197.pdf.

Chen:2000:RTP

[CN00] Zhiming Chen and Ricardo H. Nochetto. Residual type a posteriori error estimates for elliptic obstacle problems. Numerische Mathematik, 84(4):527–548, February 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0084004/00840527.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0084004/00840527.pdf.

Chan:2002:BCP

[CNY02] Raymond H. Chan, Michael K. Ng, and Andy M. Yip. The best circulant preconditioners for Hermitian Toeplitz systems II: The multiple-zero case. Numerische Mathematik, 92(1):17–40, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Copetti:2008:QSS

[Cop08] M. I. M. Copetti. A quasi-static Signorini contact problem for a thermoviscoelastic beam. Numerische Mathematik, 110(1):27–47, July 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2001:NAR

[CP01a] Carsten Carstensen and Andreas Prohl. Numerical analysis of relaxed micromagnetics by penalised finite elements. Numerische Mathematik, 90(1):65–99, November 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1090001/10900065.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1090001/10900065.pdf.

Chang:2001:CPA

[CP01b] Xiao-Wen Chang and Christopher C. Paige. Componentwise perturbation analyses for the QR factorization. Numerische Mathematik, 88(2):319–345, April 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088002/10880319.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088002/10880319.pdf.

Calvo:2006:CEM

[CP06] M. P. Calvo and C. Palencia. A class of explicit multistep exponential integrators for semilinear problems. Numerische Mathematik, 102(3):367–381, January 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=3&spage=367.

Carstensen:2000:NAY

[CR00] Carsten Carstensen and Tomás Roubíček. Numerical approximation of Young measures in non-convex variational problems. Numerische Mathematik, 84(3):395–415, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0084003/00840395.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0084003/00840395.pdf.

Caboussat:2005:AOD

[CR05] Alexandre Caboussat and Jacques Rappaz. Analysis of a onedimensional free boundary flow problem. Numerische Mathematik, 101(1):67–86, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=1&spage=67.

Chilton:2000:CSC

[CS00a] Lawrence Chilton and Manil Suri. On the construction of stable curvilinear p version elements for mixed formulations of elasticity and Stokes flow. Numerische Mathematik, 86(1):29– 48, July 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086001/00860029.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086001/00860029.pdf.

Costabel:2000:SCC

[CS00b] M. Costabel and J. Saranen. Spline collocation for convolutional parabolic boundary integral equations. Numerische Mathematik, 84(3):417–449, October 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0084003/00840417.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0084003/00840417.pdf.

Costabel:2003:SCM

[CS03] M. Costabel and J. Saranen. The spline collocation method for parabolic boundary integral equations on smooth curves. Numerische Mathematik, 93(3):549–562, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2004:PEA

[CS04] Carsten Carstensen and Stefan A. Sauter. A posteriori error analysis for elliptic PDEs on domains with complicated structures. Numerische Mathematik, 96(4):691–721, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cangiani:2005:ERM

[CS05] Andrea Cangiani and Endre Süli. Enhanced RFB method. Numerische Mathematik, 101(2):273–308, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=101&issue=2&spage=273.

Carstensen:2006:RBP

[CS06] Carsten Carstensen and Joachim Schöberl. Residual-based a posteriori error estimate for a mixed Reißner–Mindlin plate finite element method. Numerische Mathematik, 103(2):

225–250, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=2&spage=225.

Chai:2007:DWF

[CS07] Anwei Chai and Zuowei Shen. Deconvolution: a wavelet frame approach. Numerische Mathematik, 106(4):529–587, June 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cucker:2003:ECN

[CW03] Felipe Cucker and Mario Wschebor. On the expected condition number of linear programming problems. Numerische Mathematik, 94(3):419–478, May 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandler-Wilde:2002:FTG

[CWRR02] Simon N. Chandler-Wilde, Mizanur Rahman, and C. R. Ross. A fast two-grid and finite section method for a class of integral equations on the real line with application to an acoustic scattering problem in the half-plane. Numerische Mathematik, 93(1):1–51, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:2008:SAA

[CX08] Long Chen and Jinchao Xu. Stability and accuracy of adapted finite element methods for singularly perturbed problems. Numerische Mathematik, 109(2):167–191, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cao:2007:FEM

[CYY07] Yanzhao Cao, Hongtao Yang, and Li Yin. Finite element methods for semilinear elliptic stochastic partial differential equations. Numerische Mathematik, 106(2):181–198, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chenais:2003:CEE

[CZ03] D. Chenais and Enrique Zuazua. Controllability of an elliptic equation and its finite difference approximation by the shape of the domain. Numerische Mathematik, 95(1):63–99, July 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Da:2008:SDT

[Da08] Xu Da. Stability of the difference type methods for linear Volterra equations in Hilbert spaces. Numerische Mathematik, 109(4):571– 595, June 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dai:2001:NPN

[Dai01] Yu-Hong Dai. New properties of a nonlinear conjugate gradient method. Numerische Mathematik, 89(1):83–98, July 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1089001/10890083.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1089001/
10890083.pdf.

Deckelnick:2002:FDN

[DD02] Klaus Deckelnick and Gerhard Dziuk. A fully discrete numerical scheme for weighted mean curvature flow. Numerische Mathematik, 91(3):423–452, May 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=91&issue=3&spage=423.

DeBouard:2004:SDS

[DD04] A. De Bouard and A. Debussche. A semi-discrete scheme for the stochastic nonlinear Schrödinger equation. Numerische Mathematik, 96(4):733–770, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:2007:FLA

[DDH07] James Demmel, Ioana Dumitriu, and Olga Holtz. Fast linear algebra is stable. Numerische Mathematik, 108(1):59–91, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:2007:FMM

[DDHK07] James Demmel, Ioana Dumitriu, Olga Holtz, and Robert Kleinberg. Fast matrix multiplication is stable. Numerische Mathematik, 106(2):199–224, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dahmen:2007:MPP

[DDP07] Wolfgang Dahmen, Shai Dekel, and Pencho Petrushev. Multilevel preconditioning for partition of unity methods: some analytic concepts. Numerische Mathematik, 107(3):503–532, September 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

delaCalleYsern:2005:EBR

[de 05] Bernardo de la Calle Ysern. Error bounds for rational quadrature formulae of analytic functions. Numerische Mathematik, 101(2):251–271, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=2&spage=251.

Droniou:2006:MFV

[DE06] Jérôme Droniou and Robert Eymard. A mixed finite volume scheme for anisotropic diffusion problems on any grid. Numerische Mathematik, 105(1):35–71, November 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Despres:2001:LSC

[Des01] Bruno Després. Lagrangian systems of conservation laws. Numerische Mathematik, 89(1):99–134, July 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://link.springer-ny.com/link/service/journals/00211/
bibs/1089001/10890099.htm;   http://link.springer-ny.
com/link/service/journals/00211/papers/1089001/10890099.
pdf.

Despres:2008:FVT

[Des08] Bruno Després. Finite volume transport schemes. Numerische Mathematik, 108(4):529–556, February 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deuring:2009:FEE

[Deu09] Paul Deuring. Finite element error estimates for 3D exterior incompressible flow with nonzero velocity at infinity. Numerische Mathematik, 114(2):233–270, December 2009. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=114&issue=
2&spage=233.

Dai:2005:PBB

[DF05] Yu-Hong Dai and Roger Fletcher. Projected Barzilai–Borwein methods for large-scale box-constrained quadratic programming. Numerische Mathematik, 100(1):21–47, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dujardin:2007:NFL

[DF07] Guillaume Dujardin and Erwan Faou. Normal form and long time analysis of splitting schemes for the linear Schrödinger equation with small potential. Numerische Mathematik, 108(2):223–262, December 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deFrutos:2000:PGM

[dFGAN00] Javier de Frutos, Bosco García-Archilla, and Julia Novo. A postprocessed Galerkin method with Chebyshev or Legendre polynomials. Numerische Mathematik, 86(3):419–442, September 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086003/00860419.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086003/00860419.pdf.

dHalluin:2004:PMA

[dFL04] Y. d’Halluin, P. A. Forsyth, and G. Labahn. A penalty method for American options with jump diffusion processes. Numerische Mathematik, 97(2):321–352, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DiFiore:2003:MAQ

[DFLZ03] Carmine Di Fiore, Stefano Fanelli, Filomena Lepore, and Paolo Zellini. Matrix algebras in Quasi-Newton methods for unconstrained minimization. Numerische Mathematik, 94(3):479–500, May 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

daFonseca:2005:EIT

[dFP05] C. M. da Fonseca and J. Petronilho. Explicit inverse of a tridiagonal k-Toeplitz matrix. Numerische Mathematik, 100(3):457–482, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deckelnick:2009:FEA

[DGH09] Klaus Deckelnick, Andreas Günther, and Michael Hinze. Finite element approximation of elliptic control problems with constraints on the gradient. Numerische Mathematik, 111(3):335–350, January 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=3&spage=335.

Dahmen:2001:MSC

[DGMM01] Wolfgang Dahmen, Birgit Gottschlich-Müller, and Siegfried Müller. Multiresolution schemes for conservation laws. Numerische Mathematik, 88(3):399–443, May 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://link.springer-ny.com/link/service/journals/00211/
bibs/1088003/10880399.htm;   http://link.springer-ny.
com/link/service/journals/00211/papers/1088003/10880399.
pdf.

Domnguez:2007:HNA

[DGS07] V. Domínguez, I. G. Graham, and V. P. Smyshlyaev. A hybrid numerical-asymptotic boundary integral method for highfrequency acoustic scattering. Numerische Mathematik, 106(3): 471–510, May 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Daruis:2002:IQF

[DGV02] Leyla Daruis and Pablo González-Vera. Interpolatory quadrature formulas on the unit circle for Chebyshev weight functions. Numerische Mathematik, 90(4):641–664, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://link.springer.de/link/service/
journals/00211/bibs/2090004/20900641.htm;   http://link.
springer.de/link/service/journals/00211/papers/2090004/
20900641.pdf.

Deckelnick:2004:SEE

[DH04] Klaus Deckelnick and Michael Hinze. Semidiscretization and error estimates for distributed control of the instationary Navier– Stokes equations. Numerische Mathematik, 97(2):297–320, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dziuk:2006:FEA

[DH06] Gerhard Dziuk and John E. Hutchinson. Finite element approximations to surfaces of prescribed variable mean curvature. Numerische Mathematik, 102(4):611–648, February 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=4&spage=611.

Deano:2009:CGQ

[DH09] Alfredo Deaño and Daan Huybrechs. Complex Gaussian quadrature of oscillatory integrals. Numerische Mathematik, 112(2):197– 219, April 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=2&spage=197.

Duran:2007:CNG

[DHN07] Mario Durán, Ricardo Hein, and Jean-Claude Nédélec. Computing numerically the Green’s function of the half-plane Helmholtz operator with impedance boundary conditions. Numerische Mathematik, 107(2):295–314, August 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duran:2000:FEA

[DHNL^{+}00] R. G. Durán, L. Hervella-Nieto, E. Liberman, R. Rodríguez, and J. Solomin. Finite element analysis of the vibration problem of a plate coupled with a fluid. Numerische Mathematik, 86(4):591– 616, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086004/00860591.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086004/00860591.pdf.

Demkowicz:2001:ACF

[DI01] Leszek Demkowicz and Frank Ihlenburg. Analysis of a coupled finite-infinite element method for exterior Helmholtz problems. Numerische Mathematik, 88(1):43–73, March 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088001/10880043.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088001/
10880043.pdf.

Driver:2007:IZS

[DJ07] Kathy Driver and Kerstin Jordaan. Interlacing of zeros of shifted sequences of one-parameter orthogonal polynomials. Numerische Mathematik, 107(4):615–624, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dahmen:2001:ABC

[DK01] Wolfgang Dahmen and Angela Kunoth. Appending boundary conditions by Lagrange multipliers: Analysis of the LBB condition. Numerische Mathematik, 88(1):9–42, March 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088001/10880009.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088001/
10880009.pdf.

Demmel:2004:ASW

[DK04] James Demmel and Plamen Koev. Accurate SVDs of weakly diagonally dominant M-matrices. Numerische Mathematik, 98(1): 99–104, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dopico:2009:ISJ

[DKM09] Froilán M. Dopico, Plamen Koev, and Juan M. Molera. Implicit standard Jacobi gives high relative accuracy. Numerische Mathematik, 113(4):519–553, October 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=4&spage=
519.

Ysern:2007:RQF

[dlCYGV07] B. de la Calle Ysern and P. González-Vera. Rational quadrature formulae on the unit circle with arbitrary poles. Numerische Mathematik, 107(4):559–587, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Descombes:2004:OSN

[DM04] Stéphane Descombes and Marc Massot. Operator splitting for nonlinear reaction-diffusion systems with an entropic structure : singular perturbation and order reduction. Numerische Mathematik, 97(4):667–698, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dawson:2000:CGA

[DMC00] Clint N. Dawson and Mónica L. Martínez-Canales. A characteristic-Galerkin approximation to a system of shallow water equations. Numerische Mathematik, 86(2):239–256, August 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0086002/00860239.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0086002/
00860239.pdf.

Dorfler:2002:SDO

[DN02] Willy Dörfler and Ricardo H. Nochetto. Small data oscillation implies the saturation assumption. Numerische Mathematik, 91(1): 1–12, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dontchev:2001:CNM

[DQQ01] Asen L. Dontchev, Houduo Qi, and Liqun Qi. Convergence of Newton’s method for convex best interpolation. Numerische Mathematik, 87(3):435–456, January 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087003/10870435.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087003/10870435.pdf.

Descombes:2003:CPR

[DR03] Stéphane Descombes and Magali Ribot. Convergence of the Peaceman–Rachford approximation for reaction-diffusion systems. Numerische Mathematik, 95(3):503–525, September 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dedner:2004:NAE

[DR04a] Andreas Dedner and Christian Rohde. Numerical approximation of entropy solutions for hyperbolic integro-differential equations. Numerische Mathematik, 97(3):441–471, May 2004. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dodu:2004:IDF

[DR04b] Fabrice Dodu and Christophe Rabut. Irrotational or divergencefree interpolation. Numerische Mathematik, 98(3):477–498, September 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Diening:2007:IOO

[DR07] L. Diening and M. Ruzika. Interpolation operators in Orlicz– Sobolev spaces. Numerische Mathematik, 107(1):107–129, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dahmen:2008:AEC

[DRSZ08] Wolfgang Dahmen, Thorsten Rohwedder, Reinhold Schneider, and Andreas Zeiser. Adaptive eigenvalue computation: complexity estimates. Numerische Mathematik, 110(3):277–312, September 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=3&spage=277.

Dryja:2005:NNA

[Dry05] Maksymilian Dryja. A Neumann–Neumann algorithm for a mortar discretization of elliptic problems with discontinuous coefficients. Numerische Mathematik, 99(4):645–656, February 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

delSastre:2008:EEP

[dSB08] Pedro Galán del Sastre and Rodolfo Bermejo. Error estimates of proper orthogonal decomposition eigenvectors and Galerkin projection for a general dynamical system arising in fluid models. Numerische Mathematik, 110(1):49–81, July 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dubois:2002:DHS

[DSS02] François Dubois, Michel Salaün, and Stéphanie Salmon. Discrete harmonics for stream function-vorticity Stokes problem. Numerische Mathematik, 92(4):711–742, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Degani:2005:CEC

[DST05] Ilan Degani, Jeremy Schiff, and David J. Tannor. Commuting extensions and cubature formulae. Numerische Mathematik, 101(3):479–500, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=3&spage=479.

Dick:2006:GLR

[DSWW06] Josef Dick, Ian H. Sloan, Xiaoqun Wang, and Henryk Woźniakowski. Good lattice rules in weighted Korobov spaces with general weights. Numerische Mathematik, 103(1):63–97, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=103&issue=1&spage=
63.

Dahmen:2000:NFW

[DSX00] Wolfgang Dahmen, Reinhold Schneider, and Yuesheng Xu. Nonlinear functionals of wavelet expansions — adaptive reconstruction and fast evaluation. Numerische Mathematik, 86(1):49– 101, July 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086001/00860049.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086001/00860049.pdf.

Dryja:2007:DDD

[DT07] Maksymilian Dryja and Xuemin Tu. A domain decomposition discretization of parabolic problems. Numerische Mathematik, 107 (4):625–640, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DeSwardt:2000:GTQ

[DV00] S. A. De Swardt and J. M. De Villiers. Gregory type quadrature based on quadratic nodal spline interpolation. Numerische Mathematik, 85(1):129–153, March 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085001/00850129.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085001/00850129.pdf.

daVeiga:2004:UEE

[dV04] L. Beirão da Veiga. Uniform error estimates for a class of intermediate cylindrical shell problems. Numerische Mathematik, 96(4): 661–689, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Delvaux:2005:ORF

[DV05a] Steven Delvaux and Marc Van Barel. Orthonormal rational function vectors. Numerische Mathematik, 100(3):409–440, May 2005.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dieci:2005:ECL

[DV05b] Luca Dieci and Erik S. Van Vleck. On the error in computing Lyapunov exponents by QR Methods. Numerische Mathematik, 101(4):619–642, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=4&spage=619.

daVeiga:2008:RBE

[dV08a] L. Beirão da Veiga. A residual based error estimator for the Mimetic Finite Difference method. Numerische Mathematik, 108 (3):387–406, January 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dolejsi:2008:ABD

[DV08b] Vít Dolejší and Miloslav Vlasák. Analysis of a BDF DGFE scheme for nonlinear convection diffusion problems. Numerische Mathematik, 110(4):405–447, October 2008. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=110&issue=4&spage=
405.

daVeiga:2009:CAH

[dVLM09] L. Beirão da Veiga, K. Lipnikov, and G. Manzini. Convergence analysis of the high-order mimetic finite difference method. Numerische Mathematik, 113(3):325–356, September 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=3&spage=325.

daVeiga:2006:PDB

[dVLP06] L. Beirão da Veiga, C. Lovadina, and L. F. Pavarino. Positive definite balancing Neumann–Neumann preconditioners for nearly incompressible elasticity. Numerische Mathematik, 104(3):271–296, September 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

daVeiga:2007:PEE

[dVNS07] L. Beirão da Veiga, J. Niiranen, and R. Stenberg. A posteriori error estimates for the Morley plate bending element. Numerische Mathematik, 106(2):165–179, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dolzmann:2000:ENA

[DW00] G. Dolzmann and N. J. Walkington. Estimates for numerical approximations of rank one convex envelopes. Numerische Mathematik, 85(4):647–663, June 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085004/00850647.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085004/00850647.pdf.

Dai:2008:COC

[DXZ08] Xiaoying Dai, Jinchao Xu, and Aihui Zhou. Convergence and optimal complexity of adaptive finite element eigenvalue computations. Numerische Mathematik, 110(3):313–355, September 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=3&spage=313.

Demanet:2009:WAT

[DY09] Laurent Demanet and Lexing Ying. Wave atoms and time upscaling of wave equations. Numerische Mathematik, 113(1):1– 71, July 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=1&spage=1.

Dziuk:2008:CPW

[Dzi08] Gerhard Dziuk. Computational parametric Willmore flow. Numerische Mathematik, 111(1):55–80, November 2008. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=111&issue=
1&spage=55.

Engelborghs:2002:SPP

[ED02] K. Engelborghs and E. J. Doedel. Stability of piecewise polynomial collocation for computing periodic solutions of delay differential equations. Numerische Mathematik, 91(4):627–648, June 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eymard:2006:FVS

[EFG06] Robert Eymard, Jürgen Fuhrmann, and Klaus Gärtner. A finite volume scheme for nonlinear parabolic equations derived from one-dimensional local Dirichlet problems. Numerische Mathematik, 102(3):463–495, January 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=3&spage=463.

Eymard:2002:CFV

[EGHM02] Robert Eymard, Thierry Gallouït, Raphaèle Herbin, and Anthony Michel. Convergence of a finite volume scheme for nonlinear degenerate parabolic equations. Numerische Mathematik, 92(1):41–82, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eppler:2006:CFB

[EH06] Karsten Eppler and Helmut Harbrecht. Coupling of FEM and BEM in shape optimization. Numerische Mathematik, 104(1):47– 68, July 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ezquerro:2008:UMU

[EH08] J. A. Ezquerro and M. A. Hernández. The Ulm method under mild differentiability conditions. Numerische Mathematik, 109(2):193– 207, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eymard:2006:CFV

[EHV06] Robert Eymard, Danielle Hilhorst, and Martin Vohralk. A combined finite volume-nonconforming/mixed-hybrid finite element scheme for degenerate parabolic problems. Numerische Mathematik, 105(1):73–131, November 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Evje:2000:DAB

[EK00] Steinar Evje and Kenneth Hvistendahl Karlsen. Discrete approximations of BV solutions to doubly nonlinear degenerate parabolic equations. Numerische Mathematik, 86(3):377–417, September 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086003/00860377.htm;

 http://link.springer-ny.com/link/service/journals/00211/
papers/0086003/00860377.pdf.

Engelke:2002:IST

[EK02] Stephan Engelke and Christian Kanzow. Improved smoothingtype methods for the solution of linear programs. Numerische Mathematik, 90(3):487–507, January 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
2090003/20900487.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/2090003/20900487.pdf.

Ebmeyer:2005:QNI

[EL05] Carsten Ebmeyer and W. B. Liu. Quasi-norm interpolation error estimates for the piecewise linear finite element approximation of p-Laplacian problems. Numerische Mathematik, 100(2):233–258, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Esmaeili:2001:CAA

[EMAS01] Hamid Esmaeili, Nezam Mahdavi-Amiri, and Emilio Spedicato. A class of ABS algorithms for Diophantine linear systems. Numerische Mathematik, 90(1):101–115, November 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090001/10900101.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090001/10900101.pdf.

Engelberg:2000:NCG

[EN00] Shlomo Engelberg and Aaron E. Naiman. A note on conjugate gradient convergence — Part III. Numerische Mathematik, 85(4): 685–696, June 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085004/00850685.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085004/00850685.pdf.

Egger:2005:PLI

[EN05] Herbert Egger and Andreas Neubauer. Preconditioning Landweber iteration in Hilbert scales. Numerische Mathematik, 101(4):643–662, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=4&spage=643.

Eisinberg:2001:VMI

[EPS01] Alfredo Eisinberg, Paolo Pugliese, and Nicola Salerno. Vandermonde matrices on integer nodes: the rectangular case. Numerische Mathematik, 87(4):663–674, February 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1087004/10870663.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1087004/
10870663.pdf.

Ecevit:2009:AMS

[ER09] Fatih Ecevit and Fernando Reitich. Analysis of multiple scattering iterations for high-frequency scattering problems. I: the two-dimensional case. Numerische Mathematik, 114(2):271–354, December 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=114&issue=2&spage=271.

Ervedoza:2009:SCA

[Erv09] Sylvain Ervedoza. Spectral conditions for admissibility and observability of wave systems: applications to finite element schemes. Numerische Mathematik, 113(3):377–415, September 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=3&spage=377.

Engleder:2008:SBE

[ES08] S. Engleder and O. Steinbach. Stabilized boundary element methods for exterior Helmholtz problems. Numerische Mathematik, 110 (2):145–160, August 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=110&issue=2&spage=145.

Elman:2002:PAS

[ESW02] Howard C. Elman, David J. Silvester, and Andrew J. Wathen. Performance and analysis of saddle point preconditioners for the discrete steady-state Navier–Stokes equations. Numerische Mathematik, 90(4):665–688, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.de/link/service/journals/00211/bibs/2090004/
20900665.htm;   http://link.springer.de/link/service/
journals/00211/papers/2090004/20900665.pdf.

Eirola:2004:NTE

[EvP04] Timo Eirola and Jan von Pfaler. Numerical Taylor expansions for invariant manifolds. Numerische Mathematik, 99(1):25–46, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Efendiev:2002:MFE

[EW02] Yalchin R. Efendiev and Xiao-Hui Wu. Multiscale finite element for problems with highly oscillatory coefficients. Numerische Mathematik, 90(3):459–486, January 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
2090003/20900459.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/2090003/20900459.pdf.

Eck:2003:RBE

[EW03] C. Eck and W. L. Wendland. A residual-based error estimator for BEM-discretizations of contact problems. Numerische Mathematik, 95(2):253–282, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ervedoza:2008:PML

[EZ08] S. Ervedoza and E. Zuazua. Perfectly matched layers in 1-d: energy decay for continuous and semi-discrete waves. Numerische Mathematik, 109(4):597–634, June 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Faermann:2002:LAS

[Fae02] Birgit Faermann. Localization of the Aronszajn–Slobodeckij norm and application to adaptive boundary element methods: Part II. The three-dimensional case. Numerische Mathematik, 92(3): 467–499, September 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fiorot:2002:BAC

[FG02] Jean-Charles Fiorot and Olivier Gibaru. Blowing up: application to G^{2}-continuous 8-sided filling patch. Numerische Mathematik, 92 (2):257–287, August 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fernandez-Garcia:2002:FCP

[FGSn02] J. R. Fernández-García, M. Sofonea, and J. M. Via no. A frictionless contact problem for elastic-viscoplastic materials with normal compliance: Numerical analysis and computational experiments. Numerische Mathematik, 90(4):689– 719, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/2090004/20900689.htm;
 http://link.springer.de/link/service/journals/00211/
papers/2090004/20900689.pdf.

Floater:2007:BRI

[FH07] Michael S. Floater and Kai Hormann. Barycentric rational interpolation with no poles and high rates of approximation. Numerische Mathematik, 107(2):315–331, August 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fernandez:2003:NAE

[FHV03] José R. Fernández, Patrick Hild, and Juan M. Viaño. Numerical approximation of the elastic-viscoplastic contact problem with non-matching meshes. Numerische Mathematik, 94(3):501–522, May 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Filbet:2006:FVS

[Fil06] Francis Filbet. A finite volume scheme for the Patlak–Keller– Segel chemotaxis model. Numerische Mathematik, 104(4):457– 488, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fischer:2002:RQU

[Fis02] J.-W. Fischer. Romberg quadrature using the Bulirsch sequence. Numerische Mathematik, 90(3):509–519, January 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/2090003/20900509.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/2090003/
20900509.pdf.

Fang:2006:OQP

[FL06] Gensun Fang and Xuehua Li. Optimal quadrature problem on classes defined by kernels satisfying certain oscillation properties. Numerische Mathematik, 105(1):133–158, November 2006. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fernandez-Mendez:2003:CFE

[FMDH03] Sonia Fernández-Méndez, Pedro Díez, and Antonio Huerta. Convergence of finite elements enriched with mesh-less methods. Numerische Mathematik, 96(1):43–59, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Farhat:2000:TLD

[FML00] Charbel Farhat, Antonini Macedo, and Michel Lesoinne. A two-level domain decomposition method for the iterative solution of high frequency exterior Helmholtz problems. Numerische Mathematik, 85(2):283–308, April 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085002/00850283.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085002/00850283.pdf.

Frenod:2007:TSN

[FMS07] Emmanuel Frénod, Alexandre Mouton, and Eric Sonnendrücker. Two-scale numerical simulation of the weakly compressible 1D isentropic Euler equations. Numerische Mathematik, 108(2):263– 293, December 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Feng:2007:EAF

[FNP07] Xiaobing Feng, Michael Neilan, and Andreas Prohl. Error analysis of finite element approximations of the inverse mean curvature flow arising from the general relativity. Numerische Mathematik, 108 (1):93–119, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fang:2009:PEE

[FO09] F. Fang and C. W. Oosterlee. Pricing early-exercise and discrete barrier options by Fourier-cosine series expansions. Numerische Mathematik, 114(1):27–62, November 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?

genre=article&issn=0029-599X&volume=114&issue=1&spage=
27.

Foucart:2009:ACF

[Fou09] Simon Foucart. Allometry constants of finite-dimensional spaces: theory and computations. Numerische Mathematik, 112(4):535– 564, June 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=4&spage=535.

Formaggia:2001:NAP

[FP01] L. Formaggia and S. Perotto. New anisotropic a priori error estimates. Numerische Mathematik, 89(4):641–667, October 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089004/10890641.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089004/10890641.pdf.

Feng:2003:NAA

[FP03a] Xiaobing Feng and Andreas Prohl. Numerical analysis of the Allen–Cahn equation and approximation for mean curvature flows. Numerische Mathematik, 94(1):33–65, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Formaggia:2003:AEE

[FP03b] L. Formaggia and S. Perotto. Anisotropic error estimates for elliptic problems. Numerische Mathematik, 94(1):67–92, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Feng:2004:EAM

[FP04] Xiaobing Feng and Andreas Prohl. Error analysis of a mixed finite element method for the Cahn–Hilliard equation. Numerische Mathematik, 99(1):47–84, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ferreira:2009:CLA

[FP09] Carla Ferreira and Beresford Parlett. Convergence of LR algorithm for a one-point spectrum tridiagonal matrix. Numerische Mathematik, 113(3):417–431, September 2009. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=113&issue=
3&spage=417.

Furihata:2001:SCF

[Fur01] Daisuke Furihata. A stable and conservative finite difference scheme for the Cahn–Hilliard equation. Numerische Mathematik, 87(4):675–699, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087004/10870675.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087004/10870675.pdf.

Fierro:2006:PEE

[FV06] Francesca Fierro and Andreas Veeser. A posteriori error estimators, gradient recovery by averaging, and superconvergence. Numerische Mathematik, 103(2):267–298, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=103&issue=2&spage=267.

Feng:2005:RCR

[FvOP05] Xiaobing Feng, Markus von Oehsen, and Andreas Prohl. Rate of convergence of regularization procedures and finite element approximations for the total variation flow. Numerische Mathematik, 100(3):441–456, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gallagher:2002:MAS

[Gal02] Isabelle Gallagher. Mathematical analysis of a structurepreserving approximation of the bidimensional vorticity equation. Numerische Mathematik, 91(2):223–236, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gallice:2003:PES

[Gal03] Gérard Gallice. Positive and entropy stable Godunov-type schemes for gas dynamics and MHD equations in Lagrangian or Eulerian coordinates. Numerische Mathematik, 94(4): 673–713, June 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=94&issue=4&spage=673.

Gautschi:2001:NBF

[Gau01] Walter Gautschi. Note: Barycentric formulae for cardinal (SINC-) interpolants by Jean-Paul Berrut. Numerische Mathematik, 87(4): 791–792, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1087004/10870791.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1087004/10870791.pdf.

Gautschi:2005:HNG

[Gau05] Walter Gautschi. A historical note on Gauss–Kronrod quadrature. Numerische Mathematik, 100(3):483–484, May 2005. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Geissert:2007:ADM

[Gei07] Matthias Geissert. Applications of discrete maximal L_{p} regularity for finite element operators. Numerische Mathematik, 108(1): 121–149, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gerbeau:2000:SFE

[Ger00] J.-F. Gerbeau. A stabilized finite element method for the incompressible magnetohydrodynamic equations. Numerische Mathematik, 87(1):83–111, November 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0087001/00870083.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0087001/00870083.pdf.

Galiano:2003:SDT

[GGJ03] Gonzalo Galiano, María L. Garzón, and Ansgar Jüngel. Semidiscretization in time and numerical convergence of solutions of a nonlinear cross-diffusion population model. Numerische Mathematik, 93(4):655–673, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:2000:QRR

[GGL00] Walter Gautschi, Laura Gori, and M. Laura Lo Cascio. Quadrature rules for rational functions. Numerische Mathematik, 86(4): 617–633, October 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0086004/00860617.

htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0086004/00860617.pdf.

Girault:2001:BMF

[GGLV01] V. Girault, R. Glowinski, H. López, and J.-P. Vila. A boundary multiplier/fictitious domain method for the steady incompressible Navier–Stokes equations. Numerische Mathematik, 88(1):75– 103, March 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088001/10880075.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088001/10880075.pdf.

Gaiffe:2002:DDS

[GGM02] Stéphanie Gaiffe, Roland Glowinski, and Roland Masson. Domain decomposition and splitting methods for Mortar mixed finite element approximations to parabolic equations. Numerische Mathematik, 93(1):53–75, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gaudioso:2009:PLA

[GGM09] M. Gaudioso, E. Gorgone, and M. F. Monaco. Piecewise linear approximations in nonconvex nonsmooth optimization. Numerische Mathematik, 113(1):73–88, July 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=1&spage=
73.

Guillen-Gonzalez:2005:BFE

[GGRG05] F. Guillén-González and D. Rodríguez-Gómez. Bubble finite elements for the primitive equations of the ocean. Numerische Mathematik, 101(4):689–728, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=4&spage=689.

Gradinaru:2003:MFE

[GH03] V. Gradinaru and R. Hiptmair. Mixed finite elements on sparse grids. Numerische Mathematik, 93(3):471–495, January 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guo:2004:ORC

[GH04] Benqi Guo and Nobert Heuer. The optimal rate of convergence of the p-version of the boundary element method in two dimensions. Numerische Mathematik, 98(3):499–538, September 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Galantai:2008:PBP

[GH08] A. Galántai and C. J. Hegedüs. Perturbation bounds for polynomials. Numerische Mathematik, 109(1):77–100, March 2008. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gavrilyuk:2002:MAO

[GHK02] Ivan P. Gavrilyuk, Wolfgang Hackbusch, and Boris N. Khoromskij. H-matrix approximation for the operator exponential with applications. Numerische Mathematik, 92(1):83–111, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Glowinski:2005:FEA

[GHL^{+}05] Roland Glowinski, Jiwen He, Alexei Lozinski, Jacques Rappaz, and Joël Wagner. Finite element approximation of multi-scale elliptic problems using patches of elements. Numerische Mathematik, 101(4):663–687, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=4&spage=663.

Gatica:2009:RBP

[GHM09] Gabriel N. Gatica, George C. Hsiao, and Salim Meddahi. A residual-based a posteriori error estimator for a two-dimensional fluid solid interaction problem. Numerische Mathematik, 114(1): 63–106, November 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=114&issue=1&spage=63.

Grikurov:2003:CSM

[GHNP03] Valery E. Grikurov, Erkki Heikkola, Pekka Neittaanmäki, and Boris A. Plamenevskii. On computation of scattering matrices and on surface waves for diffraction gratings. Numerische Mathematik, 94(2):269–288, April 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Graham:2000:DBE

[GHS00a] I. G. Graham, W. Hackbusch, and S. A. Sauter. Discrete boundary element methods on general meshes in 3D. Numerische Mathematik, 86(1):103–137, July 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0086001/00860103.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0086001/00860103.pdf.

Graham:2000:HGB

[GHS00b] I. G. Graham, W. Hackbusch, and S. A. Sauter. Hybrid Galerkin boundary elements: theory and implementation. Numerische Mathematik, 86(1):139–172, July 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0086001/00860139.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0086001/00860139.pdf.

Gosse:2002:CRI

[GJ02] Laurent Gosse and François James. Convergence results for an inhomogeneous system arising in various high frequency approximations. Numerische Mathematik, 90(4):721– 753, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/2090004/20900721.htm;
 http://link.springer.de/link/service/journals/00211/
papers/2090004/20900721.pdf.

Grune:2001:HON

[GK01] L. Grüne and P. E. Kloeden. Higher order numerical schemes for affinely controlled nonlinear systems. Numerische Mathematik, 89 (4):669–690, October 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1089004/10890669.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1089004/10890669.pdf.

Giladi:2002:STD

[GK02] Eldar Giladi and Herbert B. Keller. Space–time domain decomposition for parabolic problems. Numerische Mathematik, 93(2): 279–313, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gopalakrishnan:2003:EMD

[GK03a] J. Gopalakrishnan and G. Kanschat. Erratum: A multilevel discontinuous Galerkin method. Numerische Mathematik, 95(3):551, September 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [GK03b].

Gopalakrishnan:2003:MDG

[GK03b] J. Gopalakrishnan and G. Kanschat. A multilevel discontinuous Galerkin method. Numerische Mathematik, 95(3):527–550, September 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [GK03a].

Grasedyck:2009:DDB

[GKB09] Lars Grasedyck, Ronald Kriemann, and Sabine Le Borne. Domain decomposition based HH-LU preconditioning. Numerische Mathematik, 112(4):565–600, June 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=112&issue=4&spage=
565.

Glader:2005:NAM

[GL05] Christer Glader and Mikael Lindström. A new algorithm for meromorphic Nevanlinna–Pick interpolation. Numerische Mathematik, 100(1):49–69, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gervasio:2001:HCV

[GLQ01] P. Gervasio, J.-L. Lions, and A. Quarteroni. Heterogeneous coupling by virtual control methods. Numerische Mathematik, 90(2): 241–264, December 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1090002/10900241.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1090002/10900241.pdf.

Guzman:2009:HEG

[GLRS09] J. Guzmán, D. Leykekhman, J. Rossmann, and A. H. Schatz. Hölder estimates for Green’s functions on convex polyhedral domains and their applications to finite element methods. Numerische Mathematik, 112(2):221–243, April 2009. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=112&issue=
2&spage=221.

Giraud:2005:REA

[GLRvdE05] Luc Giraud, Julien Langou, Miroslav Rozložník, and Jasper van den Eshof. Rounding error analysis of the classical Gram–Schmidt orthogonalization process. Numerische Mathematik, 101(1):87–100, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=1&spage=87.

Graham:2007:DDM

[GLS07] I. G. Graham, P. O. Lechner, and R. Scheichl. Domain decomposition for multiscale PDEs. Numerische Mathematik, 106(4):589– 626, June 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guo:2006:SPD

[GLX06] Xiao-Xia Guo, Wen-Wei Lin, and Shu-Fang Xu. A structurepreserving doubling algorithm for nonsymmetric algebraic Riccati equation. Numerische Mathematik, 103(3):393–412, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=103&issue=3&spage=
393.

Gatica:2005:CMF

[GM05] Gabriel N. Gatica and Salim Meddahi. On the coupling of MIXED–FEM and BEM for an exterior Helmholtz problem in the plane. Numerische Mathematik, 100(4):663–695, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=100&issue=4&spage=
663.

Gia:2006:POL

[GM06] Q. T. Le Gia and H. N. Mhaskar. Polynomial operators and local approximation of solutions of pseudo-differential equations on the sphere. Numerische Mathematik, 103(2):299–322, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=103&issue=2&spage=

Guillaume:2002:CIM

[GMZ02] Philippe Guillaume, Mohamed Masmoudi, and Anis Zeglaoui. From compressible to incompressible materials via an asymptotic expansion. Numerische Mathematik, 91(4):649–673, June 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gudi:2008:HDG

[GNP08] Thirupathi Gudi, Neela Nataraj, and Amiya K. Pani. hp-Discontinuous Galerkin methods for strongly nonlinear elliptic boundary value problems. Numerische Mathematik, 109(2):233– 268, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gonzalez-Pinto:2003:DSS

[GP03] S. González-Pinto. Differential systems with semi-stable equilibria and numerical methods. Numerische Mathematik, 96(2):253–268, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guermond:2008:MMH

[GP08] Jean-Luc Guermond and Bojan Popov. L^{1}-minimization methods for Hamilton–Jacobi equations: the one-dimensional case. Numerische Mathematik, 109(2):269–284, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gallardo:2005:GDM

[GPC05] José M. Gallardo, Carlos Parés, and Manuel Castro. A generalized duality method for solving variational inequalities. Applications to some nonlinear Dirichlet problems. Numerische Mathematik, 100 (2):259–291, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gonzalez-Pinto:2004:SRK

[GPHA04] S. González-Pinto and D. Hernández-Abreu. Stable Runge–Kutta integrations for differential systems with semi-stable equilibria. Numerische Mathematik, 97(3):473–491, May 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldfeld:2003:BNN

[GPW03] Paulo Goldfeld, Luca F. Pavarino, and Olof B. Widlund. Balancing Neumann–Neumann preconditioners for mixed approximations of heterogeneous problems in linear elasticity. Numerische Mathematik, 95(2):283–324, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grun:2000:NPC

[GR00] Günther Grün and Martin Rumpf. Nonnegativity preserving convergent schemes for the thin film equation. Numerische Mathematik, 87(1):113–152, November 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0087001/00870113.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0087001/00870113.pdf.

Godlewski:2004:NIC

[GR04] E. Godlewski and P.-A. Raviart. The numerical interface coupling of nonlinear hyperbolic systems of conservation laws: I. The scalar case. Numerische Mathematik, 97(1):81–130, March 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:2005:SPS

[Gri05a] R. D. Grigorieff. Superapproximation for projections on spline spaces. Numerische Mathematik, 99(4):657–668, February 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grimm:2005:NGT

[Gri05b] Volker Grimm. A note on the Gautschi-type method for oscillatory second-order differential equations. Numerische Mathematik, 102(1):61–66, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=1&spage=61.

Grimm:2005:EBG

[Gri05c] Volker Grimm. On error bounds for the Gautschi-type exponential integrator applied to oscillatory second-order differential equations. Numerische Mathematik, 100(1):71–89, March 2005. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grohs:2009:SEP

[Gro09] Philipp Grohs. Smoothness equivalence properties of univariate subdivision schemes and their projection analogues. Numerische Mathematik, 113(2):163–180, August 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=2&spage=
163.

Grune:2004:EEA

[Grü04] Lars Grüne. Error estimation and adaptive discretization for the discrete stochastic Hamilton–Jacobi–Bellman equation. Numerische Mathematik, 99(1):85–112, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guo:2000:LGM

[GS00] Ben-Yu Guo and Jie Shen. Laguerre–Galerkin method for nonlinear partial differential equations on a semi-infinite interval. Numerische Mathematik, 86(4):635–654, October 2000. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0086004/00860635.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0086004/
00860635.pdf.

Graham:2002:FDS

[GS02] I. G. Graham and I. H. Sloan. Fully discrete spectral boundary integral methods for Helmholtz problems on smooth closed surfaces in R^{3}. Numerische Mathematik, 92(2):289–323, August 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grau-Sanchez:2007:IES

[GS07] Miquel Grau-Sánchez. Improvements of the efficiency of some three-step iterative like-Newton methods. Numerische Mathematik, 107(1):131–146, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gatica:2008:CIS

[GS08] Gabriel N. Gatica and Francisco-Javier Sayas. Characterizing the inf-sup condition on product spaces. Numerische Mathematik, 109 (2):209–231, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gil:2004:IRC

[GST04] Amparo Gil, Javier Segura, and Nico M. Temme. Integral representations for computing real parabolic cylinder functions.

Numerische Mathematik, 98(1):105–134, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gosse:2004:APW

[GT04] Laurent Gosse and Giuseppe Toscani. Asymptotic-preserving & well-balanced schemes for radiative transfer and the Rosseland approximation. Numerische Mathematik, 98(2):223–250, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garvie:2007:FEA

[GT07] Marcus R. Garvie and Catalin Trenchea. Finite element approximation of spatially extended predator-prey interactions with the Holling type II functional response. Numerische Mathematik, 107 (4):641–667, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garcke:2005:NAC

[GW05] Harald Garcke and Ulrich Weikard. Numerical approximation of the Cahn–Larché equation. Numerische Mathematik, 100(4):639–662, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=100&issue=4&spage=639.

Girault:2008:NDD

[GW08] V. Girault and M. F. Wheeler. Numerical discretization of a Darcy Forchheimer model. Numerische Mathematik, 110(2):161– 198, August 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=2&spage=161.

Guglielmi:2001:ZSV

[GZ01] N. Guglielmi and M. Zennaro. On the zero-stability of variable stepsize multistep methods: the spectral radius approach. Numerische Mathematik, 88(3):445–458, May 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088003/10880445.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088003/
10880445.pdf.

Hackbusch:2007:ACI

[Hac07] Wolfgang Hackbusch. Approximation of coalescence integrals in population balance models with local mass conservation. Numerische Mathematik, 106(4):627–657, June 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:2008:ECN

[Hac08] W. Hackbusch. Efficient convolution with the Newton potential in d dimensions. Numerische Mathematik, 110(4):449–489, October 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=4&spage=449.

Hager:2000:RKM

[Hag00] William W. Hager. Runge–Kutta methods in optimal control and the transformed adjoint system. Numerische Mathematik, 87(2): 247–282, December 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0087002/00870247.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0087002/00870247.pdf.

Hairer:2003:GMH

[Hai03] Ernst Hairer. Global modified Hamiltonian for constrained symplectic integrators. Numerische Mathematik, 95(2):325–336, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haider:2009:SAC

[HCC09] F. Haider, J.-P. Croisille, and B. Courbet. Stability analysis of the cell centered finite-volume Muscl method on unstructured grids. Numerische Mathematik, 113(4):555–600, October 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=4&spage=555.

Dai:1999:GCM

[hDxY99] Yu hong Dai and Ya xiang Yuan. Global convergence of the method of shortest residuals. Numerische Mathematik, 83(4):581– 598, October 1999. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/9083004/90830581.htm;

 http://link.springer-ny.com/link/service/journals/00211/
papers/9083004/90830581.pdf. See comments [Pyt02].

Heider:2009:LLS

[Hei09] Pascal Heider. A local least-squares method for solving nonlinear partial differential equations of second order. Numerische Mathematik, 111(3):351–375, January 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=111&issue=3&spage=
351.

Hesse:2006:LBW

[Hes06] Kerstin Hesse. A lower bound for the worst-case cubature error on spheres of arbitrary dimension. Numerische Mathematik, 103(3):413–433, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=3&spage=413.

Heuer:1998:ISM

[Heu98] Norbert Heuer. An iterative substructuring method for the pversion of the boundary element method for hypersingular integral operators in three dimensions. Numerische Mathematik, 79(3):371–396, May 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/8079003/80790371.htm;
http://science.springer.de/nmee/bibs/8079003/80790371.

htm. See erratum [Heu01b].

Heuer:2001:ASM

[Heu01a] Norbert Heuer. Additive Schwarz method for the p-version of the boundary element method for the single layer potential operator on a plane screen. Numerische Mathematik, 88(3):485– 511, May 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088003/10880485.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088003/10880485.pdf.

Heuer:2001:EIS

[Heu01b] Norbert Heuer. Erratum: “An iterative substructuring method for the p-version of the boundary element method for hypersingular integral operators in three dimensions” Numer. Math.

(1998) 79: 371–396. Numerische Mathematik, 87(4):793–794, February 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1087004/10870793.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1087004/10870793.pdf. See [Heu98].

Han:2000:AAB

[HHW00] Houde Han, Changhong He, and Xiaonan Wu. Analysis of artificial boundary conditions for exterior boundary value problems in three dimensions. Numerische Mathematik, 85(3):367– 386, May 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085003/00850367.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085003/00850367.pdf.

Hill:2006:NSG

[Hil06] Adrian T. Hill. Nonlinear stability of general linear methods. Numerische Mathematik, 103(4):611–629, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hiptmair:2001:DHO

[Hip01] R. Hiptmair. Discrete Hodge operators. Numerische Mathematik, 90(2):265–289, December 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1090002/10900265.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1090002/10900265.pdf.

Hong:2009:ACC

[HJL09] Jialin Hong, Shanshan Jiang, and Chun Li. Accuracy of classical conservation laws for Hamiltonian PDEs under Runge– Kutta discretizations. Numerische Mathematik, 112(1):1–23, March 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=1&spage=1.

Higham:2005:NMN

[HK05] Desmond J. Higham and Peter E. Kloeden. Numerical methods for nonlinear stochastic differential equations with jumps. Numerische Mathematik, 101(1):101–119, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=101&issue=1&spage=101.

Haasdonk:2001:CSL

[HKR01] Bernard Haasdonk, Dietmar Kröner, and Christian Rohde. Convergence of a staggered Lax–Friedrichs scheme for nonlinear conservation laws on unstructured two-dimensional grids. Numerische Mathematik, 88(3):459–484, May 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088003/10880459.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088003/10880459.pdf.

Hackbusch:2007:AGB

[HKS07] Wolfgang Hackbusch, Boris N. Khoromskij, and Stefan Sauter. Adaptive Galerkin boundary element methods with panel clustering. Numerische Mathematik, 105(4):603–631, February 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:2008:AIS

[HKT08] Wolfgang Hackbusch, Boris N. Khoromskij, and Eugene E. Tyrtyshnikov. Approximate iterations for structured matrices. Numerische Mathematik, 109(3):365–383, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:2004:SMM

[HL04a] Ernst Hairer and Christian Lubich. Symmetric multistep methods over long times. Numerische Mathematik, 97(4):699–723, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

He:2004:ABT

[HL04b] Yinnian He and Kaitai Li. Asymptotic behavior and time discretization analysis for the non-stationary Navier–Stokes problem. Numerische Mathematik, 98(4):647–673, October 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hansbo:2005:LMM

[HLPS05] Peter Hansbo, Carlo Lovadina, Ilaria Perugia, and Giancarlo Sangalli. A Lagrange multiplier method for the finite element solution of elliptic interface problems using non-matching meshes.

Numerische Mathematik, 100(1):91–115, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

He:2005:MLS

[HLS05] Yinnian He, Kam-Moon Liu, and Weiwei Sun. Multi-level spectral Galerkin method for the Navier–Stokes problem I: Spatial discretization. Numerische Mathematik, 101(3):501– 522, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=3&spage=501.

He:2003:SAO

[HLW03] Bingsheng He, Li-Zhi Liao, and Shengli Wang. Self-adaptive operator splitting methods for monotone variational inequalities. Numerische Mathematik, 94(4):715–737, June 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=94&issue=4&spage=715.

He:2004:NAM

[HMMC04] Yinnian He, Huanling Miao, R. M. M. Mattheij, and Zhangxin Chen. Numerical analysis of a modified finite element nonlinear Galerkin method. Numerische Mathematik, 97(4):725–756, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Handlovicova:2003:SIC

[HMS03] Angela Handlovičová, Karol Mikula, and Fiorella Sgallari. Semiimplicit complementary volume scheme for solving level set like equations in image processing and curve evolution. Numerische Mathematik, 93(4):675–695, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:2007:PEM

[HMY07] Desmond J. Higham, Xuerong Mao, and Chenggui Yuan. Preserving exponential mean-square stability in the simulation of hybrid stochastic differential equations. Numerische Mathematik, 108(2): 295–325, December 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Huhtanen:2000:MDI

[HN00] Marko Huhtanen and Olavi Nevanlinna. Minimal decompositions and iterative methods. Numerische Mathematik, 86(2):257–281, August 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086002/00860257.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086002/00860257.pdf.

Hild:2005:PEE

[HN05] Patrick Hild and Serge Nicaise. A posteriori error estimations of residual type for Signorini’s problem. Numerische Mathematik, 101(3):523–549, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=3&spage=523.

Hansen:2008:DSE

[HO08] Eskil Hansen and Alexander Ostermann. Dimension splitting for evolution equations. Numerische Mathematik, 108(4):557–570, February 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henning:2009:HMF

[HO09] Patrick Henning and Mario Ohlberger. The heterogeneous multiscale finite element method for elliptic homogenization problems in perforated domains. Numerische Mathematik, 113(4):601–629, October 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=4&spage=601.

Holst:2008:PEE

[Hol08] Stefan Holst. An a priori error estimate for a monotone mixed finite-element discretization of a convection-diffusion problem. Numerische Mathematik, 109(1):101–119, March 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Horak:2004:CMP

[Hor04] Jiří Horák. Constrained mountain pass algorithm for the numerical solution of semilinear elliptic problems. Numerische Mathematik, 98(2):251–276, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Havu:2001:AFE

[HP01] V. Havu and J. Pitkäranta. An analysis of finite element locking in a parameter dependent model problem. Numerische Mathematik, 89(4):691–714, October 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1089004/10890691.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089004/10890691.pdf.

Harbrecht:2002:BWA

[HPPS02] H. Harbrecht, F. Paiva, C. Pérez, and R. Schneider. Biorthogonal wavelet approximation for the coupling of FEM-BEM. Numerische Mathematik, 92(2):325–356, August 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Harbrecht:2003:SRP

[HPS03] Helmut Harbrecht, Sergei Pereverzev, and Reinhold Schneider. Self-regularization by projection for noisy pseudodifferential equations of negative order. Numerische Mathematik, 95(1):123–143, July 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Houston:2005:IPM

[HPSS05] Paul Houston, Ilaria Perugia, Anna Schneebeli, and Dominik Schötzau. Interior penalty method for the indefinite timeharmonic Maxwell equations. Numerische Mathematik, 100(3): 485–518, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Han:2000:CAP

[HR00] Weimin Han and B. Daya Reddy. Convergence of approximations to the primal problem in plasticity under conditions of minimal regularity. Numerische Mathematik, 87(2):283–315, December 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087002/00870283.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087002/00870283.pdf.

Hernandez:2004:FEA

[HR04a] Erwin Hernández and Rodolfo Rodríguez. Finite element approximation of spectral acoustic problems on curved domains.

Numerische Mathematik, 97(1):131–158, March 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hintermuller:2004:LSA

[HR04b] Michael Hintermüller and Wolfgang Ring. A level set approach for the solution of a state-constrained optimal control problem. Numerische Mathematik, 98(1):135–166, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hollig:2002:MMW

[HRW02] Klaus Höllig, Ulrich Reif, and Joachim Wipper. Multigrid methods with web-splines. Numerische Mathematik, 91(2):237–255, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hohage:2005:NSH

[HS05] Thorsten Hohage and Francisco-Javier Sayas. Numerical solution of a heat diffusion problem by boundary element methods using the Laplace transform. Numerische Mathematik, 102(1):67–92, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=1&spage=67.

Hadjidimos:2006:OSI

[HS06] A. Hadjidimos and N. S. Stylianopoulos. Optimal semi-iterative methods for complex SOR with results from potential theory. Numerische Mathematik, 103(4):591–610, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hong:2008:GFM

[HS08] Jialin Hong and Yajuan Sun. Generating functions of multisymplectic RK methods via DW Hamilton Jacobi equations. Numerische Mathematik, 110(4):491–519, October 2008. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=110&issue=
4&spage=491.

Heuer:2009:CRB

[HS09a] Norbert Heuer and Francisco-Javier Sayas. Crouzeix–Raviart boundary elements. Numerische Mathematik, 112(3):381–401, May 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=3&spage=381.

Hu:2009:NPE

[HS09b] Jun Hu and Zhongci Shi. A new a posteriori error estimate for the Morley element. Numerische Mathematik, 112(1):25– 40, March 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=1&spage=25.

Harbrecht:2008:MFS

[HSS08a] Helmut Harbrecht, Reinhold Schneider, and Christoph Schwab. Multilevel frames for sparse tensor product spaces. Numerische Mathematik, 110(2):199–220, August 2008. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=110&issue=2&spage=
199.

Harbrecht:2008:SSM

[HSS08b] Helmut Harbrecht, Reinhold Schneider, and Christoph Schwab. Sparse second moment analysis for elliptic problems in stochastic domains. Numerische Mathematik, 109(3):385–414, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Humpherys:2006:ECA

[HSZ06] Jeffrey Humpherys, Björn Sandstede, and Kevin Zumbrun. Efficient computation of analytic bases in Evans function analysis of large systems. Numerische Mathematik, 103(4):631–642, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hintermuller:2008:MIS

[HTY08] M. Hintermüller, F. Tröltzsch, and I. Yousept. Mesh-independence of semismooth Newton methods for Lavrentiev-regularized state constrained nonlinear optimal control problems. Numerische Mathematik, 108(4):571–603, February 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Huang:2009:DDS

[Hua09] Chengming Huang. Delay-dependent stability of high order Runge–Kutta methods. Numerische Mathematik, 111(3):377–387, January 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=3&spage=377.

Huhtanen:2007:AOE

[Huh07] Marko Huhtanen. Averaging operators for exponential splittings. Numerische Mathematik, 106(3):511–528, May 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wang:2002:QFE

[hW02] Lie heng Wang. On the quadratic finite element approximation to the obstacle problem. Numerische Mathematik, 92(4):771–778, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Huang:2002:EBL

[HWW02] Jinyang Huang, Jinghua Wang, and Gerald Warnecke. Error bounds for the large time step Glimm scheme applied to scalar conservation laws. Numerische Mathematik, 91(1):13–34, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hiptmair:2006:ASP

[HWZ06] R. Hiptmair, G. Widmer, and J. Zou. Auxiliary space preconditioning in H_{0}(curl; ω). Numerische Mathematik, 103 (3):435–459, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=3&spage=435.

Han:2008:OSM

[HXY08] Deren Han, Wei Xu, and Hai Yang. An operator splitting method for variational inequalities with partially unknown mappings. Numerische Mathematik, 111(2):207–237, December 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=2&spage=207.

He:2008:LPF

[HXZL08] Yinnian He, Jinchao Xu, Aihui Zhou, and Jian Li. Local and parallel finite element algorithms for the stokes problem. Numerische Mathematik, 109(3):415–434, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hu:2002:TNV

[HZ02] Qiya Hu and Jun Zou. Two new variants of nonlinear inexact Uzawa algorithms for saddle-point problems. Numerische Mathematik, 93(2):333–359, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ikonen:2008:OSM

[IT08] Samuli Ikonen and Jari Toivanen. Operator splitting methods for pricing American options under stochastic volatility. Numerische Mathematik, 111(2):299–324, December 2008. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=111&issue=
2&spage=299.

Jerome:2008:TPC

[Jer08] Joseph W. Jerome. A trapping principle and convergence result for finite element approximate solutions of steady reaction/diffusion systems. Numerische Mathematik, 109(1):121–142, March 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jia:2002:CCA

[JJL02] Rong-Qing Jia, Qingtang Jiang, and S. L. Lee. Convergence of cascade algorithms in Sobolev spaces and integrals of wavelets. Numerische Mathematik, 91(3):453–473, May 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=91&issue=3&spage=453.

Jakobsen:2009:EEA

[JKC09] Espen R. Jakobsen, Kenneth H. Karlsen, and Claudia La Chioma. Error estimates for approximate solutions to Bellman equations associated with controlled jump-diffusions. Numerische Mathematik, 113(2):221–255, August 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=2&spage=
221.

Jentzen:2009:PAS

[JKN09] A. Jentzen, P. E. Kloeden, and A. Neuenkirch. Pathwise approximation of stochastic differential equations on domains: higher order convergence rates without global Lipschitz coefficients. Numerische Mathematik, 112(1):41–64, March 2009. CODEN NUMMA7. ISSN ???? URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=112&issue=1&spage=41.

Jahnke:2003:NIQ

[JL03] Tobias Jahnke and Christian Lubich. Numerical integrators for quantum dynamics close to the adiabatic limit. Numerische Mathematik, 94(2):289–314, April 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:2000:IOA

[Joh00] Michael J. Johnson. An improved order of approximation for thin-plate spline interpolation in the unit disc. Numerische Mathematik, 84(3):451–474, January 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0084003/00840451.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0084003/00840451.pdf.

Johnson:2004:EAR

[Joh04] Michael J. Johnson. An error analysis for radial basis function interpolation. Numerische Mathematik, 98(4):675–694, October 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jin:2009:DPS

[JT09] Qinian Jin and Ulrich Tautenhahn. On the discrepancy principle for some Newton type methods for solving nonlinear inverse problems. Numerische Mathematik, 111(4):509–558, February 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=4&spage=509.

Jungel:2005:DMM

[JU05] Ansgar Jüngel and Andreas Unterreiter. Discrete minimum and maximum principles for finite element approximations of nonmonotone elliptic equations. Numerische Mathematik, 99(3):485– 508, January 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jovanovic:2001:CFD

[JV01] B. S. Jovanovic and L. G. Vulkov. On the convergence of finite difference schemes for the heat equation with concentrated capacity. Numerische Mathematik, 89(4):715–734, October 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089004/10890715.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089004/10890715.pdf.

Junk:2009:CLB

[JY09] Michael Junk and Zhaoxia Yang. Convergence of lattice Boltzmann methods for Navier–Stokes flows in periodic and bounded domains. Numerische Mathematik, 112(1):65–87, March 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=1&spage=65.

Kanzow:2001:SFE

[Kan01] Christian Kanzow. Strictly feasible equation-based methods for mixed complementarity problems. Numerische Mathematik, 89(1): 135–160, July 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089001/10890135.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089001/10890135.pdf.

Kersey:2003:NI

[Ker03] Scott N. Kersey. Near-interpolation. Numerische Mathematik, 94 (3):523–540, May 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kim:2005:EEM

[Kim05a] Kwang Y. Kim. Error estimates for a mixed finite volume method for the p-Laplacian problem. Numerische Mathematik, 101(1):121–142, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=101&issue=1&spage=121.

Kimn:2005:CTO

[Kim05b] Jung-Han Kimn. A convergence theory for an overlapping Schwarz algorithm using discontinuous iterates. Numerische Mathematik, 100(1):117–139, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Korotov:2000:FEA

[KK00] Sergey Korotov and Michal Křížek. Finite element analysis of variational crimes for a quasilinear elliptic problem in 3D. Numerische Mathematik, 84(4):549–576, February 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0084004/00840549.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0084004/00840549.pdf.

Karatson:2005:DMP

[KK05] J. Karátson and S. Korotov. Discrete maximum principles for finite element solutions of nonlinear elliptic problems with mixed boundary conditions. Numerische Mathematik, 99(4):669–698, February 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kellogg:2000:AFE

[KL00a] Bruce Kellogg and Biyue Liu. The analysis of a finite element method for the Navier–Stokes equations with compressibility. Numerische Mathematik, 87(1):153–170, November 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087001/00870153.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087001/00870153.pdf.

Koch:2000:CBM

[KL00b] Tino Koch and Jörg Liesen. The conformal ‘bratwurst’ maps and associated Faber polynomials. Numerische Mathematik, 86(1): 173–191, July 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086001/00860173.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086001/00860173.pdf.

Kugler:2003:MVI

[KL03] P. Kügler and A. Leitão. Mean value iterations for nonlinear elliptic Cauchy problems. Numerische Mathematik, 96(2):269–293, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kirsch:2002:FEM

[KM02] Andreas Kirsch and Peter Monk. A finite element method for approximating electromagnetic scattering from a conducting object. Numerische Mathematik, 92(3):501–534, September 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Krogstad:2009:GPC

[KMKZ09] Stein Krogstad, Hans Z. Munthe-Kaas, and Antonella Zanna. Generalized polar coordinates on Lie groups and numerical integrators. Numerische Mathematik, 112(2):161–187, April 2009. CODEN NUMMA7. ISSN ???? URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=112&issue=2&spage=161.

Karlsen:2009:CFV

[KMR09] Kenneth Hvistendahl Karlsen, Siddhartha Mishra, and Nils Henrik Risebro. Convergence of finite volume schemes for triangular systems of conservation laws. Numerische Mathematik, 111(4): 559–589, February 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=4&spage=559.

Kunkel:2004:SCU

[KMS04] Peter Kunkel, Volker Mehrmann, and Ronald Stöver. Symmetric collocation for unstructured nonlinear differential-algebraic equations of arbitrary index. Numerische Mathematik, 98(2):277–304, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kirkland:2008:OCN

[KNS08] Stephen J. Kirkland, Michael Neumann, and Nung-Sing Sze. On optimal condition numbers for Markov chains. Numerische Mathematik, 110(4):521–537, October 2008. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=110&issue=4&spage=
521.

Kim:2009:NVM

[KNWN09] Myoungnyoun Kim, Mitsuhiro T. Nakao, Yoshitaka Watanabe, and Takaaki Nishida. A numerical verification method of bifurcating solutions for 3-dimensional Rayleigh Bénard problems. Numerische Mathematik, 111(3):389–406, January 2009. CODEN NUMMA7. ISSN ???? URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=111&issue=3&spage=389.

Kobelkov:2000:EPU

[KO00] Georgij M. Kobelkov and Maxim A. Olshanskii. Effective preconditioning of Uzawa type schemes for a generalized Stokes problem. Numerische Mathematik, 86(3):443–470, September 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086003/00860443.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086003/00860443.pdf.

Koch:2005:ACE

[Koc05] Othmar Koch. Asymptotically correct error estimation for collocation methods applied to singular boundary value problems. Numerische Mathematik, 101(1):143–164, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=101&issue=1&spage=143.

Kornhuber:2002:CNL

[Kor02] Ralf Kornhuber. On constrained Newton linearization and multigrid for variational inequalities. Numerische Mathematik, 91(4): 699–721, June 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kruzik:2001:YMA

[KP01a] Martin Kruzík and Andreas Prohl. Young measure approximation in micromagnetics. Numerische Mathematik, 90(2):291–307, December 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090002/10900291.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090002/10900291.pdf.

Kurganov:2001:TOS

[KP01b] Alexander Kurganov and Guergana Petrova. A third-order semidiscrete genuinely multidimensional central scheme for hyperbolic conservation laws and related problems. Numerische Mathematik, 88(4):683–729, June 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1088004/10880683.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1088004/10880683.pdf.

Kroger:2008:STS

[KP08] Tim Kröger and Tobias Preusser. Stability of the 8-tetrahedra shortest-interior-edge partitioning method. Numerische Mathematik, 109(3):435–457, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kvasov:2003:LTP

[KPS03] Dmitri E. Kvasov, Clara Pizzuti, and Yaroslav D. Sergeyev. Local tuning and partition strategies for diagonal GO methods. Numerische Mathematik, 94(1):93–106, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kunik:2004:KSR

[KQW04] Matthias Kunik, Shamsul Qamar, and Gerald Warnecke. Kinetic schemes for the relativistic gas dynamics. Numerische Mathematik, 97(1):159–191, March 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kim:2009:SAS

[KR09] Sun-Mi Kim and Lothar Reichel. Sensitivity analysis for Szeg polynomials. Numerische Mathematik, 113(2):265–279, August 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=2&spage=265.

Kressner:2006:PKS

[Kre06] Daniel Kressner. A periodic Krylov–Schur algorithm for large matrix products. Numerische Mathematik, 103(3): 461–483, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=3&spage=461.

Kressner:2009:BNM

[Kre09] Daniel Kressner. A block Newton method for nonlinear eigenvalue problems. Numerische Mathematik, 111(3):355–372, January 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=3&spage=355.

Knabner:2001:IIM

[KS01] Peter Knabner and Gerhard Summ. The invertibility of the isoparametric mapping for pyramidal and prismatic finite elements. Numerische Mathematik, 88(4):661–681, June 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088004/10880661.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088004/
10880661.pdf.

Kaltenbacher:2002:MGM

[KS02a] Barbara Kaltenbacher and Josef Schicho. A multi-grid method with a priori and a posteriori level choice for the regularization of nonlinear ill-posed problems. Numerische Mathematik, 93(1): 77–107, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kaltenbacher:2002:SPV

[KS02b] Barbara Kaltenbacher and Joachim Schöberl. A saddle point variational formulation for projection-regularized parameter identification. Numerische Mathematik, 91(4):675–697, June 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kunkel:2002:SCM

[KS02c] Peter Kunkel and Ronald Stöver. Symmetric collocation methods for linear differential-algebraic boundary value problems. Numerische Mathematik, 91(3):475–501, May 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=91&issue=3&spage=475.

Krebs:2007:PVF

[KS07] Andreas Krebs and Ernst P. Stephan. A p-version finite element method for nonlinear elliptic variational inequalities in 2D. Numerische Mathematik, 105(3):457–480, January 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kugler:2005:CRA

[Küg05] Philipp Kügler. Convergence rate analysis of a derivative free Landweber iteration for parameter identification in certain elliptic PDEs. Numerische Mathematik, 101(1):165–184, July 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=101&issue=1&spage=
165.

Kunert:2000:PRE

[Kun00] Gerd Kunert. An a posteriori residual error estimator for the finite element method on anisotropic tetrahedral meshes. Numerische Mathematik, 86(3):471–490, September 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0086003/00860471.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0086003/00860471.pdf.

Kuther:2003:PEE

[Küt03] Marc Küther. A priori error estimates for approximate solutions to convex conservation laws. Numerische Mathematik, 93(4):697– 727, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kunert:2000:ERD

[KV00] G. Kunert and R. Verfürth. Edge residuals dominate a posteriori error estimates for linear finite element methods on anisotropic triangular and tetrahedral meshes. Numerische Mathematik, 86 (2):283–303, August 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0086002/00860283.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0086002/00860283.pdf.

Kunisch:2001:GPO

[KV01] K. Kunisch and S. Volkwein. Galerkin proper orthogonal decomposition methods for parabolic problems. Numerische Mathematik, 90(1):117–148, November 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1090001/10900117.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1090001/10900117.pdf.

Klausen:2006:RCM

[KW06] Runhild A. Klausen and Ragnar Winther. Robust convergence of multi point flux approximation on rough grids. Numerische Mathematik, 104(3):317–337, September 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kumar:2008:CASb

[KW08a] Jitendra Kumar and Gerald Warnecke. Convergence analysis of sectional methods for solving breakage population balance equations — I: the fixed pivot technique. Numerische Mathematik, 111 (1):81–108, November 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=1&spage=81.

Kumar:2008:CAS

[KW08b] Jitendra Kumar and Gerald Warnecke. Convergence analysis of sectional methods for solving breakage population balance equations — II: the cell average technique. Numerische Mathematik, 110(4):539–559, October 2008. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=110&issue=4&spage=
539.

Kweon:2000:OOC

[Kwe00] Jae Ryong Kweon. An optimal order convergence for a weak formulation of the compressible Stokes system with inflow boundary condition. Numerische Mathematik, 86(2):305–318, August 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086002/00860305.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086002/00860305.pdf.

Lacoste:2000:SME

[Lac00] Patrick Lacoste. Solution of Maxwell equation in axisymmetric geometry by Fourier series decomposition and by use of H (rot) conforming finite element. Numerische Mathematik, 84(4):577–609, February 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0084004/00840577.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0084004/00840577.pdf.

Leborgne:2002:OCS

[Leb02] G. Leborgne. An optimally consistent stabilization of the infsup condition. Numerische Mathematik, 91(1):35–56, March 2002.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Leykekhman:2004:PLE

[Ley04] Dmitriy Leykekhman. Pointwise localized error estimates for parabolic finite element equations. Numerische Mathematik, 96 (3):583–600, January 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lopez-Fernandez:2005:FRK

[LFLPS05] María López-Fernández, Christian Lubich, Cesar Palencia, and Achim Schädle. Fast Runge–Kutta approximation of inhomogeneous parabolic equations. Numerische Mathematik, 102 (2):277–291, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=2&spage=277.

Liang:2007:RCB

[LHJB07] Jin Liang, Bei Hu, Lishang Jiang, and Baojun Bian. On the rate of convergence of the binomial tree scheme for American options. Numerische Mathematik, 107(2):333–352, August 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:2001:MTM

[Li01] Zhiping Li. A mesh transformation method for computing microstructures. Numerische Mathematik, 89(3):511–533, September 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089003/10890511.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089003/10890511.pdf.

Li:2005:DPF

[Li05]  Jing Li.  A Dual–Primal FETI method for incompress-

ible Stokes equations. Numerische Mathematik, 102(2):257– 275, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=2&spage=257.

Linss:2008:RCC

[Lin08] Torsten Linß. Robust convergence of a compact fourthorder finite difference scheme for reaction diffusion problems. Numerische Mathematik, 111(2):239–249, December 2008. CODEN NUMMA7. ISSN ???? URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=111&issue=2&spage=239.

Lin:2009:DGL

[Lin09] Runchang Lin. Discontinuous Galerkin least-squares finite element methods for singularly perturbed reaction-diffusion problems with discontinuous coefficients and boundary singularities. Numerische Mathematik, 112(2):295–318, April 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=112&issue=2&spage=
295.

Liu:2000:FEA

[Liu00] Wenbin Liu. Finite element approximation of a nonlinear elliptic equation arising from bimaterial problems in elastic-plastic mechanics. Numerische Mathematik, 86(3):491–506, September 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086003/00860491.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086003/00860491.pdf.

Lazarov:2007:BPS

[LLP07] Raytcho D. Lazarov, Shuai Lu, and Sergei V. Pereverzev. On the balancing principle for some problems of numerical analysis. Numerische Mathematik, 106(4):659–689, June 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:2003:NCG

[LLW03] Zhilin Li, Tao Lin, and Xiaohui Wu. New Cartesian grid methods for interface problems using the finite element formulation. Numerische Mathematik, 96(1):61–98, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeTallec:2000:NAL

[LM00] Patrick Le Tallec and Saloua Mani. Numerical analysis of a linearised fluid-structure interaction problem. Numerische Mathematik, 87(2):317–354, December 2000. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0087002/00870317.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0087002/00870317.pdf.

Lyness:2005:AET

[LM05] J. N. Lyness and G. Monegato. Asymptotic expansions for twodimensional hypersingular integrals. Numerische Mathematik, 100 (2):293–329, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Larson:2008:PEE

[LM08] Mats G. Larson and Axel Målqvist. A posteriori error estimates for mixed finite element approximations of elliptic problems. Numerische Mathematik, 108(3):487–500, January 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lions:2002:WPA

[LMV02] Jacques-Louis Lions, Jérôme Métral, and Olivier Vacus. Wellposed absorbing layer for hyperbolic problems. Numerische Mathematik, 92(3):535–562, September 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Larson:2004:AFD

[LN04a] Mats G. Larson and A. Jonas Niklasson. Analysis of a family of discontinuous Galerkin methods for elliptic problems: the one dimensional case. Numerische Mathematik, 99(1):113–130, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lu:2004:SNC

[LN04b] Linzhang Lu and Michael K. Ng. On sufficient and necessary conditions for the Jacobi matrix inverse eigenvalue problem. Numerische Mathematik, 98(1):167–176, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lutterkort:2001:TLE

[LP01] David Lutterkort and Jörg Peters. Tight linear envelopes for splines. Numerische Mathematik, 89(4):735–748, October 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089004/10890735.htm;

 http://link.springer-ny.com/link/service/journals/00211/
papers/1089004/10890735.pdf.

Liesen:2008:NSP

[LP08] Jörg Liesen and Beresford N. Parlett. On nonsymmetric saddle point matrices that allow conjugate gradient iterations. Numerische Mathematik, 108(4):605–624, February 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lee:2009:DIS

[LP09a] Chang-Ock Lee and Eun-Hee Park. A dual iterative substructuring method with a penalty term. Numerische Mathematik, 112 (1):89–113, March 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=112&issue=1&spage=89.

Lu:2009:SRS

[LP09b] Shuai Lu and Sergei V. Pereverzev. Sparse recovery by the standard Tikhonov method. Numerische Mathematik, 112(3):403– 424, May 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=3&spage=403.

Lazarov:2003:AOI

[LPSV03] R. D. Lazarov, J. E. Pasciak, J. Schöberl, and P. S. Vassilevski. Almost optimal interior penalty discontinuous approximations of symmetric elliptic problems on non-matching grids. Numerische Mathematik, 96(2):295–315, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lechleiter:2009:TGC

[LR09] Armin Lechleiter and Andreas Rieder. Towards a general convergence theory for inexact Newton regularizations. Numerische Mathematik, 112(3):521–548, May 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=112&issue=3&spage=
521.

Lamichhane:2006:CIL

[LRW06] B. P. Lamichhane, B. D. Reddy, and B. I. Wohlmuth. Convergence in the incompressible limit of finite element approximations based on the Hu–Washizu formulation. Numerische Mathematik, 104 (2):151–175, August 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lai:2001:MES

[LS01a] Ming-Jun Lai and Larry L. Schumaker. Macro-elements and stable local bases for splines on Clough–Tocher triangulations. Numerische Mathematik, 88(1):105–119, March 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088001/10880105.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088001/
10880105.pdf.

Lattanzio:2001:CRS

[LS01b] Corrado Lattanzio and Denis Serre. Convergence of a relaxation scheme for hyperbolic systems of conservation laws. Numerische Mathematik, 88(1):121–134, March 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088001/10880121.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088001/10880121.pdf.

Linss:2001:SSM

[LS01c] Torsten Linß and Martin Stynes. The SDFEM on Shishkin meshes for linear convection-diffusion problems. Numerische Mathematik, 87(3):457–484, January 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087003/10870457.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087003/10870457.pdf.

Laliena:2008:TAA

[LS08] Antonio R. Laliena and Francisco-Javier Sayas. Theoretical aspects of the application of convolution quadrature to scattering of acoustic waves. Numerische Mathematik, 110(4):637–678, October 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=4&spage=637.

Loisel:2009:GCO

[LS09] Sébastien Loisel and Daniel B. Szyld. On the geometric convergence of optimized Schwarz methods with applications to elliptic problems. Numerische Mathematik, 112(4):697–728, June 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=112&issue=4&spage=697.

Leblond:2002:WRA

[LSW02] J. Leblond, E. B. Saff, and F. Wielonsky. Weighted H^{2} rational approximation and consistency. Numerische Mathematik, 90(3):521– 553, January 2002. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/2090003/20900521.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/2090003/20900521.pdf.

Lamichhane:2005:HOM

[LSW05] B. P. Lamichhane, R. P. Stevenson, and B. I. Wohlmuth. Higher order mortar finite element methods in 3D with dual Lagrange multiplier bases. Numerische Mathematik, 102(1): 93–121, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=1&spage=93.

Lipnikov:2009:LFM

[LSY09] Konstantin Lipnikov, Mikhail Shashkov, and Ivan Yotov. Local flux mimetic finite difference methods. Numerische Mathematik, 112(1):115–152, March 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=112&issue=1&spage=
115.

Lin:2001:SEE

[LT01] Chi-Tien Lin and Eitan Tadmor. L^{1}-stability and error estimates for approximate Hamilton–Jacobi solutions. Numerische Mathematik, 87(4):701–735, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087004/10870701.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087004/10870701.pdf.

Lui:2002:LMI

[Lui02] S. H. Lui. On linear monotone iteration and Schwarz methods for nonlinear elliptic PDEs. Numerische Mathematik, 93(1):109–129, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lonsing:2004:SBP

[LV04a] Marco Lonsing and Rüdiger Verfürth. On the stability of BDMS and PEERS elements. Numerische Mathematik, 99(1):131–140, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lonsing:2004:PEE

[LV04b] Marco Lonsing and Rüdiger Verfürth. A posteriori error estimators for mixed finite element methods in linear elasticity. Numerische Mathematik, 97(4):757–778, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lai:2000:SDD

[LW00] Ming-Jun Lai and Paul Wenston. On Schwarz’s domain decomposition methods for elliptic boundary value problems. Numerische Mathematik, 84(3):475–495, January 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0084003/00840475.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0084003/00840475.pdf.

Lee:2009:RMM

[LWC09] Young-Ju Lee, Jinbiao Wu, and Jinru Chen. Robust multigrid method for the planar linear elasticity problems. Numerische Mathematik, 111(4):473–496, February 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=111&issue=4&spage=
473.

Liu:2001:CRD

[LWW01] Hailiang Liu, Jinghua Wang, and Gerald Warnecke. Convergence rates to discrete shocks for nonconvex conservation laws. Numerische Mathematik, 88(3):513–541, May 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1088003/10880513.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1088003/
10880513.pdf.

Li:2009:CSN

[LX09] Zhiping Li and Xianmin Xu. Convergence and stability of a numerical method for micromagnetics. Numerische Mathematik, 112 (2):245–265, April 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=112&issue=2&spage=245.

Liu:2001:QNP

[LY01] Wenbin Liu and Ningning Yan. Quasi-norm a priori and a posteriori error estimates for the nonconforming approximation of p-Laplacian. Numerische Mathematik, 89(2):341–378, August 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089002/10890341.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089002/10890341.pdf.

Liu:2003:PEE

[LY03] Wenbin Liu and Ningning Yan. A posteriori error estimates for optimal control problems governed by parabolic equations. Numerische Mathematik, 93(3):497–521, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:2003:NSH

[LYC03] Xiang-Gui Li, Wei Yan, and C. K. Chan. Numerical schemes for Hamilton–Jacobi equations on unstructured meshes. Numerische Mathematik, 94(2):315–331, April 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyly:2000:CBS

[Lyl00] Mikko Lyly. On the connection between some linear triangular Reissner–Mindlin plate bending elements. Numerische Mathematik, 85(1):77–107, March 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085001/00850077.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085001/00850077.pdf.

Lin:2001:SFE

[LYZ01] Qun Lin, Ningning Yan, and Aihui Zhou. A sparse finite element method with high accuracy Part I. Numerische Mathematik, 88(4): 731–742, June 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088004/10880731.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088004/10880731.pdf.

Li:2009:RCG

[LZ09] Ren-Cang Li and Wei Zhang. The rate of convergence of GMRES on a tridiagonal Toeplitz linear system. Numerische Mathematik, 112(2):267–293, April 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=112&issue=2&spage=267.

Mahmood:2009:SNC

[Mah09] Mohammed Shuker Mahmood. Solution of nonlinear convectiondiffusion problems by a conservative Galerkin-characteristics method. Numerische Mathematik, 112(4):601–636, June 2009.

CODEN NUMMA7. ISSN ???? URL http://www.springerlink.
com/openurl.asp?genre=article&issn=0029-599X&volume=112&
issue=4&spage=601.

Marusic:2001:FSO

[Mar01] Miljenko Marušić. A fourth/second order accurate collocation method for singularly perturbed two-point boundary value problems using tension splines. Numerische Mathematik, 88(1):135– 158, March 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088001/10880135.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088001/10880135.pdf.

Marcinkowski:2002:MEM

[Mar02] Leszek Marcinkowski. A mortar element method for some discretizations of a plate problem. Numerische Mathematik, 93(2): 361–386, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maset:2000:SRK

[Mas00] S. Maset. Stability of Runge–Kutta methods for linear delay differential equations. Numerische Mathematik, 87(2):355–371, December 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087002/00870355.htm;

 http://link.springer-ny.com/link/service/journals/00211/
papers/0087002/00870355.pdf.

Maset:2002:IRK

[Mas02] S. Maset. Instability of Runge–Kutta methods when applied to linear systems of delay differential equations. Numerische Mathematik, 90(3):555–562, January 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
2090003/20900555.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/2090003/20900555.pdf.

Matejas:2000:QCS

[Mat00] J. Matejaš. Quadratic convergence of scaled matrices in Jacobi method. Numerische Mathematik, 87(1):171–199, November 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087001/00870171.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087001/00870171.pdf.

Matos:2001:RCP

[Mat01] Ana C. Matos. Recursive computation of Padé-Legendre approximants and some acceleration properties. Numerische Mathematik, 89(3):535–560, September 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1089003/10890535.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089003/10890535.pdf.

Maury:2006:TSS

[Mau06] B. Maury. A time-stepping scheme for inelastic collisions. Numerische Mathematik, 102(4):649–679, February 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=4&spage=649.

Maxim:2007:AEP

[Max07] Andrei Maxim. Asymptotic exactness of an a posteriori error estimator based on the equilibrated residual method. Numerische Mathematik, 106(2):225–253, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mazure:2008:DEQ

[Maz08a] Marie-Laurence Mazure. On dimension elevation in Quasi Extended Chebyshev spaces. Numerische Mathematik, 109(3):459– 475, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mazure:2008:WSD

[Maz08b] Marie-Laurence Mazure. Which spaces for design? Numerische Mathematik, 110(3):357–392, September 2008. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=110&issue=
3&spage=357.

Matache:2000:GPF

[MBS00] A. M. Matache, I. Babuška, and C. Schwab. Generalized p-FEM in homogenization. Numerische Mathematik, 86(2):319– 375, August 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086002/00860319.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086002/00860319.pdf.

Milovanovic:2004:UCG

[MC04] Gradimir V. Milovanović and Aleksandar S. Cvetković. Uniqueness and computation of Gaussian interval quadrature formula for Jacobi weight function. Numerische Mathematik, 99(1):141–162, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Milovanovic:2006:GRG

[MC06] Gradimir V. Milovanović and Aleksandar S. Cvetković. Gauss– Radau and Gauss–Lobatto interval quadrature rules for Jacobi weight function. Numerische Mathematik, 102(3):523–542, January 2006. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=102&issue=
3&spage=523.

Ming:2002:OEB

[McS02] Pingbing Ming and Zhong ci Shi. Optimal L^{2} error bounds for MITC3 type element. Numerische Mathematik, 91(1):77–91, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mao:2009:HAA

[McS09a] Shipeng Mao and Zhong ci Shi. High accuracy analysis of two nonconforming plate elements. Numerische Mathematik, 111(3): 407–443, January 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=3&spage=407.

Milovanovi:2009:QFM

[MCS09b] Gradimir V. Milovanovi, Aleksandar S. Cvetkovi, and Marija P. Stani. Quadrature formulae with multiple nodes and a maximal trigonometric degree of exactness. Numerische Mathematik, 112 (3):425–448, May 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=112&issue=3&spage=425.

Medjo:2001:NFS

[Med01] T. Tachim Medjo. New formulations of a Stokes type problem related to the primitive equations of the atmosphere and applications. Numerische Mathematik, 87(3):503–522, January 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1087003/10870503.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1087003/10870503.pdf.

Mastroianni:2003:PCS

[MFR03] G. Mastroianni, C. Frammartino, and A. Rathsfeld. On polynomial collocation for second kind integral equations with fixed singularities of Mellin type. Numerische Mathematik, 94(2):333– 365, April 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Micu:2002:UBC

[Mic02] Sorin Micu. Uniform boundary controllability of a semi-discrete 1-D wave equation. Numerische Mathematik, 91(4):723–768, June 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Markovsky:2004:CLS

[MKV04] I. Markovsky, A. Kukush, and S. Van Huffel. Consistent least squares fitting of ellipsoids. Numerische Mathematik, 98(1):177– 194, July 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

McLean:2007:SOA

[MM07] William McLean and Kassem Mustapha. A second-order accurate numerical method for a fractional wave equation. Numerische Mathematik, 105(3):481–510, January 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Makazaga:2009:NCS

[MM09] J. Makazaga and A. Murua. A new class of symplectic integration schemes based on generating functions. Numerische Mathematik, 113(4):631–642, October 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=4&spage=
631.

Makridakis:2006:PEA

[MN06] Charalambos Makridakis and Ricardo H. Nochetto. A posteriori error analysis for higher order dissipative methods for evolution problems. Numerische Mathematik, 104(4):489–514, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morales:2008:AFS

[MNS08] José Luis Morales, Jorge Nocedal, and Mikhail Smelyanskiy. An algorithm for the fast solution of symmetric linear complementarity problems. Numerische Mathematik, 111(2):251–266, December 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=2&spage=251.

Moore:2001:IEB

[Moo01] Peter K. Moore. Interpolation error-based a posteriori error estimation for two-point boundary value problems and parabolic equations in one space dimension. Numerische Mathematik, 90(1): 149–177, November 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1090001/10900149.

htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1090001/10900149.pdf.

Moore:2003:ALP

[Moo03] Peter K. Moore. Applications of Lobatto polynomials to an adaptive finite element method: A posteriori error estimates for HP-Adaptivity and Grid-to-Grid Interpolation. Numerische Mathematik, 94(2):367–401, April 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moszynski:2001:AFF

[Mos01] Krzysztof Moszyński. Approximation with frequency filter for backward parabolic equations. Numerische Mathematik, 88(1): 159–183, March 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1088001/10880159.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1088001/10880159.pdf.

Marban:2002:NRH

[MP02] J. M. Marbán and C. Palencia. On the numerical recovery of a holomorphic mapping from a finite set of approximate values. Numerische Mathematik, 91(1):57–75, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Markowich:2005:FCE

[MP05] Peter A. Markowich and Lorenzo Pareschi. Fast conservative and entropic numerical methods for the Boson Boltzmann equation. Numerische Mathematik, 99(3):509–532, January 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maass:2001:ADT

[MPRS01] Peter Maaß, Sergei V. Pereverzev, Ronny Ramlau, and Sergei G. Solodky. An adaptive discretization for Tikhonov–Phillips regularization with a posteriori parameter selection. Numerische Mathematik, 87(3):485–502, January 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087003/10870485.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087003/10870485.pdf.

Mikula:2001:SIF

[MR01] Karol Mikula and Narisoa Ramarosy. Semi-implicit finite volume scheme for solving nonlinear diffusion equations in image processing. Numerische Mathematik, 89(3):561–590, September 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089003/10890561.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089003/10890561.pdf.

Moore:2003:BEA

[MR03] Brian Moore and Sebastian Reich. Backward error analysis for multi-symplectic integration methods. Numerische Mathematik, 95(4):625–652, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monegato:2000:GPA

[MS00a] G. Monegato and L. Scuderi. Global polynomial approximation for Symm’s equation on polygons. Numerische Mathematik, 86 (4):655–683, October 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0086004/00860655.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0086004/00860655.pdf.

Mulansky:2000:CBS

[MS00b]  Bernd Mulansky and Jochen W. Schmidt.  Composition

based staircase algorithm and constrained interpolation with boundary conditions. Numerische Mathematik, 85(3):387–408, May 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085003/00850387.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085003/00850387.pdf.

Maymeskul:2002:APL

[MSS02] V. V. Maymeskul, E. B. Saff, and N. S. Stylianopoulos. L^{2}- approximations of power and logarithmic functions with applications to numerical conformal mapping. Numerische Mathematik, 91(3):503–542, May 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=91&issue=3&spage=503.

Maday:2006:MTD

[MST06a] Yvon Maday, Julien Salomon, and Gabriel Turinici. Monotonic time-discretized schemes in quantum control. Numerische Mathematik, 103(2):323–338, April 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=103&issue=2&spage=323.

McLean:2006:TDL

[MST06b] William McLean, Ian H. Sloan, and Vidar Thomée. Time discretization via Laplace transformation of an integro-differential equation of parabolic type. Numerische Mathematik, 102(3): 497–522, January 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=3&spage=497.

Moon:2003:CRA

[MSTZ03a] Kyoung-Sook Moon, Anders Szepessy, Raúl Tempone, and Georgios E. Zouraris. Convergence rates for adaptive approximation of ordinary differential equations. Numerische Mathematik, 96(1): 99–129, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moon:2003:VPA

[MSTZ03b] Kyoung-Sook Moon, Anders Szepessy, Raúl Tempone, and Georgios E. Zouraris. A variational principle for adaptive approximation of ordinary differential equations. Numerische Mathematik, 96(1):131–152, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Matache:2006:LCS

[MSW06] A.-M. Matache, C. Schwab, and T. P. Wihler. Linear complexity solution of parabolic integro-differential equations. Numerische Mathematik, 104(1):69–102, July 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mandel:2001:CDP

[MT01] Jan Mandel and Radek Tezaur. On the convergence of a dualprimal substructuring method. Numerische Mathematik, 88(3): 543–558, May 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088003/10880543.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088003/10880543.pdf.

Maday:2003:EBQ

[MT03] Y. Maday and G. Turinici. Error bars and quadratically convergent methods for the numerical simulation of the Hartree–Fock equations. Numerische Mathematik, 94(4):739–770, June 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=94&issue=4&spage=
739.

Matthies:2005:ISS

[MT05] Gunar Matthies and Lutz Tobiska. Inf-sup stable non-conforming finite elements of arbitrary order on triangles. Numerische Mathematik, 102(2):293–309, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=2&spage=293.

Muller:2001:CCH

[Mul01] Keith E. Muller. Computing the confluent hypergeometric function, M(a, b, x). Numerische Mathematik, 90(1):179–196, November 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090001/10900179.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090001/10900179.pdf.

Muller:2003:CPR

[Mül03] Stefan Müller. Critique of the paper ’the relaxation nonquasiconvex variational integrals’. Numerische Mathematik, 93 (3):523–525, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Merlet:2007:EEF

[MV07] Benoît Merlet and Julien Vovelle. Error estimate for finite volume scheme. Numerische Mathematik, 106(1):129–155, March 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mardal:2004:UPT

[MW04] Kent-Andre Mardal and Ragnar Winther. Uniform preconditioners for the time dependent Stokes problem. Numerische Mathematik, 98(2):305–327, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [MW06].

Mardal:2006:EUP

[MW06] Kent-Andre Mardal and Ragnar Winther. Erratum: Uniform preconditioners for the time dependent Stokes problem, Num. Math. 98(2) 305–327 (2004). Numerische Mathematik, 103(1):171–172, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=1&spage=171. See [MW04].

Wu:2005:SCT

[mWS05] ming Wu and Weiwei Sun. The superconvergence of the composite trapezoidal rule for Hadamard finite part integrals. Numerische Mathematik, 102(2):343–363, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=2&spage=343.

Ming:2006:MEF

[MX06] Wang Ming and Jinchao Xu. The Morley element for fourth order elliptic equations in any dimensions. Numerische Mathematik, 103(1):155–169, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=103&issue=1&spage=155.

Malajovich:2001:TGI

[MZ01] Gregorio Malajovich and Jorge P. Zubelli. Tangent Graeffe iteration. Numerische Mathematik, 89(4):749–782, October 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089004/10890749.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089004/10890749.pdf.

Nabben:2003:CBM

[Nab03] Reinhard Nabben. Comparisons between multiplicative and additive Schwarz iterations in domain decomposition methods.

Numerische Mathematik, 95(1):145–162, July 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nordbotten:2007:MCV

[NAE07] J. M. Nordbotten, I. Aavatsmark, and G. T. Eigestad. Monotonicity of control volume methods. Numerische Mathematik, 106 (2):255–288, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Noutsos:2006:BBT

[NCV06] D. Noutsos, S. Serra Capizzano, and P. Vassalos. Block band Toeplitz preconditioners derived from generating function approximations: analysis and applications. Numerische Mathematik, 104 (3):339–376, September 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Naiman:2000:NCG

[NE00] Aaron E. Naiman and Shlomo Engelberg. A note on conjugate gradient convergence — Part II. Numerische Mathematik, 85(4): 665–683, June 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085004/00850665.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085004/00850665.pdf.

Negri:2003:FEA

[Neg03] Matteo Negri. A finite element approximation of the Griffith’s model in fracture mechanics. Numerische Mathematik, 95(4):653– 687, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Negulescu:2008:NAM

[Neg08] Claudia Negulescu. Numerical analysis of a multiscale finite element scheme for the resolution of the stationary Schrödinger equation. Numerische Mathematik, 108(4):625–652, February 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neubauer:2000:LIN

[Neu00] Andreas Neubauer. On Landweber iteration for nonlinear ill-posed problems in Hilbert scales. Numerische Mathematik, 85(2):309– 328, April 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085002/00850309.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085002/00850309.pdf.

Nievergelt:2002:FAF

[Nie02] Yves Nievergelt. A finite algorithm to fit geometrically all midrange lines, circles, planes, spheres, hyperplanes, and hyperspheres. Numerische Mathematik, 91(2):257–303, April 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nicholls:2006:EAE

[NN06] David P. Nicholls and Nilima Nigam. Error analysis of an enhanced DtN–FE method for exterior scattering problems. Numerische Mathematik, 105(2):267–298, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Noel:2004:PEE

[Noë04] Virginie Bonnaillie Noël. A posteriori error estimator for the eigenvalue problem associated to the Schrödinger operator with magnetic field. Numerische Mathematik, 99(2):325–348, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Notaris:2006:ENG

[Not06] Sotirios E. Notaris. The error norm of Gauss–Kronrod quadrature formulae for weight functions of Bernstein–Szegő type. Numerische Mathematik, 103(1):99–127, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=103&issue=1&spage=99.

Novruzi:2001:EEB

[Nov01] Arian Novruzi. C^{1} error estimation on the boundary for an exterior Neumann problem in R^{3}. Numerische Mathematik, 88(1):185– 201, March 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088001/10880185.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088001/10880185.pdf.

Nochetto:2004:ORP

[NP04] Ricardo H. Nochetto and Jae-Hong Pyo. Optimal relaxation parameter for the Uzawa Method. Numerische Mathematik, 98(4): 695–702, October 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nicholls:2003:ACD

[NR03] David P. Nicholls and Fernando Reitich. Analytic continuation of Dirichlet–Neumann operators. Numerische Mathematik, 94 (1):107–146, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nochetto:2006:PPE

[NSSV06] Ricardo H. Nochetto, Alfred Schmidt, Kunibert G. Siebert, and Andreas Veeser. Pointwise a posteriori error estimates for monotone semi-linear equations. Numerische Mathematik, 104(4):515– 538, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nochetto:2003:PPE

[NSV03] Ricardo H. Nochetto, Kunibert G. Siebert, and Andreas Veeser. Pointwise a posteriori error control for elliptic obstacle problems. Numerische Mathematik, 95(1):163–195, July 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Novo:2001:EMU

[NTW01] Julia Novo, Edriss S. Titi, and Shannon Wynne. Efficient methods using high accuracy approximate inertial manifolds. Numerische Mathematik, 87(3):523–554, January 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087003/10870523.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087003/10870523.pdf.

Natterer:2005:MSI

[NW05] Frank Natterer and Frank Wübbeling. Marching schemes for inverse acoustic scattering problems. Numerische Mathematik, 100(4):697–710, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=100&issue=4&spage=697.

Oberman:2004:CMD

[Obe04] Adam M. Oberman. A convergent monotone difference scheme for motion of level sets by mean curvature. Numerische Mathematik, 99(2):365–379, December 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ohlberger:2001:PEE

[Ohl01] Mario Ohlberger. A posteriori error estimate for finite volume approximations to singularly perturbed nonlinear convectiondiffusion equations. Numerische Mathematik, 87(4):737–761, February 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1087004/10870737.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1087004/10870737.pdf.

Olver:2006:QMH

[Olv06] Sheehan Olver. On the quadrature of multivariate highly oscillatory integrals over non-polytope domains. Numerische Mathematik, 103(4):643–665, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Owren:2001:IMB

[OM01] Brynjulf Owren and Arne Marthinsen. Integration methods based on canonical coordinates of the second kind. Numerische Mathematik, 87(4):763–790, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087004/10870763.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087004/10870763.pdf.

Olshanskii:2006:UPP

[OPR06] Maxim A. Olshanskii, Jörg Peters, and Arnold Reusken. Uniform preconditioners for a parameter dependent saddle point problem with application to generalized Stokes interface equations. Numerische Mathematik, 105(1):159–191, November 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oishi:2002:FVS

[OR02] Shin’Ichi Oishi and Siegfried M. Rump. Fast verification of solutions of matrix equations. Numerische Mathematik, 90(4): 755–773, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/2090004/20900755.htm;
 http://link.springer.de/link/service/journals/00211/
papers/2090004/20900755.pdf.

Olshanskii:2006:ASI

[OR06] Maxim A. Olshanskii and Arnold Reusken. Analysis of a Stokes interface problem. Numerische Mathematik, 103(1): 129–149, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=103&issue=1&spage=129.

OLeary:2007:SGQ

[OST07] Dianne P. OLeary, Zdenk Strakos, and Petr Tichý. On sensitivity of Gauss–Christoffel quadrature. Numerische Mathematik, 107(1): 147–174, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oswald:2008:OMP

[Osw08] P. Oswald. Optimality of multilevel preconditioning for nonconforming P_{1} finite elements. Numerische Mathematik, 111(2):267– 291, December 2008. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=111&issue=2&spage=267.

Ovall:2006:AEF

[Ova06] Jeffrey S. Ovall. Asymptotically exact functional error estimators based on superconvergent gradient recovery. Numerische Mathematik, 102(3):543–558, January 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=102&issue=3&spage=543.

Oliver:2004:AMC

[OWW04] Marcel Oliver, Matthew West, and Claudia Wulff. Approximate momentum conservation for spatial semidiscretizations of semilinear wave equations. Numerische Mathematik, 97(3):493–535, May 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pampel:2001:NAC

[Pam01] Thorsten Pampel. Numerical approximation of connecting orbits with asymptotic rate. Numerische Mathematik, 90(2):309–348, December 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090002/10900309.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090002/10900309.pdf.

Pasquini:2000:ACZ

[Pas00] L. Pasquini. Accurate computation of the zeros of the generalized Bessel polynomials. Numerische Mathematik, 86(3):507–538, September 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086003/00860507.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086003/00860507.pdf.

Patrick:2009:EAV

[PC09] George W. Patrick and Charles Cuell. Error analysis of variational integrators of unconstrained Lagrangian systems. Numerische Mathematik, 113(2):243–264, August 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=2&spage=
243.

Pares:2002:CBM

[PCM02] Carlos Parés, Manuel Castro, and Jorge Macías. On the convergence of the Bermúdez-Moreno algorithm with constant parameters. Numerische Mathematik, 92(1):113–128, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pena:2002:OSB

[Peñ02] J. M. Peña. On the optimal stability of bases of univariate functions. Numerische Mathematik, 91(2):305–318, April 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pena:2003:AGC

[Peñ03] J. M. Peña. On an alternative to Gerschgorin circles and ovals of Cassini. Numerische Mathematik, 95(2):337–345, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pena:2006:NOS

[Peñ06] J. M. Peña. A note on the optimal stability of bases of univariate functions. Numerische Mathematik, 103(1):151–154, March 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=103&issue=1&spage=
151.

Peradze:2005:NAN

[Per05] Jemal Peradze. A numerical algorithm for the nonlinear Kirchhoff string equation. Numerische Mathematik, 102(2): 311–342, December 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=2&spage=311.

Peters:2002:PN

[Pet02] Martin Peters. Publisher’s note. Numerische Mathematik, 91(2): 195–196, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Petras:2003:SCG

[Pet03] Knut Petras. Smolyak cubature of given polynomial degree with few nodes for increasing dimension. Numerische Mathematik, 93 (4):729–753, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Prieto:2007:CCS

[PGL07] U. Fidalgo Prieto, J. R. Illán González, and G. López Lagomasino. Convergence and computation of simultaneous rational quadrature formulas. Numerische Mathematik, 106(1):99–128, March 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pennanen:2005:ECD

[PK05] Teemu Pennanen and Matti Koivu. Epi-convergent discretizations of stochastic programs via integration quadratures. Numerische Mathematik, 100(1):141–163, March 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pinto:2008:CAS

[PM08] Martin Campos Pinto and Michel Mehrenberger. Convergence of an adaptive semi-Lagrangian scheme for the Vlasov–Poisson system. Numerische Mathematik, 108(3):407–444, January 2008. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pares:2001:DMA

[PMC01] Carlos Parés, Jorge Macías, and Manuel Castro. Duality methods with an automatic choice of parameters: Application to shallow water equations in conservative form. Numerische Mathematik, 89(1):161–189, July 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1089001/10890161.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1089001/10890161.pdf.

Potthast:2004:NNI

[Pot04] Roland Potthast. A new non-iterative singular sources method for the reconstruction of piecewise constant media. Numerische Mathematik, 98(4):703–730, October 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peherstorfer:2003:SPG

[PP03] Franz Peherstorfer and Knut Petras. Stieltjes polynomials and Gauss–Kronrod quadrature for Jacobi weight functions. Numerische Mathematik, 95(4):689–706, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plonka:2001:NFT

[PR01] Gerlind Plonka and Amos Ron. A new factorization technique of the matrix mask of univariate refinable functions. Numerische Mathematik, 87(3):555–595, January 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1087003/10870555.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1087003/10870555.pdf.

Potzsche:2009:CNI

[PR09] Christian Pötzsche and Martin Rasmussen. Computation of nonautonomous invariant and inertial manifolds. Numerische Mathematik, 112(3):449–483, May 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=112&issue=3&spage=

Prohl:2000:FOP

[Pro00] Andreas Prohl. A first order projection-based time-splitting scheme for computing chemically reacting flows. Numerische Mathematik, 84(4):649–677, February 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0084004/00840649.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0084004/00840649.pdf.

Pitkaranta:2000:ULE

[PS00] Juhani Pitkäranta and Manil Suri. Upper and lower error bounds for plate-bending finite elements. Numerische Mathematik, 84(4): 611–648, February 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0084004/00840611.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0084004/00840611.pdf.

Paige:2002:BLS

[PS02a] Christopher C. Paige and Zdenek Strakoš. Bounds for the least squares distance using scaled total least squares. Numerische Mathematik, 91(1):93–115, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paige:2002:STL

[PS02b] Christopher C. Paige and Zdenek Strakoš. Scaled total least squares fundamentals. Numerische Mathematik, 91(1):117–146, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pechstein:2009:AFM

[PS09a] Clemens Pechstein and Robert Scheichl. Analysis of FETI methods for multiscale PDEs. Numerische Mathematik, 114(2):293– 333, December 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=114&issue=2&spage=293.

Prohl:2009:CDN

[PS09b] Andreas Prohl and Markus Schmuck. Convergent discretizations for the Nernst–Planck–Poisson system. Numerische Mathematik, 111(4):591–630, February 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=111&issue=4&spage=
591.

Potts:2004:FCR

[PSN04] Daniel Potts, Gabriele Steidl, and Arthur Nieslony. Fast convolution with radial kernels at nonequispaced knots. Numerische Mathematik, 98(2):329–351, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Popov:2006:OSS

[PT06] Bojan Popov and Ognian Trifonov. One-sided stability and convergence of the Nessyahu–Tadmor scheme. Numerische Mathematik, 104(4):539–559, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pareschi:2003:SMN

[PTV03] L. Pareschi, G. Toscani, and C. Villani. Spectral methods for the non cut-off Boltzmann equation and numerical grazing collision limit. Numerische Mathematik, 93(3):527–548, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Putinar:2002:DPA

[Put02] Mihai Putinar. On a diagonal Padé approximation in two complex variables. Numerische Mathematik, 93(1):131–152, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plum:2003:OPE

[PW03] Michael Plum and Christian Wieners. Optimal a priori estimates for interface problems. Numerische Mathematik, 95(4):735–759, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Perthame:2005:TOD

[PW05a] Benoît Perthame and Michael Westdickenberg. Total oscillation diminishing property for scalar conservation laws. Numerische Mathematik, 100(2):331–349, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plaskota:2005:AAE

[PW05b] Leszek Plaskota and Grzegorz W. Wasilkowski. Adaption allows efficient integration of functions with unknown singularities. Numerische Mathematik, 102(1):123–144, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=1&spage=123.

Pop:2002:NAD

[PY02] Iuliu Sorin Pop and Wen-An Yong. A numerical approach to degenerate parabolic equations. Numerische Mathematik, 92(2):357– 381, August 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pytlak:2002:CGC

[Pyt02] R. Pytlak. Comments on “Global convergence of the method of shortest residuals” by Y. Dai and Y. Yuan. Numerische Mathematik, 91(2):319–321, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [hDxY99].

Qi:2003:FTD

[QQ03] Houduo Qi and Liqun Qi. Finite termination of a dual Newton method for convex best C^{1} interpolation and smoothing. Numerische Mathematik, 96(2):317–337, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rodriguez-Aros:2007:NAF

[RAVS07] Á. Rodríguez-Arós, J. M. Viaño, and M. Sofonea. Numerical analysis of a frictional contact problem for viscoelastic materials with long-term memory. Numerische Mathematik, 108(2):327–358, December 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reusken:2002:CAM

[Reu02] Arnold Reusken. Convergence analysis of a multigrid method for convection-diffusion equations. Numerische Mathematik, 91(2): 323–349, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rebollo:2004:SST

[RG04] T. Chacón Rebollo and D. Rodríguez Gómez. A stabilized space– time discretization for the primitive equations in oceanography.

Numerische Mathematik, 98(3):427–475, September 2004. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rieder:2001:CRI

[Rie01] Andreas Rieder. On convergence rates of inexact Newton regularizations. Numerische Mathematik, 88(2):347–365, April 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088002/10880347.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088002/10880347.pdf.

Radu:2008:EEM

[RPK08] Florin A. Radu, Iuliu Sorin Pop, and Peter Knabner. Error estimates for a mixed finite element discretization of some degenerate parabolic equations. Numerische Mathematik, 109(2):285–311, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robbe:2002:RBP

[RS02] M. Robbé and M. Sadkane. Riccati-based preconditioner for computing invariant subspaces of large matrices. Numerische Mathematik, 92(1):129–159, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rieder:2004:AIA

[RS04] Andreas Rieder and Thomas Schuster. The approximate inverse in action III: 3D-Doppler tomography. Numerische Mathematik, 97(2):353–378, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rieder:2007:OFD

[RS07] Andreas Rieder and Arne Schneck. Optimality of the fully discrete filtered backprojection algorithm for tomographic inversion. Numerische Mathematik, 108(1):151–175, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rech:2006:TSC

[RSS06] M. Rech, S. Sauter, and A. Smolianski. Two-scale composite finite element method for Dirichlet problems on complicated domains. Numerische Mathematik, 102(4):681–708, February 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=102&issue=4&spage=681.

Riviere:2003:DGF

[RSWW03] Béatrice Rivière, Simon Shaw, Mary F. Wheeler, and J. R. Whiteman. Discontinuous Galerkin finite element methods for linear elasticity and quasistatic linear viscoelasticity. Numerische Mathematik, 95(2):347–376, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rui:2002:SOC

[RT02] Hongxing Rui and Masahisa Tabata. A second order characteristic finite element scheme for convection-diffusion problems. Numerische Mathematik, 92(1):161–177, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ramlau:2006:TBP

[RT06] Ronny Ramlau and Gerd Teschke. A Tikhonov-based projection iteration for nonlinear ill-posed problems with sparsity constraints. Numerische Mathematik, 104(2):177–203, August 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rodriguez:2009:FEC

[RVH09] Ana Alonso Rodríguez, Alberto Valli, and Rafael Vázquez Hernández. A formulation of the eddy current problem in the presence of electric ports. Numerische Mathematik, 113(4):643– 672, October 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=113&issue=4&spage=643.

Rahman:2005:ASM

[RXH05] Talal Rahman, Xuejun Xu, and Ronald Hoppe. Additive Schwarz methods for the Crouzeix–Raviart mortar finite element for elliptic problems with discontinuous coefficients. Numerische Mathematik, 101(3):551–572, September 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=3&spage=551.

Saranen:2007:BES

[SA07] J. Saranen and J. Anttila. The boundary element spline collocation for nonuniform meshes on the torus. Numerische Mathematik, 106(2):289–302, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sangalli:2001:RPE

[San01] Giancarlo Sangalli. A robust a posteriori estimator for the Residual-free Bubbles method applied to advection-diffusion problems. Numerische Mathematik, 89(2):379–399, August 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1089002/10890379.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1089002/
10890379.pdf.

Sangalli:2004:AAD

[San04] Giancarlo Sangalli. Analysis of the advection-diffusion operator using fractional order norms. Numerische Mathematik, 97(4):779– 796, June 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smoktunowicz:2006:NEA

[SBL06] Alicja Smoktunowicz, Jesse L. Barlow, and Julien Langou. A note on the error analysis of classical Gram–Schmidt. Numerische Mathematik, 105(2):299–313, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Serra-Capizzano:2007:SRB

[SC07] Stefano Serra-Capizzano. Simplification of a result on banded Toeplitz matrices and BVM methods. Numerische Mathematik, 107(1):175–179, July 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schotzau:2004:MFE

[Sch04] Dominik Schötzau. Mixed finite element methods for stationary incompressible magnetohydrodynamics. Numerische Mathematik, 96(4):771–800, February 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:2009:APC

[Sch09] Reinhold Schneider. Analysis of the projected coupled cluster method in electronic structure calculation. Numerische Mathematik, 113(3):433–471, September 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=3&spage=
433.

Strau:2008:EAR

[SHS08] Frank Strau, Vincent Heuveline, and Ben Schweizer. Existence and approximation results for shape optimization problems in rotor dynamics. Numerische Mathematik, 109(2):313–332, April 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sidi:2004:EME

[Sid04] Avram Sidi. Euler–Maclaurin expansions for integrals with endpoint singularities: a new perspective. Numerische Mathematik, 98(2):371–387, August 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sidi:2005:AAV

[Sid05] Avram Sidi. Analysis of Atkinson’s variable transformation for numerical integration over smooth surfaces in R^{3}. Numerische Mathematik, 100(3):519–536, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Simonnet:2000:VCT

[Sim00] E. Simonnet. A variational crime on the tangent bundle of the sphere. Numerische Mathematik, 85(3):409–431, May 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085003/00850409.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085003/00850409.pdf.

Sidi:2003:OPS

[SK03] Avram Sidi and Yuliya Kanevsky. Orthogonal polynomials and semi-iterative methods for the Drazin-inverse solution of singular linear systems. Numerische Mathematik, 93(3):563–581, January 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneid:2004:PEE

[SKR04] Eckhard Schneid, Peter Knabner, and Florin Radu. A priori error estimates for a mixed finite element discretization of the Richards’ equation. Numerische Mathematik, 98(2):353–370, August 2004.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Santos-Leon:2002:CRH

[SL02] J. C. Santos-León. Computation of the Riesz–Herglotz transform and its application to quadrature formulas over the unit circle. Numerische Mathematik, 93(1):153–175, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Song:2000:SNS

[Son00] Yongzhong Song. Semiconvergence of nonnegative splittings for singular matrices. Numerische Mathematik, 85(1):109–127, March 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085001/00850109.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085001/00850109.pdf.

Song:2002:CTS

[Son02] Yongzhong Song. Comparison theorems for splittings of matrices. Numerische Mathematik, 92(3):563–591, September 2002. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder-Pander:2000:GMA

[SPSF00] Friederike Schröder-Pander, Thomas Sonar, and Oliver Friedrich. Generalized multiresolution analysis on unstructured grids. Numerische Mathematik, 86(4):685–715, October 2000. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0086004/00860685.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0086004/
00860685.pdf.

Scherer:2000:EQE

[SS00] Rudolf Scherer and Thomas Schira. Estimating quadrature errors for analytic functions using kernel representations and biorthogonal systems. Numerische Mathematik, 84(3):497–518, January 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0084003/00840497.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0084003/00840497.pdf.

Simoncini:2005:ENO

[SS05] Valeria Simoncini and Daniel B. Szyld. The effect of non-optimal bases on the convergence of Krylov subspace methods. Numerische Mathematik, 100(4):711–733, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=100&issue=4&spage=711.

Simoncini:2008:NCN

[SS08] Valeria Simoncini and Daniel B. Szyld. New conditions for nonstagnation of minimal residual methods. Numerische Mathematik, 109(3):477–487, May 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schwab:2003:SFE

[ST03] Christoph Schwab and Radu-Alexandru Todor. Sparse finite elements for elliptic problems with stochastic loading. Numerische Mathematik, 95(4):707–734, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Suzuki:2005:CAT

[ST05] Takashi Suzuki and Takuya Tsuchiya. Convergence analysis of trial free boundary methods for the two-dimensional filtration problem. Numerische Mathematik, 100(3):537–564, May 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Segura:2008:NSS

[ST08] Javier Segura and Nico M. Temme. Numerically satisfactory solutions of Kummer recurrence relations. Numerische Mathematik, 111(1):109–119, November 2008. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=111&issue=1&spage=
109.

Steinbach:2000:HBE

[Ste00] Olaf Steinbach. On a hybrid boundary element method. Numerische Mathematik, 84(4):679–695, February 2000. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/0084004/00840679.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0084004/
00840679.pdf.

Steinbach:2001:SPF

[Ste01a] Olaf Steinbach. On the stability of the L_{2} projection in fractional Sobolev spaces. Numerische Mathematik, 88(2):367–379, April 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088002/10880367.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088002/10880367.pdf.

Stewart:2001:EGM

[Ste01b] G. W. Stewart. On the eigensystems of graded matrices. Numerische Mathematik, 90(2):349–370, December 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/
journals/00211/bibs/1090002/10900349.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/1090002/
10900349.pdf.

Steinbach:2002:GPS

[Ste02a] Olaf Steinbach. On a generalized L_{2} projection and some related stability estimates in Sobolev spaces. Numerische Mathematik, 90 (4):775–786, February 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.de/

link/service/journals/00211/bibs/2090004/20900775.htm;
 http://link.springer.de/link/service/journals/00211/
papers/2090004/20900775.pdf.

Stevenson:2002:NFE

[Ste02b] Rob Stevenson. Nonconforming finite elements and the cascadic multi-grid method. Numerische Mathematik, 91(2):351–387, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Steinbach:2003:RBE

[Ste03a] O. Steinbach. A robust boundary element method for nearly incompressible linear elasticity. Numerische Mathematik, 95(3): 553–562, September 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:2003:PMP

[Ste03b] G. W. Stewart. On the powers of a matrix with perturbations. Numerische Mathematik, 96(2):363–376, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stevenson:2005:USP

[Ste05] Rob P. Stevenson. The uniform saturation property for a singularly perturbed reaction-diffusion equation. Numerische Mathematik, 101(2):355–379, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=2&spage=355.

Stiemer:2005:AOE

[Sti05] Marcus Stiemer. On the approximation order of extremal point methods for hyperbolic minimal energy problems. Numerische Mathematik, 99(3):533–555, January 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stover:2001:CMS

[Stö01] Ronald Stöver. Collocation methods for solving linear differentialalgebraic boundary value problems. Numerische Mathematik, 88 (4):771–795, June 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1088004/10880771.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1088004/10880771.pdf.

Strobach:2008:UPA

[Str08] Peter Strobach. Updating the principal angle decomposition. Numerische Mathematik, 110(1):83–112, July 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strobach:2009:SRH

[Str09] Peter Strobach. Square-root Householder subspace tracking. Numerische Mathematik, 113(1):89–121, July 2009. CODEN

NUMMA7. ISSN ???? URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=113&issue=
1&spage=89.

Samelson:2007:FON

[STWW07] Roger Samelson, Roger Temam, Cheng Wang, and Shouhong Wang. A fourth-order numerical method for the planetary geostrophic equations with inviscid geostrophic balance. Numerische Mathematik, 107(4):669–705, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Steinbach:2009:BEM

[SU09] O. Steinbach and G. Unger. A boundary element method for the Dirichlet eigenvalue problem of the Laplace operator. Numerische Mathematik, 113(2):281–298, August 2009. CODEN NUMMA7.

ISSN ???? URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=113&issue=2&spage=
281.

Sun:2007:QIM

[Sun07] Yajuan Sun. Quadratic invariants and multi-symplecticity of partitioned Runge–Kutta methods for Hamiltonian PDEs. Numerische Mathematik, 106(4):691–715, June 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shaw:2000:NNE

[SW00] Simon Shaw and J. R. Whiteman. Negative norm error control for second-kind convolution Volterra equations. Numerische Mathematik, 85(2):329–341, April 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085002/00850329.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085002/00850329.pdf.

Shaw:2001:OLT

[SW01] Simon Shaw and John R. Whiteman. Optimal long-time L_{p}(0, T) stability and semidiscrete error estimates for the Volterra formulation of the linear quasistatic viscoelasticity problem. Numerische Mathematik, 88(4):743–770, June 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088004/10880743.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088004/10880743.pdf.

Schotzau:2003:ECM

[SW03] Dominik Schötzau and Thomas P. Wihler. Exponential convergence of mixed hp-DGFEM for Stokes flow in polygons. Numerische Mathematik, 96(2):339–361, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schotzau:2008:PEE

[SW08] Dominik Schötzau and Thomas P. Wihler. A posteriori error estimation for hp-version time-stepping methods for parabolic partial differential equations. Numerische Mathematik, 110(3):475–509, September 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=3&spage=475.

Schoberl:2003:STS

[SZ03] Joachim Schöberl and Walter Zulehner. On Schwarz-type smoothers for saddle point problems. Numerische Mathematik, 95 (2):377–399, August 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sokolowski:2005:MTD

[SŻ05] J. Sokołowski and A. Żochowski. Modelling of topological derivatives for contact problems. Numerische Mathematik, 102(1):145–179, November 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=1&spage=145.

Simon:2009:STS

[SZ09] René Simon and Walter Zulehner. On Schwarz-type smoothers for saddle point problems with applications to PDE-constrained optimization problems. Numerische Mathematik, 111(3):445–468, January 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=3&spage=445.

Szeftel:2006:ABC

[Sze06] Jérémie Szeftel. Absorbing boundary conditions for onedimensional nonlinear Schrödinger equations. Numerische Mathematik, 104(1):103–127, July 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tai:2003:RCS

[Tai03] Xue-Cheng Tai. Rate of convergence for some constraint decomposition methods for nonlinear variational inequalities. Numerische Mathematik, 93(4):755–786, February 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tran:2008:PPE

[TGSS08] T. Tran, Q. T. Le Gia, I. H. Sloan, and E. P. Stephan. Preconditioners for pseudodifferential equations on the sphere with radial basis functions. Numerische Mathematik, 111(1):141–163, November 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=1&spage=141.

Timmermann:2000:CMA

[Tim00] Gisela Timmermann. A cascadic multigrid algorithm for semilinear elliptic problems. Numerische Mathematik, 86(4):717–731, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086004/00860717.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086004/00860717.pdf.

Torrilhon:2007:EOE

[TJ07] Manuel Torrilhon and Rolf Jeltsch. Essentially optimal explicit Runge–Kutta methods with application to hyperbolic-parabolic equations. Numerische Mathematik, 106(2):303–334, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tanaka:2004:NSC

[TMO04] Ken’ichiro Tanaka, Sunao Murashige, and Shin’ichi Oishi. On necessary and sufficient conditions for numerical verification of double turning points. Numerische Mathematik, 97(3):537–554, May 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Toselli:2000:OSM

[Tos00] Andrea Toselli. Overlapping Schwarz methods for Maxwell’s equations in three dimensions. Numerische Mathematik, 86(4):733– 752, October 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0086004/00860733.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0086004/00860733.pdf.

Tran:2000:ASA

[TS00] Thanh Tran and Ernst P. Stephan. Additive Schwarz algorithms for the p-version of the Galerkin boundary element method. Numerische Mathematik, 85(3):433–468, May 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer-ny.com/link/service/

journals/00211/bibs/0085003/00850433.htm;   http://link.
springer-ny.com/link/service/journals/00211/papers/0085003/
00850433.pdf.

Toselli:2003:MHF

[TS03] Andrea Toselli and Christoph Schwab. Mixed HP-finite element approximations on geometric edge and boundary layer meshes in three dimensions. Numerische Mathematik, 94 (4):771–801, June 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=94&issue=4&spage=771.

Tsuchiya:2003:PFE

[Tsu03] Takuya Tsuchiya. Precise finite element error analysis by Yamamoto’s explicit inversion formula for tridiagonal matrices — an extension of Babuška–Osborn’s theorems. Numerische Mathematik, 94(3):541–572, May 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tabata:2005:EEF

[TT05] Masahisa Tabata and Daisuke Tagami. Error estimates of finite element methods for nonstationary thermal convection problems with temperature-dependent coefficients. Numerische Mathematik, 100(2):351–372, April 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=100&issue=2&spage=351.

Thomee:2000:SAM

[TW00] V. Thomée and L. B. Wahlbin. Stability and analyticity in maximum-norm for simplicial Lagrange finite element semidiscretizations of parabolic equations with Dirichlet boundary conditions. Numerische Mathematik, 87(2):373–389, December 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0087002/00870373.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0087002/00870373.pdf. See erratum [TW01].

Thomee:2001:ESA

[TW01] V. Thomée and L. B. Wahlbin. Erratum: Stability and analyticity in maximum-norm for simplicial Lagrange finite element semidiscretizations of parabolic equations with Dirichlet boundary conditions, Numer. Math. 87 (2), 2000, pp. 373–389. Numerische Mathematik, 89(1):191, July 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [TW00].

Tebou:2003:UEL

[TZ03] Louis Roder Tcheugoué Tébou and Enrique Zuazua. Uniform exponential long time decay for the space semi-discretization of a locally damped wave equation via an artificial numerical viscosity. Numerische Mathematik, 95(3):563–598, September 2003. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Uschmajew:2008:WPC

[Usc08] André Uschmajew. Well-posedness of convex maximization problems on Stiefel manifolds and orthogonal tensor product approximations. Numerische Mathematik, 110(2):309–331, August 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=110&issue=2&spage=309.

Vanek:2001:CAM

[VBM01] Petr Vaněk, Marian Brezina, and Jan Mandel. Convergence of algebraic multigrid based on smoothed aggregation. Numerische Mathematik, 88(3):559–579, May 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1088003/10880559.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1088003/10880559.pdf.

Varga:2001:ZPS

[VC01] Richard S. Varga and Amos J. Carpenter. Zeros of the partial sums of cos(z) and sin(z) II. Numerische Mathematik, 90(2):371–400, December 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090002/10900371.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090002/10900371.pdf.

VanDeun:2006:SRI

[VC06] Joris Van Deun and Ronald Cools. A stable recurrence for the incomplete gamma function with imaginary second argument. Numerische Mathematik, 104(4):445–456, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vecchio:2001:VDE

[Vec01] Antonia Vecchio. Volterra discrete equations: summability of the fundamental matrix. Numerische Mathematik, 89(4):783–794, October 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1089004/10890783.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1089004/10890783.pdf.

Veeser:2002:CAF

[Vee02] Andreas Veeser. Convergent adaptive finite elements for the nonlinear Laplacian. Numerische Mathematik, 92(4):743–770, October 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonGolitschek:2002:EBM

[vGLS02] Manfred von Golitschek, Ming-Jun Lai, and Larry L. Schumaker. Error bounds for minimal energy bivariate polynomial splines. Numerische Mathematik, 93(2):315–331, December 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vater:2009:SCG

[VK09] Stefan Vater and Rupert Klein. Stability of a Cartesian grid projection method for zero Froude number shallow water flows. Numerische Mathematik, 112(1):123–161, March 2009. CODEN NUMMA7. ISSN ???? URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=112&issue=1&spage=123.

Vohralik:2008:RFB

[Voh08] Martin Vohralík. Residual flux-based a posteriori error estimates for finite volume and related locally conservative methods. Numerische Mathematik, 111(1):121–158, November 2008. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=111&issue=1&spage=121.

Vovelle:2002:CFV

[Vov02] Julien Vovelle. Convergence of finite volume monotone schemes for scalar conservation laws on bounded domains. Numerische Mathematik, 90(3):563–596, January 2002. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
2090003/20900563.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/2090003/20900563.pdf.

Vila:2003:CEF

[VV03] Jean-Paul Vila and Philippe Villedieu. Convergence of an explicit finite volume scheme for first order symmetric systems. Numerische Mathematik, 94(3):573–602, May 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vandebril:2004:QMC

[VVM04] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. A QRmethod for computing the singular values via semiseparable matrices. Numerische Mathematik, 99(1):163–195, November 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vandebril:2009:RII

[VVM09] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. Rational QR-iteration without inversion. Numerische Mathematik, 113 (4):561–575, October 2009. CODEN NUMMA7. ISSN ???? URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=113&issue=4&spage=561.

Vandebril:2006:CPO

[VVVM06a] Raf Vandebril, Ellen Van Camp, Marc Van Barel, and Nicola Mastronardi. On the convergence properties of the orthogonal similarity transformations to tridiagonal and semiseparable (plus diagonal) form. Numerische Mathematik, 104(2):205–239, August 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vandebril:2006:OST

[VVVM06b] Raf Vandebril, Ellen Van Camp, Marc Van Barel, and Nicola Mastronardi. Orthogonal similarity transformation of a symmetric matrix into a diagonal-plus-semiseparable one with free choice of the diagonal. Numerische Mathematik, 102 (4):709–726, February 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=102&issue=4&spage=709.

Waldron:2000:MSE

[Wal00] Shayne Waldron. Minimally supported error representations and approximation by the constants. Numerische Mathematik, 85(3): 469–484, May 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085003/00850469.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085003/00850469.pdf.

Wang:2007:NCZ

[WcSX07] Ming Wang, Zhong ci Shi, and Jinchao Xu. A new class of Zienkiewicz-type non-conforming element in any dimensions. Numerische Mathematik, 106(2):335–347, April 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wegmann:2002:NCC

[Weg02] R. Wegmann. Numerical calculation of capillary-gravity waves. Numerische Mathematik, 92(2):383–400, August 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wensch:2001:EML

[Wen01] Jörg Wensch. Extrapolation methods in Lie groups. Numerische Mathematik, 89(3):591–604, September 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
1089003/10890591.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/1089003/10890591.pdf.

Wilderotter:2003:AFE

[Wil03] Olga Wilderotter. An adaptive finite element method for singular parabolic equations. Numerische Mathematik, 96(2):377–399, December 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wang:2002:AFD

[WL02] Cheng Wang and Jian-Guo Liu. Analysis of finite difference schemes for unsteady Navier–Stokes equations in vorticity formulation. Numerische Mathematik, 91(3):543–576, May 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=91&issue=3&spage=

Wang:2004:AFO

[WLJ04] Cheng Wang, Jian-Guo Liu, and Hangs Johnston. Analysis of a fourth order finite difference method for the incompressible Boussinesq equations. Numerische Mathematik, 97(3):555–594, May 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Washio:2000:EAP

[WO00] T. Washio and C. W. Oosterlee. Error analysis for a potential problem on locally refined grids. Numerische Mathematik, 86(3): 539–563, September 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0086003/00860539.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0086003/00860539.pdf.

Wang:2000:FDR

[WP00] Jianhua Wang and C. V. Pao. Finite difference reaction-diffusion equations with nonlinear diffusion coefficients. Numerische Mathematik, 85(3):485–502, May 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085003/00850485.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085003/00850485.pdf.

Wang:2006:TDF

[WP06] Yuan-Ming Wang and C. V. Pao. Time-delayed finite difference reaction-diffusion systems with nonquasimonotone functions. Numerische Mathematik, 103(3):485–513, May 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=103&issue=3&spage=485.

Wendland:2005:AIA

[WR05] Holger Wendland and Christian Rieger. Approximate interpolation with applications to selecting smoothing parameters. Numerische Mathematik, 101(4):729–748, October 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=101&issue=4&spage=729.

Wu:2008:SNR

[WS08] Jiming Wu and Weiwei Sun. The superconvergence of Newton– Cotes rules for the Hadamard finite-part integral on an interval. Numerische Mathematik, 109(1):143–165, March 2008. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weideman:2007:KPF

[WT07] J. A. C. Weideman and L. N. Trefethen. The kink phenomenon in Fejér and Clenshaw–Curtis quadrature. Numerische Mathematik, 107(4):707–727, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wang:2002:UAE

[WW02] Z. Wang and R. Wong. Uniform asymptotic expansion of J_{ν}(νa) via a difference equation. Numerische Mathematik, 91(1):147–193, March 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wang:2003:AES

[WW03] Z. Wang and R. Wong. Asymptotic expansions for second-order linear difference equations with a turning point. Numerische Mathematik, 94(1):147–194, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wang:2006:SIQ

[WY06] Zhou-Hong Wang and Ya-Xiang Yuan. A subspace implementation of quasi-Newton trust region methods for unconstrained optimization. Numerische Mathematik, 104(2):241–269, August 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Xu:2001:MME

[XC01] Xuejun Xu and Jinru Chen. Multigrid for the mortar element method for P1 nonconforming element. Numerische Mathematik, 88(2):381–398, April 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/1088002/10880381.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/1088002/10880381.pdf.

Xue:2000:BPT

[XG00] Jungong Xue and Weiguo Gao. Blockwise perturbation theory for block p-cyclic stochastic matrices. Numerische Mathematik, 84 (3):519–526, January 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-ny.

com/link/service/journals/00211/bibs/0084003/00840519.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/0084003/00840519.pdf.

Xiang:2007:EFT

[Xia07] Shuhuang Xiang. Efficient Filon-type methods for ∫_{a}^{b} f(x)e^{iωg(x)} dx. Numerische Mathematik, 105(4):633–658, February 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Xu:2000:CCF

[Xu00] Yuan Xu. Constructing cubature formulae by the method of reproducing kernel. Numerische Mathematik, 85(1):155–173, March 2000. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/0085001/00850155.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/0085001/00850155.pdf.

Xue:2009:ERP

[XY09] Jungong Xue and Qiang Ye. Entrywise relative perturbation bounds for exponentials of essentially non-negative matrices. Numerische Mathematik, 113(3):393–403, September 2009. CODEN NUMMA7. ISSN ???? URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=113&issue=3&spage=393.

Xu:2003:SOB

[XZ03] Jinchao Xu and Ludmil Zikatanov. Some observations on Babuška and Brezzi theories. Numerische Mathematik, 94(1):195–202, March 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yang:2002:ISM

[Yan02] Daoqi Yang. Iterative schemes for mixed finite element methods with applications to elasticity and compressible flow problems. Numerische Mathematik, 93(1):177–200, November 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:2001:PSP

[Yse01] Harry Yserentant. The propagation of sound in particle models of compressible fluids. Numerische Mathematik, 88(3):581– 601, May 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1088003/10880581.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1088003/10880581.pdf.

Yserentant:2004:RES

[Yse04] Harry Yserentant. On the regularity of the electronic Schrödinger equation in Hilbert spaces of mixed derivatives. Numerische Mathematik, 98(4):731–759, October 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:2005:SGS

[Yse05] H. Yserentant. Sparse grid spaces for the numerical solution of the electronic Schrödinger equation. Numerische Mathematik, 101(2):381–389, August 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=101&issue=2&spage=381.

Yserentant:2007:HCS

[Yse07] Harry Yserentant. The hyperbolic cross space approximation of electronic wavefunctions. Numerische Mathematik, 105(4):659– 690, February 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yun:2003:PIF

[Yun03] Jae Heon Yun. Performance of ILU factorization preconditioners based on multisplittings. Numerische Mathematik, 95(4):761–779, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yang:2006:QLS

[YW06] Danping Yang and Liqiu Wang. Quasi-least-squares finite element method for steady flow and heat transfer with system rotation. Numerische Mathematik, 104(3):377–411, September 2006. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zheng:2006:AFE

[ZCW06] Weiying Zheng, Zhiming Chen, and Long Wang. An adaptive finite element method for the H-Ψ formulation of time-dependent eddy current problems. Numerische Mathematik, 103(4):667–689, June 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2003:EEN

[Zha03] Yongmin Zhang. Error estimates for the numerical approximation of time-dependent flow of Bingham fluid in cylindrical pipes by the regularization method. Numerische Mathematik, 96(1):153–184, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2004:NRQ

[Zha04] Shangyou Zhang. On the nested refinement of quadrilateral and hexahedral finite elements and the affine approximation. Numerische Mathematik, 98(3):559–579, September 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2007:CFB

[Zha07] Yongmin Zhang. Convergence of free boundaries in discrete obstacle problems. Numerische Mathematik, 106(1):157–164, March 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zheng:2006:NSM

[Zhe06] Chunxiong Zheng. Numerical simulation of a modified KdV equation on the whole real axis. Numerische Mathematik, 105(2): 315–335, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2000:IEE

[ZK00] Jing Zhang and Fumio Kikuchi. Interpolation error estimates of a modified 8-node serendipity finite element. Numerische Mathematik, 85(3):503–524, May 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0085003/00850503.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0085003/00850503.pdf.

Zhang:2003:UZP

[ZL03] Zhimin Zhang and Runchang Lin. Ultraconvergence of ZZ patch recovery at mesh symmetry points. Numerische Mathematik, 95 (4):781–801, October 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2004:GCA

[ZLW04] J. Z. Zhang, G. S. Liu, and S. Y. Wang. A globally convergent approximately active search algorithm for solving mathematical programs with linear complementarity constraints. Numerische Mathematik, 98(3):539–558, September 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2000:NSC

[ZOS00] Shao-Liang Zhang, Yoshio Oyanagi, and Masaaki Sugihara. Necessary and sufficient conditions for the convergence of Orthomin(k) on singular and inconsistent linear systems. Numerische Mathematik, 87(2):391–405, December 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/service/journals/00211/bibs/
0087002/00870391.htm;   http://link.springer-ny.com/link/
service/journals/00211/papers/0087002/00870391.pdf.

Zarin:2005:IPD

[ZR05] Helena Zarin and Hans-Görg Roos. Interior penalty discontinuous approximations of convection-diffusion problems with parabolic layers. Numerische Mathematik, 100(4):735–759, June 2005. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=100&issue=4&spage=
735.

Sun:2004:SOA

[zSlZ04] Zhi zhong Sun and You lan Zhu. A second order accurate difference scheme for the heat equation with concentrated capacity. Numerische Mathematik, 97(2):379–395, April 2004. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:2002:NPT

[ZX02] Jianzhong Zhang and Naihua Xiu. New projection-type methods for monotone LCP with finite termination. Numerische Mathematik, 92(1):179–195, July 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhu:2003:STF

[ZZ03] Qiding Zhu and Qinghua Zhao. SPR technique and finite element correction. Numerische Mathematik, 96(1):185–196, November 2003. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhou:2009:NIM

[ZZ09] Shuzi Zhou and Zhanyong Zou. A new iterative method for discrete HJB equations. Numerische Mathematik, 114(1):159–167, November 2009. CODEN NUMMA7. ISSN ???? URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=114&issue=1&spage=159.

Zhang:2006:GCM

[ZZL06] Li Zhang, Weijun Zhou, and Donghui Li. Global convergence of a modified Fletcher–Reeves conjugate gradient method with Armijo-type line search. Numerische Mathematik, 104(4):561–572, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).